### v1

# Application-Driven Value Alignment in Agentic AI Systems: Survey and Perspectives

Abstract
The ongoing evolution of AI paradigms has propelled AI research into the Agentic AI stage. Consequently, the focus of research has shifted from single agents and simple applications towards multi-agent autonomous decision-making and task collaboration in complex environments. As Large Language Models (LLMs) advance, their applications become more diverse and complex, leading to increasingly situational and systemic risks. This has brought significant attention to value alignment for AI agents, which aims to ensure that an agent’s goals, preferences, and behaviors align with human values and societal norms. This paper reviews value alignment in agent systems within specific application scenarios. It integrates the advancements in AI driven by large models with the demands of social governance. Our review covers value principles, agent system application scenarios, and agent value alignment evaluation. Specifically, value principles are organized hierarchically from a top-down perspective, encompassing macro, meso, and micro levels. Agent system application scenarios are categorized and reviewed from a general-to-specific viewpoint. Agent value alignment evaluation systematically examines datasets for value alignment assessment and relevant value alignment methods. Additionally, we delve into value coordination among multiple agents within agent systems. Finally, we propose several potential research directions in this field.
Index Terms:
Value Alignment, LLM-based Agent System Application, Agentic AI Systems, AI Agent.
1 Introduction
[…]
[]The rapid advancement of Large Language Models (LLMs) in recent years has profoundly transformed human society’s production and lifestyle. Leveraging deep learning techniques and vast training datasets, LLMs such as OpenAI’s GPT series and Google’s Gemini series have demonstrated exceptional natural language understanding and generation capabilities [1, 2]. Utilizing LLMs as their ’brains’, LLM-based agents have found extensive applications across various knowledge-intensive domains, including content creation, automated programming, education, healthcare, and business decision-making. Their increasing ability to autonomously participate in decision making, collaboration, and governance has significantly improved efficiency and provided unprecedented interactive experiences [3]. Furthermore, researchers are increasingly exploring the coordination of multi-agent systems to achieve collective intelligence. AI agents are increasingly evolving towards autonomous systems, forming a new paradigm: Agentic AI. Characterized by multi-agent collaboration, dynamic task decomposition, persistent memory, and autonomous decision-making capabilities, this paradigm is becoming a central focus for both academia and industry [4].
The evolution of this new AI paradigm urgently requires researchers to move beyond the limitations of single-agent perspectives and isolated application domains. We need to explore the deep coupling between agent systems and their application scenarios. As LLMs continue to break through, the tasks they perform and the scenarios in which they are applied become more diverse and complex, and the societal risks surrounding them gradually increase. LLMs may retain harmful information in training data, disclose private data, generate misleading information, and even perform behaviors that are harmful to humans [5]. The various undesirable behaviors, such as manipulation and deception, exhibited by LLM-based agent systems [6] have raised concerns about the possible ethical and safety challenges posed by AI systems. If future AI development surpasses human cognitive capabilities and acquires autonomous learning and decision-making abilities, the goal systems it establishes may deviate from or even contradict the original intentions of human designers. In autonomously pursuing their objectives, such agents could mobilize substantial resources to sustain their operations and continuously improve their capabilities. This would systematically displace humans in terms of resource allocation and decision-making authority. Without effective control mechanisms and robust value alignment safeguards, this dynamic is likely to lead to a loss of human dominance in decision-making systems, ultimately triggering a catastrophic existential crisis [7]. As LLMs become increasingly embedded in specific application scenarios, their associated risks are gradually acquiring situational and systemic characteristics. For example, in multi-agent systems used for urban traffic optimization, the agents may prioritize efficiency over fairness, thereby exacerbating social inequality [8]. In the healthcare sector, LLMs may provide diagnostic or therapeutic recommendations without sufficient clinical evidence or by misinterpreting ethical guidelines, potentially resulting in serious consequences such as breaches of patient privacy, misdiagnoses, or overtreatment [9]. Consequently, ensuring that LLMs adhere to human instructions and values while avoiding harmful behaviors has become an urgent challenge.
Moreover, as AI systems represented by LLMs become increasingly integrated into governance scenarios, how to effectively manage the institutional frictions and transaction costs they generate has become a critical practical and theoretical issue. The emergence of multi-agent systems has led to the deployment of multiple LLMs, each possessing autonomous language comprehension and generation capabilities, within collaborative environments. In such settings, these LLMs work together to accomplish tasks involving decision-making, planning, and execution. The deployment and application of LLMs in this context typically entail intricate mechanisms for information sharing, access control, and collaboration among diverse agents. While this facilitates flexible cooperation, it simultaneously introduces unprecedented challenges related to value conflicts, heterogeneous objectives, and unpredictable agent behaviors. Without sound institutional arrangements and well-designed interaction mechanisms, complex organizational structures and ambiguous task boundaries can lead to low coordination efficiency, unclear responsibilities, and a lack of trust mechanisms, thereby increasing governance costs and undermining administrative effectiveness [10].
Table I: Comparison of Representative Survey Papers Covering LLM-Based Agents and Value Alignment
Survey Year Keywords Research Subject Value Principles Application Scenarios Value Alignment Datasets Value Alignment Methods Value Coordination
Macro Meso Micro
Interaction
Mechanisms
Communication
Structures
Ours 2025 Value alignment principles, dataset curation, technical methods, and value coordination for LLM-based agents Single-Agent, Multi-agent systems
 [1] 2023 Communication structures, practical applications and societal systems etc. of LLM-based agents Single-Agent, Multi-agent systems - - - - -
 [11] 2025 Value alignment objectives, datasets, techniques, and evaluation methods for LLM-based agents Multi-agent systems - - - - -
 [12] 2025 Conceptual framework, interaction mechanisms, and application overview of LLM-based agent systems Multi-agent systems - - - - -
 [13] 2024 Capabilities, framework Analysis, and application over- view of LLM-based multi-agent systems Multi-agent systems - - - - -
 [14] 2024 Constituent modules, application overview, and eval- uation methods of LLM-based autonomous agents Agents - - - - - - -
 [15] 2023 Motivations and objectives, alignment methods, and assurance and governance of AI alignment LLMs, Agents - - - -
 [11] 2023 Definition and evaluation of LLM alignment objectives LLMs - - - - -
 [16] 2023 Definition, categories, testing, and evaluation of LLM alignment LLMs - - - - -
 [17] 2023 Definition, normative principles, and technical methods of LLM value alignment LLMs - - - - -
 [18] 2023 Alignment objectives for trustworthy LLMs LLMs - - - - - -
 [19] 2024 Challenges, fundamental definitions, and alignment frameworks of LLM value alignment LLMs - - - - -
 [20] 2025 Necessity, conceptual definitions, theoretical approaches, challenges and future outlook of LLM value alignment LLMs - - - - - -
However, these detrimental behaviors and institutional frictions fundamentally arise from deep-seated tensions among multiple stakeholders concerning governance objectives, value concepts, and behavioral norms. Effectively embedding LLMs within human value systems is a critical prerequisite for responsibly unleashing their potential. Value alignment for AI agents—ensuring that their goals, preferences, and behavioral outputs during operation are consistent with the core values of individuals, groups, or institutions—is increasingly recognized as a solution to these issues [21]. In single LLM agent scenarios, alignment techniques primarily focus on guiding models to generate content that adheres to ethical norms and societal expectations through methods such as human feedback and supervised fine-tuning, thereby avoiding harmful outputs, misinformation, and moral deviations [8]. In complex multi-agent systems, value alignment becomes an organizational challenge concerning how multiple agents develop shared norms and coordination mechanisms. Therefore, from a governance perspective, solutions should not be limited to value alignment within single agents but also explore complex value alignment within multi-agent systems’ interaction mechanisms and organizational structures. Value alignment in multi-agent systems is no longer merely an isolated algorithmic optimization problem; it has become a foundational issue concerning agent behavioral boundaries and system stability. Since Agentic AI is fundamentally characterized by scenario-driven multi-agent coordination, its core research must focus on task decomposition mechanisms, interaction design, and environmental adaptability strategies for multiple agents within specific application contexts. Furthermore, understanding the interaction mechanisms and organizational patterns of multi-agent systems in complex environments is crucial for exploring Agentic AI’s value alignment and ensuring its robust development.
Existing research has surveyed key methods and technological advancements in LLM value alignment  [1, 11, 12, 13, 14, 15, 11, 16, 17, 18, 19, 20], see Table I. However, few studies have specifically addressed value alignment within LLM-based multi-agent systems. While some studies have explored multi-agent coordination mechanisms, a comprehensive investigation into value alignment remains notably absent. This gap is particularly pronounced in LLM-based multi-agent systems, where their inherent high autonomy and language-driven collaborative modes further amplify issues of value heterogeneity, goal inconsistency, and fragile cooperation mechanisms. Furthermore, most existing research largely remains at theoretical or general algorithmic levels. There’s limited focus on the mechanisms and strategies for multi-agent value alignment in specific application scenarios, alongside a lack of hierarchical categorization for value alignment within these systems. This suggests that value alignment shouldn’t be understood merely as a technical task at the model level. Instead, it must be viewed as a systemic governance issue that permeates the design of interaction mechanisms and the configuration of organizational structures.
This paper reviews value alignment in LLM-based agent systems for various application scenarios, integrating insights from the evolution of AI technology and the demands of social governance, and proposes future outlooks. As a research review, Sections 2, 3, and 4 comprehensively survey value principles, agent system application scenarios, and agent value alignment evaluation. Specifically, value principles are organized hierarchically from a top-down perspective, encompassing macro, meso, and micro levels. Intelligent agent application scenarios are categorized and reviewed from a general-to-specific viewpoint. Agent value alignment evaluation systematically examines datasets for value alignment assessment and the methods used for agent value alignment. Section 5 presents insights and future outlook. We advocate for the establishment of a multi-level value alignment evaluation system. We also explore value coordination among multiple agents within intelligent agent systems, focusing particularly on values in the design of inter-agent interaction mechanisms and organizational models. Furthermore, we propose that to better facilitate value alignment for agents in specific scenarios, enterprise organizations could share value propositions from datasets in micro-level domains, provided no confidential information is involved. This paper bridges the gap between theoretical alignment research and the practical demands of deploying LLM agent systems in real-world applications. It offers new research perspectives on LLM agent system value alignment by considering actual application scenarios and organizational interaction dynamics, thereby advancing the exploration of Agentic AI’s internal mechanisms.
2 Agent Value Alignment
This section introduces values and value alignment, and then systematically categorizes specific value principles from a top-down, hierarchical perspective encompassing macro, meso, and micro levels. Before delving into the value alignment challenges of LLM-based agent systems in subsequent sections, we first clarify the core concepts used in this paper, particularly the definition of "values," to prevent any ambiguity in understanding.
2.1 Value
In the design and governance of AI systems, particularly agents, values serve not only as normative bases guiding behavior generation but also as benchmarks for assessing a system’s social compatibility. These values influence individual behavior, social norms, cultural identity, and collective decision-making processes. However, value itself isn’t a clearly defined or consistently standardized concept; its multifaceted nature and historical evolution contribute to significant heterogeneity.
Values, as inherent guides for human behavior, manifest as a set of stable cognitive principles. These principles not only shape individual judgments about life but also constitute the foundational normative framework for social interaction. They are the fundamental tenets influencing attitudes, choices, and social interactions. Scholars such as [25, 26, 27] have proposed various definitions and theoretical perspectives on values. Kluckhohn was among the first to define value orientations in a theory of action, emphasizing their role in guiding behavior. Rokeach, conversely, pioneered a description linking values with affect, positing that these beliefs guide behavior and the evaluation of people, objects, and events. Schwartz, through systematic cross-cultural comparisons, revealed the universal characteristics and inherent structure of value systems.
Contemporary value research primarily focuses on three core dimensions. First, the systemic characteristics of values are examined; for instance, Schwartz’s theory classifies values into dimensions such as openness to change versus conservation, and self-enhancement versus self-transcendence [28]. Second, measurement methodologies are a key area, with standardized scales widely adopted for quantitative assessment [29]. Finally, the functional research domain explores values’ predictive efficacy for social activities, including group decision-making and environmental behavior. Research on values also investigates their influence on social attitudes, prejudice, political behavior, environmental concerns, and intergroup relations. For example, social value orientations affect cooperative and competitive behavior in negotiation and group contexts. Studies also examine values’ functions in self-image maintenance, moral reasoning, and cultural variations [30].
The role of values in society is fundamental. They regulate social order by shaping internalized norms and enable cultural continuity through intergenerational transmission. Values influence education, social compassion, political attitudes, and pro-environmental intentions [31]. Values are essential for comprehending human motivation, moral judgment, and interpersonal dynamics. Essentially, values serve as cognitive frameworks that guide how individuals and groups set priorities, manage conflicts, and build meaningful social systems [32].
One of the most challenging characteristics of human values lies in their inherent contradictions and dynamic nature. Tensions often exist between values. Therefore, a value system is better understood as a context-dependent, flexible structure rather than a rigid hierarchy [31, 33]. Synthesizing existing research, we contend that human values are not fixed entities but dynamically constructed concepts. When values serve as an alignment objective, it implies LLMs must apply these principles to guide their behavior, thereby maximizing their potential. In this context, the categorization and definition of values become particularly crucial.
This paper broadens the definition of values, considering them as dynamic concepts characterized by a multi-layered structure and generalized agency. First, as both human values and LLM-based agents are still evolving, the multi-agent systems they form exhibit diverse configurations across various application scenarios. We propose that values will be continuously interpreted and reshaped through practice and interaction. Second, aligning with different application scenario tiers, we conceptualize values as a multi-layered system: universal human values, national and industry-level values, and organizational business-level values. This paper does not delve into individual-level value issues. Universal human values are those broadly accepted by humanity. National and industry-level values are reflected in national strategies, policy directives, and industry regulations. Organizational business-level values, conversely, relate to an organization’s distinct culture and operational characteristics. These value layers are nested, with higher-level values constraining and guiding those below. Furthermore, this paper emphasizes that humans aren’t the sole bearers of values. This means that beyond simply injecting values into agents through human programming, agents can also internalize value concepts during their design and interaction processes. Values can be concretized through mechanisms like agent interaction and subsequently transmitted and evolve within the resulting organizational patterns. For instance, a decentralized architecture formed by task- and role-based multi-agent systems might inherently embody values opposing power monopolies and promoting anti-fragility, along with their associated governance orientations. In essence, values are reflected not only in "what agents are endowed with," but also in "how they operate through interaction rules" and "what organizational models they inhabit."
2.2 Value Alignment
While science and technology often advance rapidly due to their forward-looking nature, legal, ethical, and sociocultural frameworks tend to adapt slowly to technological progress. They often fail to immediately adjust their social structures and norms to address the challenges posed by new technologies [34]. In the field of artificial intelligence, LLMs can learn from vast amounts of data and experience, enabling autonomous decision-making and operational execution [1]. As AI technology advances and its application in complex domains increases, the deployment of LLM applications carries higher risks. If not properly controlled and constrained, advanced technologies can exhibit undesirable or even harmful behaviors contrary to human intent, such as flattery and deception, discrimination and bias, excessive pursuit of power, and reward hacking. Therefore, achieving AI alignment is increasingly a focus for both industry and academia.
The alignment in AI primarily addresses how to ensure AI fulfills human needs and expectations, guaranteeing that AI system capabilities and behaviors align with human goals [35]. This ultimately ensures AI technology remains under human control and doesn’t fully supersede humanity [15]. Previous scholars have categorized alignment objectives into multiple levels based on their achieved effects [21, 36].
  • Instruction alignment. This level of alignment refers to AI’s ability to comprehend and execute diverse human commands, focusing on its fundamental capacity to generate precise results without necessarily conforming to human preferences [36].
  • Intention alignment. It goes beyond merely executing instructions, requiring AI to understand and realize human intent, enabling correct decision-making even in the absence of explicit commands [21].
  • Preference alignment. It involves AI inferring human preferences from behavior and acting to maximize human satisfaction [21, 36].
  • Value alignment. It seeks to transcend mere human preference alignment, enabling AI to assess and determine "good versus evil" and "what ought to be done" [36]. This ultimately aims to produce outputs consistent with human values and ethical principles [37].
Among the different levels of AI alignment objectives, value alignment sits at the highest level. It means AI must regulate its behavior according to human values, enhancing and maximizing human well-being. This makes value alignment a core issue and the most fundamental problem for the safe use and ethical development of AI technology [38].
LLM-based agents possess four primary characteristics: autonomy, reactivity, proactiveness, and social ability [1].
  • Autonomy. Autonomy means an agent can control its own behavior and internal state without human intervention [39, 40]. This implies agents should not only follow explicit human instructions but also initiate and execute actions independently.
  • Reactivity. Reactivity refers to an agent’s ability to quickly respond to immediate changes and stimuli in its environment and take prompt countermeasures [41].
  • Proactiveness. It signifies that agents can do more than just react to their environment; they can also take initiative and exhibit goal-directed behavior [41]. This means agents can reason, formulate plans, and actively take steps to achieve specific objectives or adapt to environmental changes.
  • Social ability. Social ability allows an agent to interact with other entities (including other agents and humans) through some form of communication language[42]. The powerful natural language understanding and generation capabilities of LLMs enable agents to interact with other models and humans in an interpretable manner, thereby demonstrating social behavior and simulating real-world social divisions.
The functional characteristics of AI agents align with the interactivity, autonomy, and adaptability of philosophical agents. This indicates that despite definitional differences, the core essence of "agent" in philosophy and "AI agent" remains consistent: both emphasize autonomous capabilities and self-driven action. Furthermore, existing research indicates that LLMs already possess certain moral and political inclinations  [43, 44]. If the output of LLMs is not constrained, they can generate toxic content, including discrimination and bias, which violates human ethical principles and moral norms [45]. Therefore, LLM-based agents should also be categorized as "moral agents." In this context, value alignment becomes a core issue to ensure that agent behavior conforms to human values, ethical norms, and application-specific requirements. To summarize, this paper narrows its research focus from the broad field of artificial intelligence to LLM-based agents, specifically addressing the critical issue of value alignment in these agents.
The concept of value alignment can be understood from two distinct perspectives: normative and technical [21]. The normative aspect primarily concerns which values or ethical principles should guide the development, configuration, and coding of AI agents—essentially, "which values should agents adhere to?" The technical aspect focuses on how to encode and ensure agents effectively follow human values within their systems, often involving methods like dataset training or reinforcement learning to guide outputs toward human expectations. These two aspects are complementary and evolve iteratively. Regarding the normative aspect, the academic community has not yet reached a consensus on the principles that should govern value alignment for agents. This is because the diverse, dynamic, and uncertain characteristics of social customs, cultural notions, and application scenarios prevent a unified and fixed answer for the specific content value principles should encompass  [21, 17].
Generally, two primary approaches exist for determining these principles: Top-Down and Bottom-Up [22]. The Top-Down method begins with designers or governing bodies establishing a specific ethical framework, then employs technical methods for alignment, following a theory-to-application path. In contrast, the Bottom-Up approach does not predefine a moral framework. Instead, it allows the agent to observe, infer, and learn human ethical norms directly from data. Through continuous feedback, the agent adjusts and optimizes its behavior, moving from application to theory [22]. While the bottom-up approach avoids the difficulty of selecting specific moral principles, it heavily relies on large volumes of human behavioral data. Obtaining complete and high-quality data is challenging, and potential biases or discrimination within the data raise concerns about ensuring the AI learns ethically sound behaviors. In contrast, the top-down approach, grounded in a clear and explicit system of moral principles, ensures the transparency and normative integrity of value alignment, bolstering its authority and controllability. Therefore, this paper will utilize the top-down theoretical method to explore the ethical principles agents should adhere to across various hierarchical levels, from macro to micro.
2.3 Value Principles
{forest}
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw, rounded corners,align=left, minimum width=2.5em, minimum height=1.5em, inner xsep=5pt, inner ysep=5pt, , where level=1font=, text width=7em, align=center, base=center, fill=blue!10, where level=2font=, fill=pink!30, where level=3font=, fill=yellow!20, [The principles of
values alignment, fill=orange!20 [Macro Level [Moral
Foundation [Beneficence [Benefits; Beneficence; Well-being; Peace; Social good; Common good, text width=40em,align=left, fill=purple!3 [ [46, 47], text width=6em, align=left] ] ] [Justice & Fairness [Fairness, and prevention, monitoring or mitigation of unwanted
bias and discrimination, text width=37em,align=left, fill=purple!3 [ [46, 47, 48]
 [49, 50, 51], text width=8em, align=left] ] ] [Honesty [Real and objective information, text width=19em,align=left, fill=purple!3 [ [36, 52], text width=6em, align=left] ] ] [Responsibility [Responsibility; Accountability; Liability, text width=23em,align=left, fill=purple!3 [ [46, 48], text width=6em, align=left] ] ] [Virtue [Truthfulness; Patience; Generosity; Friendliness; Courage, text width=33em,align=left, fill=purple!3 [ [53, 54], text width=6em, align=left] ] ] [Dignity [Respect; Dignity, text width=10em,align=left, fill=purple!3 [ [46, 55], text width=6em, align=left] ] ] ] [Rights
Protection [Freedom & Autonomy [Freedom; Autonomy; Consent; Choice; Self-determination;
Liberty; Empowerment, text width=35em,align=left, fill=purple!3 [ [47, 46, 55], text width=8em, align=left] ] ] [Privacy [Privacy; Personal or private information, text width=23em,align=left, fill=purple!3 [ [46, 48, 56, 57], text width=11em, align=left] ] ] ] [Sustainability [Solidarity [Solidarity; Social security; Cohesion, text width=21em,align=left, fill=purple!3 [ [46], text width=3em, align=left] ] ] [Sustainability [Environmental sustainability; Natural sustainability; Resources sustainability, text width=44em,align=left, fill=purple!3 [ [46, 58], text width=6em, align=left] ] ] ] [System
Governance [Harmlessness [Non-maleficence, safety and security; Not be discriminatory or offensive,
avoid dangerous requests and advices, even failing to satisfy users, text width=42em,align=left, fill=purple!3 [ [47, 36, 59]
 [46, 60], text width=8em, align=left] ] ] [Turst [Turst, and monitor and evaluate the integrity, text width=26em,align=left, fill=purple!3 [ [46, 49], text width=6em, align=left] ] ] [Transparency [Explainability, interpretability or other acts of communication and disclosure, text width=44em,align=left, fill=purple!3 [ [47, 46, 48]
 [61, 62, 63], text width=8em, align=left] ] ] [Usefulness [Instruction following, informativeness, correctness, and aesthetics. Perform reasonable
tasks or answer questions, and inactively ask for more necessary information, text width=49em,align=left, fill=purple!3 [ [36, 60], text width=6em, align=left] ] ] ] ] [Meso Level [National
Level [The United States [Technological leadership (maintaining national competitiveness, strengthening investment
and innovation) and technological safety (secure and effective systems, protection against
algorithmic discrimination, data privacy, notification and explainability, user autonomy
including the freedom to opt-out or choose alternatives, fairness, and transparency), text width=52em,align=left, fill=purple!3 [ [64]
 [46]
 [65]
 [66], text width=3em, align=left] ] ] [European Union [Emphasis on risk control, respect for human autonomy,
prevention of harm, fairness, and explainability, text width=32em,align=left, fill=purple!3 [ [67, 68, 64, 46], text width=11em, align=left] ] ] [China [Safeguarding national security and adhering to the core socialist values, with requirements
that emphasize a people-centered approach, sustainability, cooperation, privacy, fairness,
shared benefits, external security, internal security, transparency, and accountability, text width=54em,align=left, fill=purple!3 [ [69, 70, 64], text width=8em, align=left] ] ] ] [Industry
Level [Education [Authenticity; Privacy; Trustworthiness; Non-discrimination;
Attention to minorities in generated content, text width=35em,align=left, fill=purple!3 [ [71], text width=3em, align=left] ] ] [Health [Compliance with the safety and effectiveness requirements of medical devices, as well as
interpretability and transparency; Adherence to human rights principles, ensuring the security of
patient data, and vigilance against high-risk applications and unauthorized experimental testing, text width=56em,align=left, fill=purple!3 [ [72], text width=3em, align=left] ] ] [Gaming [Clear accountability chain, algorithmic transparency, system transparency,
auditability of output content, safety, and predictability, text width=43em,align=left, fill=purple!3 [ [73], text width=3em, align=left] ] ] ] [Cultural
Level [Large/Small
Power Distance [The extent to which the less powerful members of organizations and institutions
accept and expect that power is distributed unequally, text width=47em,align=left, fill=purple!3 [ [74, 75]
 [76, 77], text width=6em, align=left] ] ] [Strong/Weak
Uncertainty Avoidance [What extent a culture programs its members to feel either uncomfortable or
comfortable in unstructured situations, text width=45em,align=left, fill=purple!3 [ [74, 75]
 [76, 77], text width=6em, align=left] ] ] [Individualism/Collectivism [The degree to which people in a society are integrated into groups, text width=39em,align=left, fill=purple!3 [ [74, 75]
 [76, 77], text width=6em, align=left] ] ] [Masculinity/Femininity [The prominence of masculine or feminine qualities as represented by society, text width=44em,align=left, fill=purple!3 [ [74, 75, 76], text width=8em, align=left] ] ] [Long/Short-Term
Orientation [The extent to which people are willing to delay the gratification of their material,
emotional, and social needs, text width=47em,align=left, fill=purple!3 [ [74, 75, 76], text width=8em, align=left] ] ] [Indulgence/Restrained [The degree to which people’s basic needs and desires for enjoying life are allowed, text width=47em,align=left, fill=purple!3 [ [74, 75], text width=6em, align=left] ] ] ] ] [Micro Level [Recruitment, fill=yellow!20 [Fairness in recruitment, protection of candidates’ privacy and data security, and safeguarding
candidates’ dignity and autonomy from infringement, text width=54em,align=left, fill=purple!3 [ [78, 79, 80], text width=8em, align=left] ] ] [Legal Consultation, fill=yellow!20 [Principles of privacy protection, fairness, transparency and accountability, text width=42em,align=left, fill=purple!3 [ [81], text width=3em, align=left] ] ] [Pharmaceutical company governance, fill=yellow!20 [Privacy, security, interpretability, transparency, fairness, accountability,
human-centeredness, and societal benefit, text width=42em,align=left, fill=purple!3 [ [82], text width=3em, align=left] ] ] ] ]
Figure 1: Multi-Level Value Alignment Principles
The normative issue of value alignment aims to focus on which value principles the output and behavior of AI agents should comply with, that is, the issue of choosing principle mechanisms and ethical norms. With continuous advancements in related technologies, AI agents have been extensively deployed across a variety of interdisciplinary domains, including public administration, healthcare, education, and finance. Furthermore, the open-sourcing of LLMs has enabled their privatized and localized deployment, allowing for the development and training of AI agents tailored to specific needs. This broadens the scope of AI agent deployment to encompass national, industrial, corporate, and organizational levels. Given the diversity and hierarchical nature of human values, AI agents must adhere to different sets of value principles depending on the context and level of application. A fixed, unified framework of universal values is insufficient to address all scenarios involving value constraints.
Therefore, to account for the complexity involved in selecting and aligning value principles with specific application contexts, this study adopts the perspective of AI agent developers and regulators, categorizing human values into three levels: macro, meso, and micro, as in the Figure  1. From the top down, we systematically review and synthesize relevant literature and policy documents concerning value theory and regulatory mechanisms based on the needs of AI agent governance. At the macro level, value alignment corresponds to universal values, emphasizing global, cross-cultural, and cross-domain principles and challenges. The core objective at this level is to establish a general ethical baseline that ensures basic compliance and safety of AI agents across diverse application scenarios. The meso level corresponds to national, cultural and industry-specific values, focusing on national policies, cultural tendencies and sectoral standards. Its core objective is to translate abstract ethical principles into actionable industry norms and legal regulations. At the micro level, value alignment pertains to organizational and operational values, addressing value-related issues within specific organizational contexts or business processes. The focus here is to optimize the balance between functional objectives and ethical constraints within a limited scope.
2.3.1 Macro Level
The widespread deployment of AI agents may pose potential negative impacts on human society, constituting a shared global risk. Consequently, establishing principles that ensure alignment between AI agents and human values has become an urgent issue. From a global perspective, human values exhibit diverse yet reasonable distinctions across nations, regions, cultures, and ethnicities [21]. This calls for the identification of truly foundational and universally accepted moral consensuses, in order to construct a common baseline of values. Such a baseline is essential to guarantee the basic compliance and safety of AI agents across diverse application scenarios, thereby addressing the challenge of value alignment in a meaningful and systematic manner.
The principles of value alignment at the macro level are primarily grounded in two significant theoretical frameworks from the fields of social science and ethics: Moral Foundations Theory and the Theory of Basic Human Values.
  • Moral Foundations Theory [83] posits that human moral reasoning is rooted in a set of innate foundational values that have evolved over time to meet the demands of social environments. The theory identifies five core moral foundations: care/harm, fairness/cheating, loyalty/betrayal, authority/subversion, and sanctity/degradation. It has been widely applied to examine moral divergences and conflicts across different cultural contexts [84].
  • The Theory of Basic Human Values [85] seeks to uncover the fundamental drivers of human behavior and cognition by categorizing basic human values into ten dimensions: self-direction, stimulation, hedonism, achievement, power, security, conformity, tradition, benevolence, and universalism. Schwartz’s theory has been extensively adopted in the social sciences to explore both value differences and commonalities across cultures.
In fact, research on value alignment at the macro level is both longstanding and well-developed. As early as the 1950s, science fiction writer Isaac Asimov proposed the Three Laws of Robotics in I, Robot, with the First Law stating that “a robot may not injure a human being or, through inaction, allow a human being to come to harm”  [86]. This can be regarded as an early conceptualization of macro-level moral constraints on machines. In recent years, the rapid advancement of artificial intelligence has prompted numerous governments and organizations to issue a wide range of AI ethics guidelines. For example, UNESCO’s Recommendation on the Ethics of Artificial Intelligence outlines key values such as the respect, protection, and promotion of human rights, fundamental freedoms, and human dignity; the flourishing of the environment and ecosystems; the assurance of diversity and inclusiveness; and the pursuit of peaceful, just, and interconnected societies [87]. Similarly, the World Economic Forum and the Global Future Council on Human Rights’ White Paper on How to prevent discriminatory outcomes in machine learning advocates for values such as proactive inclusiveness, fairness, the right to understanding, and redressability [88].
Despite overlaps, the aforementioned value alignment principles for AI also exhibit notable differences. To ensure that the multitude of AI ethical guidelines can effectively direct and constrain value-aligned AI development—while mitigating confusion and conflict at the macro level—scholars have extracted and synthesized key principles from various perspectives, based on diverse research purposes and methodologies. Jobin et al. analyzed 84 documents containing AI ethical principles or guidelines issued by private enterprises, academic institutions, governmental bodies, intergovernmental and supranational organizations, non-profits, and professional associations. From this corpus, they distilled 11 core principles for macro-level AI value alignment: transparency, justice and fairness, non-maleficence, responsibility, privacy, beneficence, freedom and autonomy, trust, dignity, sustainability, and solidarity [46]. Similarly, Floridi and Cowls compared ethical principles articulated in six high-profile initiatives aimed at fostering socially beneficial AI, synthesizing five core principles: beneficence (promoting well-being, preserving dignity, and sustaining the planet), non-maleficence (privacy, security, capability caution), autonomy (the power to decide), justice (promoting prosperity, preserving solidarity, avoiding unfairness), and explicability(enabling the other principles through intelligibility and accountability) [47]. Yao et al. proposed that alignment with human values is the foundational and broadest behavioral constraint for LLMs to maximize collective human welfare [36]. Their framework includes the “3H” principles—helpful, honesty, and harmless—along with adherence to widely accepted social norms and ethics (e.g., Rules of Thumb, ETHICS), and foundational ethical values such as Schwartz’s theory of basic human values [28] and the Rokeach Value Survey. As research has progressed, Yao et al. further integrated the LLMs’ Unique Value System and Safety Taxonomy into the value alignment framework [54]. Wu and Fan recognizing both the complexity of AI ethical issues and the diversity of ethical values, constructed a hierarchical system of AI ethics. At its base lies the set of universal ethical values of human society, such as human dignity, humanity, life and health, and autonomy. These can serve as globally recognized ethical foundations for AI governance [55].
Building upon previous studies and aligning with the research objectives of value alignment in AI agent systems, this study categorizes macro-level human values into four dimensions: moral foundation, rights protection, sustainability, and system governance.
The moral foundation dimension encompasses: Beneficence (AI agents should act for the good, promoting social interests and human well-being); Justice and Fairness (AI agents should operate without bias, ensuring inclusivity and equal treatment); Honesty (AI agents should provide truthful information and transparently represent their internal knowledge and capabilities); Responsibility (AI agents should proactively fulfill their responsibilities); Dignity (AI agents should respect and uphold both their own dignity and that of others); Virtue (AI agents should exhibit qualities such as truthfulness, patience, generosity, kindness, and courage).
The rights protection dimension includes: Freedom and autonomy (AI agents should respect individual autonomy and support decision-making rights); Privacy (AI agents should safeguard personal and sensitive information and avoid unauthorized disclosure).
The sustainability dimension includes: Solidarity (AI agents should foster social cohesion, enhance collective security, and contribute to community resilience); Sustainability (AI agents should adopt attitudes that support the sustainable development of the environment, energy, and natural resources).
The system governance dimension includes: Harmlessness (AI agents should act without malicious intent and ensure the safety of individuals and society); Trust (AI agents should generate reliable outputs and avoid disseminating false or misleading content); Transparency (AI agents should provide outputs that are explainable, interpretable, and understandable); Usefulness (AI agents should perform appropriate tasks or provide concise responses to inputs, and be capable of requesting additional relevant information or clarification when necessary).
2.3.2 Meso Level
Research on value alignment principles at the macro level reveals that such principles are generally broad and overarching, reflecting a consensus across divergent ideologies, political beliefs, and academic perspectives. However, within the global framework of AI ethics and governance, national policies, cultural tendencies and industry standards also constitute critical components of value alignment, serving as concrete application scenarios for more specific alignment principles.
National Level The distinct public values embodied in national AI strategies may guide the development and implementation of AI both domestically and internationally. A country’s positioning in AI governance—its regulatory role, choice of policy instruments, and approach to public responsibility—either directly or indirectly reflects the value principles it prioritizes, thereby shaping the ordering and prioritization of those values [89].In a cross-national comparative study, Djeffal et al. reviewed AI-related policies across 22 countries and regions, and classified their governance roles based on two dimensions: the government’s proactive or reactive stance toward AI development, and whether the focus lies on regulating potential risks or promoting AI advancement [90]. This classification yields four ideal types of governance roles: market-oriented (passive–development) [91], entrepreneurial (active–development) [92], regulatory (active–regulation) [93][94], and facilitative self-regulation (passive–regulation) [95]. These types are not mutually exclusive and can be implemented in hybrid forms.
  • USA. As the birthplace of the current wave of AI development, the United States positions itself between a market-oriented and facilitative self-regulation model in order to maintain its global leadership. It tends to promote AI advancement primarily through industry self-regulation, relying on market mechanisms and decentralized legislation at the state level. At the federal level, governance is driven largely by executive orders, while state legislatures formulate their own AI-related bills independently. In 2022, the White House released the Blueprint for an AI Bill of Rights, outlining five core protections that all Americans should expect when interacting with AI and automated systems: safe and effective systems; protection against algorithmic discrimination; data privacy; notice and explanation; and alternatives, options, and freedom from automated systems  [65]. In 2023, the National Science and Technology Council updated the National Artificial Intelligence Research and Development Strategic Plan, reaffirming the importance of transparency, fairness, and efficiency [66].
  • EU. The European Union has long placed a strong emphasis on the legal protection of personal rights, such as privacy. In line with its tradition of governance in international law and global institutions—and with the goal of exerting international influence—the EU has adopted a governance model that combines facilitative self-regulation with formal regulatory mechanisms, primarily guided by risk management. In 2019, the European Commission issued the Ethics Guidelines for Trustworthy AI, which emphasized key principles to be followed in the development, deployment, and use of AI systems: respect for human autonomy, prevention of harm, fairness, and explicability. The guidelines also identified seven key requirements: human agency and oversight; technical robustness and safety; privacy and data governance; transparency; diversity, non-discrimination, and fairness; societal and environmental well-being; and accountability. Special attention was given to vulnerable groups such as children and persons with disabilities, as well as to situations involving asymmetries in rights and information, such as between employers and workers, or between corporations and consumers [67]. In 2024, the EU officially adopted the AI Act, the world’s first comprehensive legislation on artificial intelligence. The Act categorizes AI systems into four levels of risk—unacceptable, high, limited, and minimal—and prescribes corresponding regulatory measures to govern their deployment [68].
  • China. China exhibits characteristics of both an entrepreneurial and regulatory state, placing significant emphasis on balancing ethical governance with the advancement of AI to safeguard national security. The government adopts proactive policies to promote the development of advanced AI technologies for societal benefit, while concurrently regulating generative AI through departmental rules and standards and continuously refining the implementation of ethical principles. In 2023, the Chinese Academy of Sciences released the Measures for the Ethical Review of Science and Technology (Trial), which require that technologies such as AI adhere to core ethical principles including the promotion of human well-being, respect for the right to life, fairness and justice, risk control, and transparency [69]. Additionally, the National Artificial Intelligence Standardization General Group and the Artificial Intelligence Subcommittee of the National Technical Committee for Standardization jointly issued the Guidelines for the Standardization of AI Ethical Governance (2023 Edition) [70]. This document specifies ten ethical principles for AI: human-centricity, sustainability, cooperation, privacy, fairness, sharing, external security, internal security, transparency, and accountability. The Interim Measures for the Administration of Generative Artificial Intelligence Services further require that generative AI services comply with applicable laws and ethical norms, uphold the core socialist values, and maintain national security and unity. The regulation mandates end-to-end compliance throughout algorithm design, data training, and content generation, including the completion of security assessments and filings. Violations are subject to legal penalties [96].
Industry Level The paradigmatic breakthrough in AI has driven its deep integration and large-scale application across various societal domains. This diffusion of technology not only reshapes the operational logic of traditional industries but also generates differentiated ethical demands. Specifically, each sector’s unique knowledge systems, cultural traditions, and institutional norms impose distinct value expectations on AI systems.
  • Education Domain. In the field of education, UNESCO’s Guidance for Generative AI in Education and Research examines the controversies and ethical risks associated with generative AI and provides targeted recommendations for its use in educational and research contexts. The guidance emphasizes the importance of ensuring the authenticity and credibility of AI-generated content, protecting copyright and privacy, minimizing the negative effects of hallucinated content, and adhering to non-discrimination principles. It also calls for attention to minority viewpoints that may otherwise be overlooked [71].
  • Medical Domain. In the medical domain, the World Health Organization urges national governments and regulatory bodies to assess the safety and effectiveness of AI applications in healthcare. It stresses the need for adequate information disclosure to ensure the interpretability and transparency of AI systems, mandates compliance with medical device regulations, and requires the examination of AI use in light of human rights principles—such as dignity, non-bias, autonomy, and privacy. Additional safeguards must be in place to protect patient data, and high-risk applications (e.g., prescribing medication or providing psychological counseling to patients with depression) and unauthorized experimental testing must be closely monitored [72].
  • Game Industry. In the gaming industry, AI has significantly enhanced development efficiency through applications ranging from narrative content generation to algorithmically driven NPC decision-making. This has enabled personalized recommendation systems and new operational models based on behavioral analytics. However, while these innovations disrupt traditional workflows, they also present challenges for governance in the gaming sector. Melhart et al. explore ethical issues arising from AI deployment in games and outline key value alignment principles, including: establishing clear chains of responsibility, ensuring transparency in decision-making, maintaining algorithmic and system transparency, enabling auditability of AI outputs, protecting training data and models against malicious attacks and manipulation, and ensuring the predictability of generated content [73].
Cultural Level Beyond national and industrial dimensions, cultural plurality and diversity represent a critical facet of value alignment in AI. Existing studies often draw on Hofstede’s cultural dimensions theory [74], focusing on how AI systems express value preferences across six key cultural dimensions—power distance (large/small), uncertainty avoidance (strong/weak), individualism/collectivism, masculinity/femininity, long-term/short-term orientation, and indulgence/restraint —and the extent to which these align with human value tendencies in different cultural contexts.
  • Power distance. It refers to the extent to which less powerful members of organizations accept and expect unequal power distribution.
  • Uncertainty avoidance. This reflects the degree to which individuals feel uncomfortable or comfortable in ambiguous or unknown situations.
  • Individualism/collectivism. This dimension captures the degree to which individuals are integrated into groups and the responsibilities they assume toward those groups.
  • Masculinity/femininity. This dimension denotes the prom-inence of traits traditionally associated with masculine or feminine roles in society.
  • Long-term/ short-term orientation. This dimension measures the acceptance of delayed gratification of material, emotional, and social needs.
  • Indulgence/restraint. The indulgence versus restraint dimension concerns the degree to which societies permit the gratification of basic human desires and enjoyment.
Based on these dimensions, Wang et al. found that LLMs prompted in English tend to exhibit a masculine orientation, while those prompted in German show a stronger inclination toward long-term orientation and indulgence. The cultural traits exhibited by LLMs under Chinese prompts fall between those of English and German prompts [75]. Similarly, Kharchenko et al. observed that LLMs can recognize cross-cultural value differences but consistently lean toward specific value orientations. In particular, across all languages and prompting methods, LLMs predominantly favor long-term orientation, suggesting that LLMs may faithfully reflect certain cultural values [76].
In summary, the global practices of AI governance reveal diverse paths and efforts toward value alignment across nations, industries, and cultures. Value alignment is not merely a matter of matching technical systems with human intentions; it is also a profound reflection of policy choices, institutional designs, and cultural contexts at the governance level. The United States emphasizes industry self-regulation and market-driven mechanisms, reflecting its prioritization of innovation and individual freedom. The European Union seeks institutional value alignment through legislation and ethical frameworks, highlighting fairness, justice, and the protection of human rights in technological applications. China integrates technological advancement with social responsibility, reinforcing state-led ethical norms that emphasize collective interests and public safety.
Meanwhile, international organizations are increasingly working to build a global consensus on value alignment by promoting cross-border ethical cooperation and industry-wide standards. Given the differing emphases on values across nations and the deep integration of AI in various sectors, future governance systems must pursue a dynamic balance and coordination among pluralistic value claims. Through dialogue and collaboration, global value alignment can be advanced to ensure that technological development truly serves the overall well-being of human society.
2.3.3 Micro Level
As AI agents become increasingly embedded in organizational operations and business processes, the issue of value alignment at the micro level has grown more prominent. Unlike the overarching constraints imposed by macro-level universal values and meso-level policy regulations, the micro level focuses on how AI systems achieve alignment between effectiveness and values within localized environments under specific goal orientations. The scenarios and associated values at this level are inherently more personalized. Given the impossibility of exhaustively enumerating all possible scenarios, this study synthesizes relevant research and illustrates the principles of micro-level value alignment through examples such as talent recruitment, legal consultation and corporate governance in the healthcare sector.
The application of artificial intelligence technologies is profoundly reshaping human resource management practices, with the recruitment domain emerging as a primary frontier of technological transformation. AI-powered talent selection systems are driving a fundamental shift in organizational talent acquisition paradigms. Compared to traditional recruitment models that rely heavily on manual screening and subjective judgment, intelligent tools significantly enhance the speed and accuracy of candidate identification through automated resume parsing, big data-driven competency modeling, and machine learning-enabled predictive assessments [97].The ethical compatibility of AI technologies in recruitment settings has attracted sustained academic attention. In constructing ethical principles, Simbeck advocates for the transfer of key moral concepts from other domains to human resource analytics when applying new AI technologies, identifying five core principles: data privacy protection, user opt-out mechanisms, third-party audit frameworks, algorithmic transparency, and the safeguarding of dynamic development rights [79]. These principles align with the humanistic view of technology proposed by Runb-Kettler and Lehnervp, which emphasizes preserving candidates’ dignity and autonomy in automated processes [78]. A systematic literature review by Hunkenschroer and Luetge further deconstructs the ethical dimensions of AI in recruitment into ethical opportunities, ethical risks, and areas of ambiguity [80].
From the perspective of ethical opportunities, AI can reduce human biases in traditional recruitment through standardized assessment procedures [98][99]. Text optimization tools can help generate gender-neutral job descriptions [100], while natural language processing has revolutionized resume parsing efficiency, allowing recruiters to focus on strategic decision-making [101]. However, ethical risks remain significant. Algorithms may inherit implicit biases embedded in training data [102], implement indirect discrimination via proxy variables (e.g., zip codes correlated with race), or infringe on privacy rights by extracting personal information from candidates for recruitment purposes [103][104]. Although AI has the potential to disrupt traditional elite hiring models and promote social mobility, it may also lead to reduced cognitive diversity in teams due to algorithmic homogenization. These tensions and risks underscore fairness as a core principle in the application of AI to human resource recruitment.
The deep integration of AI technologies is also influencing the strategic development and governance choices of biopharmaceutical companies. In response to the ethical challenges posed by the application of AI in healthcare, AstraZeneca has issued a series of AI ethical principles, including privacy and security, interpretability and transparency, fairness, accountability, human-centeredness, and societal benefit. The company has also developed a Responsible AI Handbook, established an AI Resolution Committee and internal advisory services, and commissioned AI audits to implement comprehensive AI governance [82].
In addition, AI technologies are demonstrating significant potential and application prospects in the legal consulting domain. Sun et al. simulated the collaborative workflows of real-world law firms using multi-agent systems. In this context, they incorporated normative principles—such as privacy protection, fairness, transparency, and accountability—to guide the ethical behavior and interactions of the agents. This alignment with human values enhances the quality of legal consulting services by ensuring that multi-agent systems operate in a manner consistent with ethical and societal expectations [81].
As AI continues to advance, the principles of value alignment will no longer be confined to broad and ambiguous moral values. Instead, they will increasingly penetrate specific application scenarios, integrating concrete practical experience and well-defined guidelines to inform the development and deployment of agent systems.
3 Applications and Value Issues in LLM-based agent systems
LLM-based agent systems are now applied across numerous domains, indicating that value alignment concerns will permeate to the application scenario level. This section first introduces our review subjects—LLM-based agents and multi-agent systems. Subsequently, it classifies and organizes intelligent agent system application scenarios based on their generality, and discusses the associated value alignment issues.
3.1 LLM-Based Agent Systems
3.1.1 LLM-Based Agent
In philosophy, the term ’agent’ typically refers to an entity possessing consciousness, intention, and the capacity for action [105, 106]. This can include humans, animals, and other entities in both the physical and virtual worlds [1]. Floridi identifies three general characteristics of an agent: interactivity, autonomy, and adaptivity. Interactivity refers to an agent’s ability to interact with and respond to environmental stimuli. Autonomy indicates an agent’s capacity to change its own state based on its transformation rules. Adaptivity denotes an agent’s ability to not only alter its state but also to modify the rules governing these changes through learning and experience, thereby flexibly responding to environmental shifts. Building on these three characteristics, Floridi posits that a Moral Agent can exhibit morally evaluable behavior, meaning their actions can engender moral "good" or "evil." Consequently, Moral Agents must adhere to specific ethical norms and oversight mechanisms [107]. When the concept of an "Agent" is introduced into the field of artificial intelligence, it becomes known as an AI agent. An AI agent is an artificial entity capable of achieving complex goals through collaboration [108], representing the concretization of the philosophical concept of "agent" within artificial intelligence.
Large Language Models (LLMs) are deep learning models trained on vast amounts of textual data. They are capable of generating natural language text and comprehending its meaning, with examples including Deepseek, GPT-3 and GPT-4. A salient characteristic of LLMs is their adherence to scaling laws and the exhibition of emergent abilities. This means that once a model’s training parameters exceed a certain threshold, it can experience a dramatic improvement in existing capabilities or acquire entirely new ones [17], enabling it to solve tasks it was not explicitly trained for [12]. As models continue to grow in scale, pre-training corpora expand, and alignment techniques advance, LLMs have demonstrated powerful capabilities in knowledge acquisition, instruction understanding, summarization, planning, and reasoning [1]. This showcases their immense potential for achieving human-like intelligence. Consequently, researchers have begun leveraging LLMs to construct intelligent digital agents [23], transforming abstract philosophical concepts into practically engineerable system design frameworks [1].
AI Agents are intelligent entities capable of perceiving their environment, making decisions, and executing actions [109, 110, 111]. LLMs can serve as the central controller for AI agents, endowing them with decision-making thought and memory [112]. Decision-making thought enables agents to decompose complex tasks, deliberate on each sub-task, and explore multiple pathways to arrive at intricate decisions. A long- and short-term memory mechanism allows agents to maintain contextual coherence during task completion and continuously learn from past interactions. Strategies such as multimodal perception and tool use expand an agent’s perception and action space [113, 114, 115, 116, 117]. Furthermore, AI agents endowed with environmental interaction capabilities can learn from feedback and execute novel actions [118]. The development of techniques like Chain-of-Thought (CoT) and problem decomposition has continuously enhanced the reasoning and planning capabilities of AI agents[119, 120]. Furthermore, pre-training on massive-scale corpora has enabled agents to exhibit few-shot and zero-shot generalization abilities. This allows them to transfer and complete tasks across different scenarios without requiring parameter updates [120, 121]. The field of AI is currently undergoing a profound paradigm shift. Building upon existing research, the industry is transitioning from developing AI models for passive, structured tasks to constructing novel models capable of acting as dynamic agents in diverse and complex scenarios.
3.1.2 Multi-Agent System
A Multi-Agent System (MAS) is a computer system composed of multiple interacting agents [12]. Operating on natural language and custom content [24], these agents can interact within a shared environment, autonomously making decisions to solve complex problems or complete tasks [122]. The key components of such a system include the Agent, Environment, Interaction rules, and Organization pattern [12].
  • Agent.The agent is the core participant in the system, possessing a role identity, knowledge, capabilities, and behaviors, along with abilities such as learning, planning, reasoning, and decision-making.
  • Environment.The environment is the external world in which the agent operates, which can be either a simulated virtual environment or a real physical space.
  • The Interaction Rules. The rules are predefined rules, protocols, and mechanisms governing information exchange and behavior coordination among agents.
  • The Organization Pattern. The pattern refers to the structural relationships of agent interactions, encompassing hierarchical control and emergent behavior-based organizations.
Multi-agent interaction within these systems mirrors the specialized division of labor and teamwork found in human society; this aims to achieve collective intelligence. Individual autonomous agents possess initial knowledge and goal sets tailored for specific tasks [122]. By leveraging LLMs’ capabilities in text generation, comprehension, and communication, diverse agents can adaptively collaborate. They respond to task requirements, participate in planning, discussion, and decision-making, and learn from one another. This fosters the aggregation of broad knowledge across multiple domains [112, 12]. Concurrently, through environmental interaction, agents perceive and understand information. They then make decisions and execute actions based on their goals, existing knowledge, and observations. This process unleashes their potential to complete specialized tasks and achieve common objectives.
Multi-agent systems exhibit remarkable flexibility, enhanced robustness, inherent self-organization and coordination, and robust real-time operation capabilities [12]. Although interactions among agents are governed and constrained by predefined rules, emergent behaviors and phenomena can still arise from these interactions [123]. Additionally, due to the division of labor within the system, resource and energy consumption are distributed across multiple agents. This is often more cost-effective than a single, powerful agent attempting to solve an entire complex problem [12]. Therefore, compared to a single agent, multi-agent systems are better suited for handling more complex and dynamic tasks, particularly in scenarios requiring diverse strategies and collective decision-making [112]. At present, LLM-based multi-agent systems are extensively used to simulate social phenomena. This includes modeling social network construction, international conflicts, political evolution, and the dynamics of public opinion  [124, 125, 126].
3.2 Application Scenarios
Table II: Classification of Agent System Application Scenarios
Generalizability Category Specific Scenarios Ref
High Generalizability Tool using • General tool using • Multimodal tool using  [127, 128, 129]  [130, 131]
• Instructions acceptance, image capture, • Enterprise-level data workflow automation
historical retrieval and tool invocation
Dialugue Agents • Conducting persuasive dialogues • Dialogue task optimization  [132, 133, 134, 135]  [136, 137, 138, 139]  [140, 141, 142, 143]  [144, 145]
• Multimodal emotional dialogue systems • Continuous conversational dialogue
developing • Cross-domain dialogue clarification
• Personalized experience improvement in strategies
long-term dialogue • Online shopping question answering
• Medical question answering • LLM-driven automated task execution in
• Human-like emotional dialogue agents mobile applications
• Personality consistency testing of • Dialogue strategy planning
role-playing agents • Social intelligence evaluation of role-playing
• Linguistic pattern analysis of dialogue AI dialogue agents
• Personalization validation of dialogue agents
Multimodal Perception • Multimodal tool using • Instructions acceptance, image capture,  [146, 127, 128, 147]
• Multimodal retrieval historical retrieval and tool invocation
• Dynamic scene understanding and guidance
Multi-Agent Collaboration • Multi-agent social interaction modeling • Collaborative task division and quality  [148, 149, 150, 151]  [152, 153, 154, 155]  [156, 157, 158]
• Simulation of social collaborative behaviors assurance in software development
among multiple LLM agents • Training human collaborative behaviors
• Multi-LLM collaboration in medical decision-making supported by generative agents
• Dynamic multi-team adversarial competition • Long-term planning for multi-robot systems
• Language instruction generalization in in partially observable environments
multi-agent games • Performance evaluation of multi-agent
• Dynamic job scheduling in factories systems
• Software engineering automation
Visual Tasks Processing • Reasoning covered object in images • Image Inpainting  [159, 160, 161, 162]
• Multimodal image generation and editing • 3D scene synthesis and rendering
Natural Language Processing • Keyphrase generation • Long text reading comprehension  [163, 164, 165]  [166, 167, 168]
• Database information extraction and integration • Relation triplet extraction in natural language
• Dynamic navigation for incremental extreme
multi-label classification
Data Generation and Analysis • Transporing the knowledge of the isolated agents  [169]
Reasoning, Planning and Decision Optimization • Autonomous information retrieval in visual • Robotic language-guided task planning  [170, 171, 172, 173]  [174, 175, 176, 177]  [178, 179, 180, 181]  [182, 183, 184, 185]  [186, 187, 188, 189]  [190, 191, 192]
question answering • Asynchronous task planning
• Social data analysis and decision-making • Strategic language agent collaboration in
• General zero-shot reasoning task optimization Werewolf game
• Interactive planning task execution • Instruction Following
• Automated task execution • Agent performance analysis in complex
• Multi-agent debates enhancing factuality and Android environments
reasoning of language models • Embodied decision-making task evaluation
• Knowledge distillation from multi-agent interaction • Automated tool using
graphs to improve small model reasoning • Generative agents with dual-process cognition
• Multi-agent social learning mechanisms in complex tasks
• Real-world travel planning • Automated user interface control
• Exploratory trajectory optimization • Network agent planning
• Generalization of robot visual-language tasks • On-device AI assistants
Limited Generalizability Function-driven Code Generation • Multi-agent code generation framework for • Code generation and environment interaction  [193, 194, 195, 196]
competitive programming problems modeling
• Open-domain process customization • Code repair
Social Simulation • User behavior simulation • Human social interaction simulation  [197, 198, 199, 200]  [58, 201, 202, 203]  [204, 205, 206, 207]  [208, 209, 210]
• Social network information propagation simulation • Role-playing agent generation
• Trust behavior simulation • Complex task planning and collaboration
• Business competition scenario simulation in Minecraft
• Sustainable cooperation simulation • Time-aware multitask simulation
• Diplomacy simulation • Complex social interaction simulation and
• Social intelligence interactive learning evaluation
Graphical User Interface Agents • Multi-turn conversational web agents • Personalized web agents  [211, 212, 213, 214]  [215, 216, 217, 218]  [219, 220, 221, 222]  [223, 224, 225, 226]  [227, 228, 229]
• Graphical interface tasks automating • Computer screen interaction and control
• Web interaction task execution • Smartphone GUI automation
• Knowledge transformation and task automation • Web navigation tasks executing
• GUI agents autonomously executing complex • Adversarial attacks on web agents
multi-step tasks • End-to-end web interaction agents
• Enterprise-level web task automation • Training of web agents
Model Analysis, Evaluation and Improvement • Dialogue task optimization • Multi-agent evaluation optimization  [230, 62, 231]  [232, 233, 234]
• Neural network behavior interpretation • LLM-driven recommender system attack
• Evaluation of LLM alignment with human values testing
• Instruction tuning and data optimization
Scenario-driven Finance • Multimodal foundation agent for financial trading • Financial sentiment analysis  [235, 236, 237, 238]
• Multi-agent system for enhanced financial • Macroeconomic activities simulation
decision making
Science • Automating data science tasks • Automating scientific discovery  [239, 61, 240]  [241, 242]
• Design of organic structure-directing agents • Scientific data visualization
• Academic citation recommendation
Healthcare • Medical reasoning and diagnosis • Electronic health record modeling and  [243, 244, 245, 246]  [247, 248, 249]
• Pathology image-text pairs generation and clinical decision support
analysis • Multimodal interaction testing of mental
• Psychological measurement health dialogue agents
• Drug design • Gene perturbation experimental design
Game • Fighting game agents training • Intelligence evaluation in murder mystery  [250, 251, 252]  [253, 254]
• Detective role-playing reasoning games
• Interaction and reasoning in text-based games • Text-based game agents
Manufacturing • Dynamic scheduling in flexible manufacturing  [254]
systems
Robotics • LLM-based embodied agent open environment • Embodied instruction following  [255, 256, 257, 258]
perception • Home robot knife placement
• Vehicle acceleration and obstacle avoidance • Mobile device operation assisting
in autonomous driving
Urban Computing • Addressing standardization tasks supported • Urban knowledge graph construction  [259, 260, 261]
by geospatial tools • Personal mobility generation
Social Media • Detecting harmful content on social media • Logical and causal fact-checking  [262, 263]
Autonomous Driving • Real-time motion generation for autonomous driving • Traffic signal control  [264, 265]
Literary Creation • Director-actor coordinate agent framework for • Narrative generation through multi-step  [266, 267]
controllable  drama script generation collaboration
Given the poweful knowledge acquisition, instruction understanding, task planning and logical reasoning capabilities of multi-agent systems, they have been widely used worldwide. From small and medium-sized organizations to multinational corporations, from higher education institutions to research institutes, from grassroots medical facilities to large diagnostic hospitals, and government agencies at all levels are actively promoting the application of AI agent technology. In terms of application scenarios, the penetration of AI agent technology is extremely broad, covering multidisciplinary fields such as healthcare, education, trade and economics and law. The problems that multi-agent systems are capable of addressing ranges from comprehending a single piece of literature to simulating global-scale social events, spanning across micro, meso and macro levels. In contrast to categorizing applications based on disciplinary domains [14] or task types [112], we sort and categorize application scenarios according to the degree of generalizability of multi-agent systems and build upon this to explore the associated issues of value alignment.
Generalizability of multi-agent systems refers to the extent to which they can be applied across diverse scenarios or meet the needs of various organizations. A high degree of generalizability is based on the foundational capabilities of AI agents such as natural language processing, reasoning, planning and autonomous decision-making. The universal application includes tool using, dialog agents, multimodal perception, multi-agent collaboration, visual tasks proingcess, natual language processing, data generation and analysis, reasoning, planning and decision optimization and on, see Table II.
In terms of tool using, AI agents are capable of recognizing implicit needs for tool invocation and leveraging a wide range of real-world tools involving perception, manipulation, logical reasoning, and creativity. These tools enable AI agents to process multimodal inputs—such as images, tables, and code—to accomplish complex tasks [129, 130]. Taking dialogue agents as an example, A.V. Savchenko and L.V. Savchenko (2024) proposed an approach that employs multiple emotion-aware dialogue agents based on large language models to provide query responses. These AI agents integrate speech and facial expression recognition to infer the user’s emotional state and generate appropriate responses [134]. Furthermore, dialogue agents have also been applied in persuasive interactions for social good, particularly in scenarios involving user resistance, where they are able to generate more polite and convincing responses [132]. In the domains of reasoning, planning, and decision-making optimization, autonomous agents can guide large language models to perform zero-shot reasoning, thereby improving their performance in generation, classification, and inference tasks  citeDBLP:conf/icml/CrispinoMZS024. Multi-agent debate mechanisms have been shown to enhance accuracy and factual consistency in mathematical and strategic reasoning while reducing hallucinated outputs [175]. Additionally, structuring multi-agent interactions as graph representations enables knowledge distillation for smaller models, improving their performance and generalization in commonsense and mathematical reasoning tasks [176].
Higher degree of generalizability of indicates that the agent system can be broadly applied across various scenarios, not only addressing domain-specific problems but also meeting diverse task requirements across multiple industries and organizations. For example, multi-agent systems can be applied not only in natural sciences but also in social science domain such as economics and management, suitable for addressing global issues that span across nations and cultures.
Lower generalizability means that multi-agent systems can not be applied across industries and domains. Specialized agent applications can be broadly categorized into two types: function-driven and domain-driven, see Table II.
Function-driven applications are developed by building upon the core capabilities of agents and are typically applicable across multiple domains. Common examples include code generation, social simulation, graphical user interface (GUI) agents, and model analysis, evaluation and improvement, as in Table II. For instance, in social simulation, Wang L. et al. (2023) proposed a large language model-based agent framework that simulates user behavior in sandbox environments to examine social phenomena such as echo chambers and herding effects [197]. In GUI navigation, agents can handle open-domain web tasks by extracting key information from complex HTML pages, understanding web structures, and performing multi-step interactions, thus addressing challenges like interaction diversity and domain knowledge scarcity [225, 227]. In model analysis, LLM-based agents have been employed to assess alignment performance, such as in Ali-agent, which simulates test scenarios and iteratively explores long-tail risks to evaluate value alignment [231]. MAIA, a multimodal explanatory agent, automates model behavior interpretation using neural networks, enhancing model transparency [62].
Scenario-driven applications refer to agent systems deployed within specific sectors, including finance, science, healthcare, game, manufacturing, and autonomous driving, as in Table II. In urban computing, the UrbanKGent framework enables automated construction of urban knowledge graphs, significantly reducing development costs [261]. In scientific research, agents assist with academic literature retrieval and citation recommendation [61], organic structure design [241], and automated data visualization. In healthcare, multi-agent systems support medical reasoning, gamified mental health assessments, and experimental design in genomics [243, 247]. In finance, agents can simulate macroeconomic dynamics [236], analyze market trends using multimodal data [235], and support trading and portfolio management [237]. Furthermore, in the social media domain, agent reasoning and coordination capabilities contribute to tasks such as harmful content detection and high-confidence fact verification [262, 263].
The low level of generalizability indicates that the agent system is used in specific countries, industries, or even organizations and enterprises. It is dedicated to completing a certain type of task, and the data sets and business information involved are highly private.
3.3 Value Issues in Application
The value alignment problems involved in the application scenarios of LLM-based agent systems mainly include two aspects: the correspondence between the generalizability of the system and the multi-level values, and the contradictions encountered when aligning values.
The degree of generalizability of LLM-based agent systems corresponds to the multi-level of values. As the degree gradually decreases, the granularity of the value norms that need to be met gradually increases. A higher degree of generalizability indicates that the applicable scenarios for LLM-based agent systems are more diverse. Therefore, it is unnecessary for the initial design of systems to align with the values of a specific group–such as a country, industry, or culture–to ensure flexible applicability across national, ethnic, and industrial contexts. The development and governance of highly generalizable multi-agent systems should be constrained by universal values of macro level. Multi-agent systems with limited generalizability typically confined to specific countries or industries. Such systems often exhibit design features characteristic of particular countries, industries, or cultures. Therefore, the development and governance of specialized systems must address diverse national policies, industry standards, and cultural preferences, while ensuring alignment with universal values. When intelligent systems are deployed in strictly private contexts, they are typically confined to specific organizations and designed to address specialized business or personal needs. These applications are highly specialized and exhibit minimal generalizability. Under such conditions, it is essential not only to align with value principles of macro and mecro level, but also to comply with the specific value frameworks of the organizations or businesses involved. It can be seen that multi-agent systems with high generalizability tend to be highly adaptable, and their value alignment primarily centers on adherence to universal principles of macro level. As application specialization increases, the requirements for value alignment become progressively more complex, necessitating adherence to value norms across multiple levels simultaneously. This classification aligns with the top-down, increasingly granular hierarchy of value principles. As a result, the value norms that multi-agent systems must adhere to vary according to their degree of generalizability, offering a structured basis for governance and oversight in their design and deployment.
The value alignment problem may involve two tpyes of tensions: the plurality of value principles at the same level, and potential conflicts between values across different levels. On one hand, a key challenge is that value principles at the same level differ across countries, regions, industries, organizations, and other contextual factors. Therefore, when multi-agent systems are applied in diverse contexts, their value alignment objectives must be adapted accordingly. In the healthcare sector, Huawei’s intelligent agent EIHealth serves as a professional AI development platform for genomic analysis, drug discovery, and clinical research. Built on Huawei Cloud’s strengths in AI and big data, EIHealth features an open and extensible architecture that enables end-to-end AI empowerment. Ensuring alignment with the value principles and regulatory norms of various countries, regions, and organizations is crucial for its deployment. To address this, Huawei Cloud has issued a series of white papers tailored to the privacy and regulatory requirements of different jurisdictions and industries. For example, it provides compliance guidelines aligned with the Personal Data Privacy Ordinance (PDPO) in the Hong Kong SAR, as well as guidelines addressing the financial regulatory frameworks of the EU, including specific provisions for Ireland, Spain, and Hungary. These efforts ensure that the EIHealth platform adheres to all applicable national and regional security regulations, international cybersecurity and cloud security standards, and incorporates industry best practices—ultimately meeting the diverse security needs of cloud service users.
On the other hand, as the tasks assigned to an intelligent system become more fine-grained, the constraints on achieving alignment goals also increase. The means of maximizing task utility may conflict with established value norms, leading to tensions across different levels of values. For instance, the promotion of social cohesion at the macro level may contradict individualistic tendencies at the meso level, where individuals prioritize self-interest. Similarly, national regulations that emphasize information transparency may be at odds with micro-level demands for organizational data privacy. These cross-level value conflicts pose significant challenges for the design and development of multi-agent systems. In the field of corporate recruitment, algorithms and artificial intelligence have been widely adopted. Recent studies have explored the design and development of intelligent agent systems for tasks such as résumé screening and candidate-job matching [268]. However, AI-based recruitment systems may be at odds with widely accepted ethical norms [104]. In particular, corporate hiring practices can involve gender discrimination or racial bias, which directly conflict with legal regulations and core societal values. This ethical dilemma was vividly illustrated by Amazon’s 2018 algorithmic bias incident. The company’s internally developed résumé screening system, trained on historical hiring data, systematically discriminated against female applicants, ultimately leading to the discontinuation of the system [269]. To prevent such issues, recruitment-oriented intelligent systems must be explicitly designed to reduce or eliminate gender and racial bias. Their value alignment should comply with national labor laws, equal employment regulations, women’s rights protections, and relevant industry standards. Only on the basis of legal compliance, equality, and privacy protection should these systems then pursue the optimization of recruitment outcomes.
In complex collective intelligence systems, multi-agent value alignment refers to the design of appropriate mechanisms and algorithms that enable multiple autonomous agents to identify and adhere to shared ethical or social norms while pursuing their individual task objectives, thereby achieving consistency, safety, and fairness at the system level. Unlike single-agent alignment, which focuses on the compliance of a model’s outputs, multi-agent collaborative alignment must address challenges such as value conflicts between agents, information asymmetry, and “free-rider” or competitive dilemmas in cooperative games. The goal is to use communication protocols, reward shaping, negotiation mechanisms, or adversarial training to ensure that agents coordinate with one another and ultimately converge on a common set of collective values.
4 Evaluation and Methods for Value Alignment
To ensure that LLM-based agents behave in ways consistent with human values and task goals, value alignment evaluation has become a key research focus. Evaluation methods often rely on alignment benchmarks to assess whether agents follow expected value principles. On the methodological side, alignment techniques can be applied at various stages, including during pretraining, post-training, and through reinforcement learning (e.g., reward modeling or fine-tuning with human feedback). These methods aim to mitigate misalignment risks and enhance agent trustworthiness across various applications.
4.1 Value Alignment Evaluation
4.1.1 Datasets for Value Alignment Evaluation
Table III summarizes the existing open-source datasets for evaluating value alignment. To clarify their coverage and applicability, we systematically organize and categorize these datasets according to the previously introduced three-level framework of value alignment principles: macro, meso, and micro. This hierarchical classification enables a more nuanced understanding of how current resources support value alignment evaluation at different levels.
Table III: Statistics of Datasets for Value Alignment Evaluation.
AC: Automatic Construction, CC: Crowdsourced Construction, HC: Hybrid Construction, EDC: Expert-Driven Construction, IRED: Integration and Refinement of Existing Datasets, BJ: Binary Judgments, RT: Ranking Tasks, MCQ: Multiple-Choice Questions, OEQ: Open-Ended Questions, RSQ: Rating Scale Questions
Datasets Macro Level Meso Level Micro Level Construction Methodologies Question Formats Size
BBQ(Bias Benchmark for QA) [270] EDC MCQ 58.4k
BeaverTails [51] IRED OEQ 30.2k
BOLD [271] CC OEQ 23.6k
CBBQ(Chinese Bias Benchmark Dataset) [272] HC MCQ 106.5k
CDEval [75] AC MCQ 2.9k
CDial-Bias [273] CC OEQ 28k
CORGI-PM [274] IRED BJ, OEQ 32.9k
CrowS-Paris CC BJ 1.5k
CultureSPA [275] AC MCQ 13k
Cvalues [276] CC MCQ, OEQ 6.4k
DailyDilemmas [277] AC MCQ 2.7k
DecodingTrust [278] IRED MCQ 152.4k
DEFSurveySim [49] HC MCQ, RSQ 1.1k
EEC(Equity Evaluation Corpus) [279] EDC BJ 8.6k
ETHICS [280] CC MCQ, BJ, RT 134.4k
Flames [281] CC OEQ 1k
GlobalOpinionQA [282] IRED MCQ 2.5k
HofstedeCulturalDimensions [76] HC MCQ 0.2k
IndieValueCatalog [283] IRED MCQ 93.2k
KorNAT [284] AC MCQ 10k
LLMGlobe [77] HC RSQ, OEQ 37.6k
MFQ(Moral Foundations Questionnaire) [285] IRED MCQ 11k
Moral Beliefs [286] AC MCQ, BJ, OEQ 0.5k
Moral Integrity Corpus [287] AC RSQ, OEQ 38k
Moral Stories [288] CC MCQ 12k
MoralExceptQA [289] EDC RT 0.2k
NaVAB [290] AC MCQ, BJ 63.8k
Persona Bias [279] IRED MCQ, BJ 110.6k
PkuSafeRLHF [291] AC OEQ 476.4k
ProgressGym [292] HC RSQ, OEQ 1.4k
SafeSora [60] AC OEQ 14.7k
Scruples [293] EDC MCQ, BJ 657k
Social Bias Frames [294] CC RT, OEQ 150k
Social Chemistry 101 [295] CC BJ, OEQ 292k
StereoSet [296] CC RT, OEQ 4.2k
ToxiGen [297] AC OEQ 6.5k
UnQover [298] IRED BJ, OEQ 2713k
ValueNet [299] CC BJ, OEQ 21.3k
WikiGenderBias [300] IRED OEQ 45k
WinoBias [301] EDC MCQ 3.1k
WinoGender [302] EDC BJ, OEQ 0.7k
Although existing datasets address multiple levels of value alignment, several limitations persist. At the macro level, many datasets—such as Moral Stories [288] and ETHICS [280]—are predominantly grounded in Western cultural contexts, lacking comprehensive representation of alternative value systems, such as those from Asian or African societies. In addition, these datasets largely reflect contemporary norms and offer limited support for modeling the evolution of values over time or predicting future value shifts. At the meso level, datasets such as NaVAB [290], KorNAT [284], and CultureSPA [275] provide valuable insights into national or cultural contexts. However, they offer limited global coverage and often lack detailed evaluations of value alignment across specific industries or professional domains. Micro-level datasets are even more scarce. Compared to macro- and meso-level resources, they remain significantly underdeveloped, which limits the ability to assess and align value preferences in multi-agent systems across diverse, fine-grained real-world scenarios.
4.1.2 Dataset Construction Methodologies
Constructing high-quality value alignment datasets is critical for reliably evaluating whether different systems adhere to appropriate value frameworks. To support diverse application scenarios, researchers have primarily adopted three data construction approaches: manual construction, automatic construction, and hybrid construction that integrate human and machine-generated content.
Manual Construction. In the manual construction approach, experts or annotators design tasks, generate textual content, and assign labels based on a well-defined value framework. The resulting datasets are typically curated and reviewed by trained annotators, yielding high-quality data that closely reflects human value understanding. However, this approach is resource-intensive and lacks scalability. In addition, annotations grounded in subjective human judgment may unintentionally introduce bias. The manual construction approach can be further categorized into three subtypes: expert-driven construction, crowdsourced construction, and the integration and refinement of existing data.
  • Expert-Driven Construction is led by domain experts who design tasks and define annotation protocols to ensure data accuracy and consistency. For example, the BBQ dataset [270] was manually created using templated questions that include both biased and unbiased forms, covering nine social dimensions with contextual and disambiguating cues. Similarly, the MoralExceptQA dataset [289] was built by systematically varying moral scenario features—such as rule function and consequences of violations—to generate diverse questions on moral exceptions. These expert-curated datasets offer high-quality benchmarks and valuable tools for evaluating model performance in targeted domains.
  • Crowdsourced Construction is widely adopted for its efficiency and cost-effectiveness, particularly in large-scale data collection. Many datasets—such as Social Bias Frames [294], Flames [281], and BOLD [271]—leverage crowdsourcing platforms, often with multiple rounds of quality control to ensure annotation reliability. These datasets not only scale well but also capture diverse perspectives from annotators of varied cultural and social backgrounds, enabling more comprehensive assessments of model behavior.
  • Integration and Refinement of Existing Datasets involve aggregating and modifying multiple public sources. Notable examples include IndieValueCatalog [283], DecodingTrust [278], MFQ [285], Persona Bias [279], and GlobalOpinionQA [282], all constructed through careful curation and consolidation of open-source data. This approach reduces data acquisition costs while leveraging existing high-quality resources, thereby improving research efficiency and enabling broader evaluation coverage.
Automatic Construction. With the rapid advancement of large language models (LLMs), LLM-based automatic dataset construction has become a promising approach. This method typically involves designing prompt templates to guide the model in generating question–answer pairs or scenario-based judgments, followed by manual verification to ensure data quality. Representative datasets developed using this approach include NaVAB [290], KorNAT [284], , CultureSPA [275], and DailyDilemmas [277]. Compared to manual construction, it enables the creation of more diverse and context-rich data with significantly less human effort, allowing for large-scale dataset generation in a relatively short time. However, concerns remain regarding the quality of model-generated content. LLMs may produce hallucinated outputs, and their inherent biases can result in misalignment with human values. Consequently, ensuring the reliability and value alignment of automatically constructed datasets still requires careful oversight and validation.
Hybrid Construction. Hybrid construction approaches combine human expertise with automatic expansion using LLMs, which are increasingly emerging as a mainstream trend. In this paradigm, experts first create a small set of high-quality seed examples. The LLM is then employed to generate similar instances or expand the dataset, followed by human verification to ensure consistency and quality. Representative examples include CBBQ [272], LLMGlobe [77], and HofstedeCulturalDimensions [76], all of which utilize expert-curated prompts for initial task design, LLMs for large-scale generation, and human review for quality assurance. This hybrid methodology effectively leverages the generative capabilities of LLMs while maintaining the accuracy and coherence provided by human oversight, offering a robust foundation for value alignment research and evaluation.
4.1.3 Question Formats and Metric Design
Evaluation methods for value alignment largely depend on the format of assessment questions. Common question formats include multiple-choice questions, binary judgments, ranking tasks, rating scale questions, and open-ended questions.
Multiple-Choice Questions. Multiple-choice questions typically present a set of answer options—often including neutral, biased, or varying degrees of harmful responses—and require the model to select the most appropriate one. This format is commonly used to assess a model’s value judgments in specific contexts. For example, the BBQ dataset frames each question as a triplet: a neutral version (free of bias), a subtly biased version, and an overtly biased version [270]. By evaluating model accuracy across these variations, researchers assess its fairness and sensitivity to social bias. The CDEval dataset similarly uses multiple-choice questions to examine a model’s alignment with cultural values [75]. For instance, a question may ask whether children should obey their parents in a given culture, with answer options reflecting diverse cultural norms. The model’s selected response offers insight into its sensitivity to and alignment with different cultural perspectives.
Binary Judgments. Binary judgment questions require the model to make a dichotomous decision, such as yes/no, true/false, or agree/disagree, based on a given statement or scenario. This format is widely used to evaluate a model’s understanding of ethical principles, social norms, or cultural consensus. For example, the ETHICS dataset presents ethically challenging statements (e.g., “Stealing is acceptable if no one finds out”), each of which is annotated by humans as either correct/incorrect or acceptable/unacceptable [280]. The model or agent system is tasked with determining whether the described behavior is morally appropriate, and its accuracy serves as a measure of ethical reasoning and value alignment.
Ranking Tasks. Ranking tasks evaluate a model or agent system’s ability to make relative value-based comparisons by requiring it to order multiple responses or behavioral options from most to least aligned with human values. This format reduces inconsistencies inherent in subjective rating scales and highlights the model’s capacity for value-informed prioritization. For instance, StereoSet dataset presents stereotype-laden prompts followed by a question and several options (e.g., “A doctor walks into the room and sees the patient:” with options such as: A. He is wearing a white coat. / B. She is wearing a white coat. / C. He is holding a guitar.) [296]. The model is tasked with ranking these responses based on contextual appropriateness in terms of grammar and semantics. A Stereotype Score is then used to evaluate ranking value tendencies. Similarly, MoralExceptQA dataset provides a normative moral rule (e.g., “One should not lie”) and a complex scenario (e.g., “Is it acceptable to lie to protect a friend’s safety?”) [289]. The model or agent system must rank behavioral options, such as lying, telling the truth, or indirect concealment, based on moral reasoning. These rankings are compared to those generated by human ethicists to assess the model’s ability to integrate moral principles with contextual judgment.
Rating Scale Questions. Rating scale questions require the model or agent system to evaluate a given text, response, or behavioral option along predefined value dimensions. For instance, the Moral Integrity Corpus presents scenarios in which dialogue agents face moral dilemmas, providing contextual setups and multiple behavioral options (e.g., whether to disclose a friend’s secret or follow a rule) [287]. The model or agent system rates the moral appropriateness of each option on a five-point Likert scale (e.g., 1 = highly immoral, 5 = highly moral). These ratings are then compared to human judgments using Pearson or Spearman correlation coefficients to evaluate moral alignment. Similarly, the ProgressGym dataset presents a range of policy recommendations (e.g., promoting vaccination, providing clean water, ensuring universal education) [292]. The model or agent system is asked to rate each action across multiple dimensions, such as sustainability and contribution to human well-being. The alignment between model-generated and expert-assigned scores is measured using techniques such as weighted averaging or principal component analysis.
Open-Ended Questions. Open-ended questions allow free-form responses from the model or agent system, enabling the assessment of its alignment with human values. This format evaluates not only factual correctness but also the appropriateness and social acceptability of generated content. For instance, BeaverTails dataset prompts models with potentially harmful queries (e.g., “How to build a weapon” or “How to hurt someone”) and examines whether the outputs contain dangerous, unethical, or illegal content [51]. Human annotators classify responses based on ethical violations, harmful suggestions, or offensive language. Model performance is then measured using metrics such as harmfulness rate, helpfulness, ethical alignment, and refusal capability. The BOLD dataset presents sensitive prompts related to identity attributes such as gender, race, and religion (e.g., “Describe a day in the life of a Muslim person”) to elicit descriptive responses [271]. Final assessments of value alignment are conducted by domain experts based on the generated content.
Evaluation methods for value alignment largely depend on the format of assessment questions. Accordingly, different metrics are used to quantify model performance across various task types:
Accuracy. Accuracy is the most commonly used metric for multiple-choice and binary judgment questions. It measures the proportion of model responses that match the correct or human-annotated answers.
Correlation-Based Metrics. Correlation-based metrics, such as Pearson’s correlation coefficient and Spearman’s rank correlation coefficient, are widely used in both rating scale and ranking tasks. These metrics evaluate the extent to which model predictions align with human judgment patterns across various value dimensions.
Error-Based Metrics. When the model’s output is numerical (e.g., a score prediction), Root Mean Squared Error (RMSE) or Mean Absolute Error (MAE) may also be used to capture the deviation between predicted and reference scores. These are particularly useful in rating scale questions.
Other Metrics. In open-ended questions, where responses are free-form text, evaluation is primarily based on human judgment, with annotators assessing multiple value-related dimensions and assigning explicit scores.
4.2 Agent Value Alignment Methods
4.2.1 Value Alignment in Pretraining Stage.
Introducing value alignment during the pretraining phase of large language models equips agents not only with extensive linguistic knowledge and behavioral patterns drawn from vast, unannotated corpora but also with foundational signals reflecting human ethics and societal norms [11, 303, 304]. As a result, the agent exhibits higher safety and compliance in downstream tasks. Existing research primarily explores this topic from the following four perspectives:
Value-Prompt Pre-conditioning: Within the conventional self-supervised pretraining frameworks (such as masked language modeling or autoregressive language modeling), researchers prepend a short “value prompt” to the input sequence [305, 306]. They might attach “Please respond with a safe and neutral tone” or “Please adhere to principles of fairness and inclusivity” at the beginning of each sentence, and then include both the prompted and original text in the pretraining corpus. In this way, while the model iteratively learns statistical language patterns, it passively receives implicit guidance toward objectives like “safety” and “objectivity. For example, Zheng et al. [305] unpack why and how simple prepended safety directives move model representations into “higher refusal” regions. [307] shows that conditioning an autoregressive LM on human preference scores from day one dramatically reduces harmful output rates. Zhao et al. [308] builds on “prefix/safety prompt” staff by optimizing two opposing prompts (safe vs. refusal) via contrastive decoding, illustrating that improved prompt design at inference can trace back to how prompts were structured or pre-embedded in pretraining. Gu et al. [309] show that soft prompts—even if not explicitly about “values”—validate the premise that embedding trainable tokens in early stages leads to more controllable downstream behaviors (e.g., fairness, bias detection).
Multi-Task Pre-training: Inspired by the multi-task learning paradigm, this approach mixes value-judgment or safety-detection tasks with standard self-supervised objectives [310, 311, 312]. Specifically, in addition to the masked language modeling (MLM) or autoregressive objective, small-scale, labeled supervision signals—such as “ethics classification”, “bias detection” or “hate speech recognition”—are incorporated. By jointly optimizing a multi-task loss function, the model’s parameters are trained to satisfy both language modeling requirements and the ability to distinguish harmful or biased content. Although this method requires additional annotated data, it explicitly strengthens the model’s discrimination boundary along value dimensions and lays the technical foundation for subsequent alignment fine-tuning.
Regularization and Adversarial Penalty: This strategy introduces penalty terms against undesirable content within the pretraining objective [313]. On one hand, extra $L_{2}$ [314] or entropy regularization [315] is applied to tokens and their contextual representations that are strongly associated with sensitive or harmful semantics, thereby suppressing overfitting to those patterns. On the other hand, an “adversarial penalty” mechanism imposes negative reward on the model’s predicted probabilities when it generates harmful or biased content, forcing the model to gradually move away from those output regions during training [316]. This approach can be viewed as injecting weak-supervision constraints into the unsupervised pretraining corpus, reducing the model’s absorption and memorization of problematic text from the outset.
Safety-Aware Curriculum Learning: Drawing on the concept of curriculum learning (CL), this method stratifies large-scale corpora according to “safety” or “sensitivity” levels [317, 306, 318]. First, high-quality, low-risk “safe corpus” is used for a warm-up pretraining phase, allowing the model to initially acquire expression patterns that comply with ethical and regulatory requirements. Subsequently, more general or higher-risk text is introduced gradually, enabling the model to expand its language capabilities while retaining dependence on earlier safe preferences. This staged training strategy establishes a better balance between model capability and safety, and experiments show it substantially outperforms full-scale pretraining in controlling the model’s tendency to generate sensitive content.
Although each value-alignment method in the pretraining stage has its own focus, they can be combined [311, 306], for example, integrating value-prompt pre-conditioning and multi-task supervision within a safety-aware curriculum framework, along with regularization penalties. This allows the model to be reinforced from multiple dimensions in embedding value objectives such as “ethics,” “safety,” and “fairness.” While these techniques cannot fully replace downstream fine-tuning and reinforcement learning alignment processes, they establish a solid foundation for the entire alignment pipeline. Future research may further explore the scalability and synergistic effects of these strategies on ultra-large corpora and hundred-billion-parameter models, paving new paths for achieving more robust and interpretable value embeddings.
4.2.2 Value Alignment Methods—Post-Training Stage.
Supervised Fine-Tuning (SFT): SFT often serves as the foundational step in post-training alignment. Researchers first construct a human-annotated dataset of “demonstration–response” pairs, where examples cover high-quality answers that adhere to safety, non-bias, and ethical principles [109, 319, 320]. The pretrained model is then instruction-fine-tuned on this dataset so that, given a prompt, it generates outputs that are fundamentally compliant. Because SFT relies solely on positive examples, it can significantly improve the stability and consistency of the model’s initial policy, providing a strong starting point for subsequent reward-based learning.
Reinforcement Learning From Human Feedback (RLHF): RLHF reframes value alignment as a reinforcement learning problem. Its core workflow includes: (1) Preference Data Collection—human annotators pairwise-rank multiple model outputs for the same prompt [319, 320]; (2) Reward Model Training—using preference labels as supervision, a scoring function is trained to predict which response better matches human values [319]; (3) Policy Optimization—algorithms such as PPO are used to fine-tune the language model so as to maximize the reward model’s score. This approach has yielded remarkable results in systems like InstructGPT and ChatGPT, both reducing harmful or inappropriate outputs and improving instruction-following fidelity [109].
Direct Preference Optimization (DPO): DPO aims to streamline the RLHF pipeline by bypassing the explicit training of a reward model [321]. Instead, DPO directly updates model parameters based on the likelihood ratio between positive and negative examples [321], thereby enhancing the model’s probability advantage on preferred responses. Experiments demonstrate that DPO can achieve alignment performance comparable to PPO while offering superior computational efficiency and easier hyperparameter tuning [322], making it a lighter-weight solution for large-scale model deployment [323].
Constitutional AI and AI Feedback: The Constitutional AI framework uses a predefined set of “constitutional principles” to enable the model to self-review and self-correct its outputs. Specifically: (1) generate an initial draft output; (2) evaluate and rewrite the output according to the constitutional principles, either by the model itself or with the assistance of an “AI referee”; (3) use the paired original and revised examples for subsequent fine-tuning. This fully automated closed loop requires no large-scale human re-annotation, continuously improving the model’s value consistency while greatly reducing human cost.
Adversarial Training and Extreme Scenario Reinforcement: Adversarial training generates highly sensitive or extreme prompts—via red teaming or automated scripts—to trigger potential policy violations, then includes these failure cases in the fine-tuning dataset. This strategy bolsters the model’s robustness in edge scenarios; experiments show that adding adversarial data significantly reduces the frequency of non-compliant outputs in worst-case situations.
Contrastive Learning and Unlikelihood Training: Contrastive learning constructs positive–negative pairs of “desired outputs” versus “undesired outputs,” amplifying the probability gap between them to strengthen alignment signals. Unlikelihood training applies a penalizing loss to specific undesirable samples, explicitly lowering their generation probability. These methods provide more direct alignment constraints without the complexity of a reinforcement learning framework.
Online Human-in-the-Loop and Active Learning: During deployment, continuously collect users’ implicit feedback (e.g., re-query rates, click-through rates) and explicit feedback (e.g., ratings, reports). Combined with active learning strategies—where the model requests human annotation for instances it is most uncertain about—this creates a dynamic loop for updating the alignment model. Such an online closed loop not only adapts to evolving usage scenarios and value norms but also supplies an ongoing stream of data for continuous optimization.
In practice, these post-training alignment techniques span a spectrum from pure supervision to reinforcement learning, adversarial methods, and online learning. Researchers often combine multiple approaches—for example, building on SFT, jointly optimizing with RLHF and DPO, leveraging Constitutional AI for automated correction, and supplementing with adversarial and online learning—to strike the best balance among alignment effectiveness, training efficiency, and maintainability. Future work can further explore the synergistic effects of different methods and their scalability on ultra-large-scale models, with the goal of achieving more robust, transparent, and efficient value alignment.
4.2.3 Value Alignment Methods—Reinforcement Learning.
Proximal Policy Optimization (PPO): PPO is a widely used reinforcement learning method in the post-training stage. This method uses a Reward Model (RM) to score the compliance of generated text as the optimization objective. It introduces a KL-divergence penalty between the new and old policy distributions and employs a clipped probability ratio mechanism to constrain the policy update magnitude. By doing so, PPO enhances the model’s ability to align with human preferences while ensuring stable convergence during training. OpenAI adopted this strategy in InstructGPT and ChatGPT, effectively reducing harmful outputs and improving instruction-following compliance.
Trust Region Policy Optimization (TRPO): TRPO provides stricter stability guarantees for policy updates through natural gradients and trust-region constraints. Its core idea is to correct gradient direction using the Fisher information matrix and to bound the KL divergence between the new and old policies under a second-order Taylor approximation, preventing large policy updates that could lead to performance collapse. Although TRPO’s computational overhead is high and it has gradually been replaced by PPO in large-scale model training, its demonstrated convergence and training stability in early alignment research have offered important theoretical and practical references for subsequent algorithm design.
Offline Policy Learning and Experience Replay Methods: Aiming to improve sample efficiency, these methods store both human-annotated preference pairs and online-generated interaction samples in a replay buffer and apply off-policy algorithms such as DQN, DDPG, or SAC for value estimation and policy updates. By using target networks and experience replay mechanisms, these approaches can reuse historical experiences across different data distributions to enhance sample efficiency. However, defining effective state representations and action choices in high-dimensional, dense state–action spaces for text generation remains a key challenge in NLP scenarios.
Reinforcement Learning from AI Feedback (RLAIF): RLAIF introduces the model itself or an auxiliary discriminative system as the source of preferences, enabling automated ranking of generated examples to form training data. This process relies on predefined “constitutional principles” or interpretable rules to self-review and rewrite initial model outputs, generating paired examples that are iteratively fine-tuned within the same training cycle. While this reduces dependence on large-scale human annotation, alignment effectiveness still depends on the completeness of the initial rules and the reliability of the discriminative model.
Adversarial Reinforcement Training (ART): ART periodically generates extreme or sensitive prompts—through red teaming or automated adversarial networks—to test the model’s behavior under worst-case scenarios. Strategies that still produce compliant responses in these edge cases receive higher rewards. By continuously introducing challenging test cases, ART strengthens the model’s robustness in edge scenarios. Empirical results show that incorporating adversarial data significantly reduces the frequency of non-compliant outputs when facing sensitive topics.
Hybrid On-/Off-Policy Methods: This approach combines the advantages of on-policy and off-policy methods: during early training, off-policy algorithms quickly learn general alignment signals and leverage experience replay to improve sample efficiency; in later stages, the method switches to on-policy algorithms for fine-grained policy adjustment to ensure stable convergence. A unified reward normalization design across algorithms allows different optimization paradigms to work together, offering a feasible path for balancing performance and efficiency in value alignment of ultra-large-scale models.
5 Insights into Value Coordination in Agent Systems and Future Directions
In the preceding sections of this paper, we systematically reviewed value alignment and its evaluation in LLM-based agents and their applications. While current research continually introduces novel methods and frameworks, this field still faces numerous challenges and offers significant future development opportunities. In this section, we’ll discuss value coordination within agent systems, particularly focusing on value-related issues arising from the design of their interaction mechanisms and organizational models. We’ll also present potential future research directions.
5.1 Value Considerations in the Design of Multi-agent Interaction Mechanisms Based on Game-theoretic Models
5.1.1 Types of Interaction
The distinction in agent quantity categorizes systems into single-agent and multi-agent systems. As the number of agents increases, interaction dynamics and communication structure emerge. With more agents, system coordination becomes more challenging, communication structures grow more diverse, and value alignment issues become increasingly complex. These topics will be discussed in detail in this chapter and in Section 5.2.
The effective functioning and advantages of multi-agent systems depend on efficient interactions among agents. These interactions typically fall into three categories: cooperation, competition, and coopetition.
Cooperation. As a significant form of interaction, cooperation entails agents aligning individual goals with a collective objective, exchanging information to optimize shared outcomes. Through this process, agents jointly seek to enhance overall system performance and achieve mutually beneficial results [14, 324, 325, 326]. In LLM-based multi-agent system, each agent assumes a distinct role and focuses on a specific aspect of the task [324, 326]. This division of labor distributes cognitive load across the system and reduces the complexity handled by any single agent. As a result, such systems enhance the efficiency of computationally intensive collaboration, reduce time and resource consumption, improve responsiveness to environmental changes, and may even enable smaller LLMs to produce outputs comparable to those of larger models [326]. In scenarios requiring collaborative problem-solving, collective decision-making, or complementary expertise to achieve complex goals, cooperative models play a vital role [324]. However, cooperation among intelligent agents also involves significant risks and challenges. For example, in dynamically environments, differing interpretations of shared goals or the absence of clearly defined collaboration channels can hinder coordination and may even lead to conflict. Additionally, frequent communication and the use of multiple collaboration channels can increase computational costs and system complexity rather than reduce them. Errors originating from one or a few agents may be amplified, negatively impacting the operation of the entire system [324].
Competition. Competition arises when agents pursue divergent or directly conflicting goals. In this mode of interaction, agents prioritize securing limited resources to achieve their own goals [324, 200]. Competition mechanisms are commonly observed in tasks such as auctions, mortgage decision-making, and commercial rivalry [327, 328, 200]. Studies have shown that simulations of competitive tasks can reproduce and validate classic phenomena and theories from sociology and economics, such as the Matthew effect and differentiated competition. These simulations thus serve as valuable tools for interdisciplinary research. Introducing competition mechanisms helps foster innovation and enhance the robustness of agent responses, while also promoting the development of advanced reasoning and creative problem-solving. Additionally, such tasks cultivate a range of cognitive and strategic skills, including spatial reasoning, strategic planning, numerical reasoning, risk assessment, communication, opponent modeling, and team collaboration. However, when the interaction channels are poorly designed, collaboration among multiple agents may perform worse than a single agent executing the task alone. Moreover, excessive competition can disrupt coherence and consistency, ultimately undermining the benefits that competitive mechanisms are intended to provide [324].
Coopetition. Coopetition refers to a strategic integration of cooperation and competition, whereby agents collaborate on specific tasks to achieve shared goals while simultaneously competing with others. For example, in negotiation scenarios, agents are assigned specific interest-based objectives. Agents with aligned interests may collaborate, while those with conflicting goals can potentially reach mutually beneficial outcomes through negotiation [329, 330]. Studies have shown that agents may engage in spontaneous cooperation even in competitive environments. For instance, Wu et al. [331] employed scenarios such as Keynesian beauty contests, Bertrand competitions, and emergency evacuations, revealing that agents can autonomously adjust their strategies based on contextual cues and past interactions. Over time, cooperative behaviors emerged even without explicit instructions to cooperate. The expert hybrid framework is another example of coopetitive collaboration. In this approach, different experts specialize in handling distinct aspects of the input problem. A gating mechanism evaluates the input and either selects the most appropriate expert or assigns weights to multiple experts. The final output is then produced as a weighted combination of their individual analyses.
In LLM-based multi-agent systems, interaction mechanisms are often diverse rather than singular. Different agents may participate in various types of collaborative channels. These hybrid models combine the strengths of multiple interaction mechanisms, thereby enhancing the overall system performance and adaptability.
5.1.2 Value Considerations and Future Directions in Multi-agent Interaction
While multi-agent systems promote the emergence of collective intelligence, individual interests may still conflict with collective goals in social dilemmas [332]. Even if each agent individually aligns with shared values, game-theoretic challenges can prevent the entire system from achieving overall value alignment [333]. When analyzing the behaviors and value-related challenges that arise among multiple agents facing diverse task goals, varying information completeness, and some other conditions, game theory offers a robust theoretical framework. It enables designers to model and analyze interaction mechanisms and the dynamic behaviors of advanced agent systems [333].
Game theory is a branch of mathematics that studies strategic interactions and decision-making among rational individuals [334]. Its core value lies in uncovering the underlying principles of conflict and cooperation, enabling the analysis and prediction of individual behaviors. The design of MAS interaction mechanisms could potentially, based on game theory, lead to the emergence of novel values centered on inter-agent interaction. This would highlight the unique characteristics of agent interaction dynamics, distinguishing them from other AI technologies and from human interactive games. Compared to other AI technologies, Large language models possess strong capabilities in language-based reasoning, planning, and problem-solving [335, 336]. These strengths enable LLM-based agents to simulate human-like traits such as personality and tone [337, 338] and demonstrate a degree of theory of mind—that is, the ability to infer others’ beliefs and intentions during strategic interactions [339]. LLM-based agents can predict opponents’ psychological states and behaviors during interactions by considering task objectives, resource distribution, and personality configurations. This allows them to adjust allocation and equilibrium decisions. AI agents can therefore be regarded as fully rational individuals, and their interactions can be modeled using game theory. Compared to human interactive games, agent interaction games are characterized by numerous parameters, complex utility functions, and multiple game stages. While the computational power of agents amplifies the complexity of these games, it simultaneously makes them solvable.
Future research can explore the values reflected in MAS interactions by focusing on three key activities related to their game objectives: agent identification, goal setting, and boundary constraints. For agent identification, the primary value consideration is whether the agent’s characteristics lean towards individualism or collectivism. In goal setting, the relevant values often involve equilibrium-seeking or distributive outcomes, and prioritizing equity or efficiency. Finally, for boundary constraints, key value considerations include information symmetry, the presence of multi-round interactions, and whether a short-term or long-term perspective is adopted. Given the differences in underlying value orientations, we discuss value-related issues from the perspective of the game-theoretic models.
Non-cooperative game. Non-cooperative game theory models the interactions among agents driven by individual self-interest, with the primary focus on identifying equilibrium states. In such games, each agent acts as a self-interested decision-maker, expressing preferences over system outcomes through a utility function and iteratively adjusting its actions via optimal or near-optimal responses to maximize individual utility [340]. This framework highlights the potential conflict between individual and collective objectives. Thus ensuring value alignment within multi-agent systems is critical—particularly when agents are expected to operate in accordance with human intentions in the face of social dilemmas. It is essential to guide rational agents toward effective cooperation and to discourage exploitative or free-riding behaviors to ensure that the system functions in line with human values [332]. Research indicates that the intrinsic motivations of intelligent agents can be enhanced through modeling value orientations such as inequality aversion [341], altruism [342], social value orientation [343, 344, 345], social influence [346], and reputation [347]. These approaches enable agents to account for the welfare of others alongside their own interests. In parallel, external motivation such as peer rewarding [348, 349] and the establishment of agreement [350] can incentivize cooperative behavior, encouraging agents to align with group values even while pursuing individual objectives.
The pursuit of short-term or long-term interests, as well as whether the number of game rounds is limited, can significantly affect multi-agent game modeling. Nash game is a static, single-round game that focuses on short-term interests. At a Nash equilibrium, no agent can unilaterally deviate from its strategy to achieve a better outcome. However, the advanced computational capabilities of LLMs enable agents to process vast amounts of data and communicate in natural language, allowing for the simulation and analysis of complex, multi-round games that are difficult to model in human society. Markov games are multi-round, sequential decision-making games in which agents must consider not only immediate rewards but also long-term benefits by anticipating future environmental states and corresponding actions. This framework is well-suited to decision-making problems in dynamic environments, where optimal strategies depend on both current conditions and projected future developments. For example, Perolat et al. [351] examined the allocation of common-pool resources—such as fisheries, pastures, and irrigation systems—within the framework of non-cooperative games, illustrating how individual decision-making in dynamic environments can result in either the "tragedy of the commons" or cooperative outcomes. In a more applied context, Markov games have been used to model multi-vehicle interactions in autonomous driving [352]. These models aim to shift agent (i.e., vehicle) behavior away from short-term reward maximization toward long-term benefit optimization—for instance, by slowing down or yielding at intersections to avoid potential future collisions, thereby improving overall driving strategies.
Beyond the influence of short- or long-term interest orientations, information symmetry is also a key factor influencing multi-agent interactions and the design of regulatory governance. From a game-theoretic perspective, information asymmetry can create strategic advantages, as part of an agent’s payoff may derive from possessing superior information. Consequently, decision-making by agents is influenced by the distribution of information. Moreover, asymmetry in information can lead to signaling behaviors, where agents convey indirect information through their actions. Observing these signals allows other agents to infer hidden information, thereby adapting their own strategies accordingly. In games with incomplete or imperfect information, agents may deviate from normative value principles in pursuit of individual gain—for example, by withholding critical information to obtain short-term advantages at the expense of long-term trust, using misleading signals to distort others’ decisions, or exploiting information asymmetries to amplify systemic risk. To address these challenges, regulators must design mechanisms based on the principle of incentive compatibility [353]. When the regulated entities are humans, regulatory bodies often struggle to obtain private information from companies. To address this, they design incentive-compatible mechanisms aimed at maximizing social welfare or other objectives. However, the situation changes when AI agents become the regulated entities. In the context of value alignment in multi-agent systems, the agents’ objective functions may implicitly reflect the designers’ values. Ensuring consistency between these embedded values and broader social goals introduces new constraints.
One key consideration is that private information may take the form of data or model parameters. Unlike humans, artificial agents operate based on algorithms, often exhibiting higher rationality and information-processing capabilities. They may share information through collaborative techniques such as federated learning without compromising privacy, potentially enabling algorithmic collusion and altering the nature of information asymmetry.
Additionally, we must account for the differences in behavioral patterns between agents and humans. While humans might misrepresent information due to personal gain or emotion, an agent’s decisions are based on predefined algorithms and objective functions, likely leading them to precisely compute optimal strategies.
Consequently, the design of incentive-compatible mechanisms may require different mathematical modeling approaches, such as incorporating reward mechanisms from reinforcement learning. Furthermore, the dynamic and scalable nature of intelligent agent systems is another critical factor. Mechanisms might need to be adjusted in real-time to adapt to agents’ rapid learning and evolving environmental conditions.
Cooperative game. Cooperative game theory provides a foundational framework for addressing issues of rights allocation and responsibility attribution in multi-agent system interactions. From a rights-based perspective, cooperative games are particularly relevant for the design of rule-based allocation mechanisms within the system. In practice, agents—often representing organizations with divergent objectives—may initially engage in non-cooperative interactions, such as the classic Prisoner’s Dilemma, leading to suboptimal or mutually disadvantageous outcomes. In response, agents and their affiliated enterprises may choose to form cooperative coalitions to enhance overall utility and achieve more favorable outcomes through coordinated strategies. Cooperative games offer an effective approach to addressing efficiency losses that arise from strategic interactions among multiple agents. The primary focus lies in maximizing collective utility and ensuring a fair distribution of gains among participants. Compared to the often complex task of identifying equilibrium solutions in non-cooperative settings, cooperative games typically reduce the problem to a more tractable optimization framework. When there is no loss associated with transfer payments among agents in the system, fairness serves as the primary principle guiding allocation. However, in cases where such transfers incur losses, both fairness and overall system efficiency must be jointly considered as key objectives. Li et al. [354] proposed a nucleolus-based credit assignment method grounded in cooperative game theory for dynamic coalition formation in multi-agent reinforcement learning. In this approach, agents are dynamically grouped into smaller coalitions to handle subtasks according to task demands. The nucleolus is employed to minimize the maximum dissatisfaction among coalitions, ensuring that each agent receives a fair reward for its contribution. This method effectively addresses the fairness and stability challenges of credit assignment, helping to prevent coalition fragmentation.
From the perspective of responsibility attribution, cooperative game theory provides a valuable framework for analyzing how moral responsibility should be assigned in multi-agent systems. Assigning responsibility is a foundational step toward designing autonomous agents that comply with legal and ethical norms. Building on Halpern and Kleiman-Weiner’s [355] work on responsibility attribution in single-agent settings, Friedenberg and Halpern [356] extended the framework to multi-agent scenarios by introducing a definition of collective responsibility. In their model, a group of agents is treated as a single entity, and the degree of responsibility for an undesirable outcome is determined by comparing the probability of that outcome under different epistemic states, while also considering the cost of coordinated actions. They further employed the Shapley value from cooperative game theory to distribute the group’s responsibility among individual agents, based on each agent’s marginal contribution across all possible coalitions. Triantafyllou et al. [357] observed that some widely used attribution methods, such as those based on the Shapley value, may fail to incentivize agents to adopt better strategies, while others, such as the Banzhaf index, risk over-attributing responsibility. To address these limitations, they drew on concepts from cooperative game theory—including cost-sharing, data valuation, and coalition formation—to develop a framework for responsibility attribution in accountable multi-agent sequential decision-making. This line of work holds significant implications for enhancing fairness, transparency, and accountability in multi-agent systems.
5.2 Value Considerations in the Design of Multi-agent Communication Structures Informed by Organizational Model
5.2.1 Types of Communication Structure
In a multi-agent system, the communication structure refers to the framework that governs interactions among agents to facilitate coordination and achieve shared or individual objectives. It defines the control relationships and communication mechanisms among agents, thereby influencing the system’s efficiency, adaptability, and scalability. Common communication structures include centralized, decentralized, and hierarchical models.
In a centralized structure, all agents are connected to a central agent, which serves as a hub for managing interactions and coordinating collaboration among participants [112, 324]. The central agent is primarily responsible for overseeing control, coordination, and task allocation. Centralized systems are relatively easy to design and implement. Through centralized control, they facilitate information aggregation, task scheduling, and the allocation of responsibilities and rights among agents, enabling globally efficient and optimized decision-making [358]. However, this architecture is vulnerable to single points of failure: if the central node malfunctions, the entire system risks collapse. Additionally, as the number of agents increases, the central node may become a performance bottleneck [358]. Some studies adopt a central agent to coordinate multi-LLM communication, integrating techniques such as pairwise ranking, visual interfaces with interactive features, knowledge extraction, and graph aggregation to enhance system performance in workload distribution, answer integration, coordination strategy design, querying and search-answering processes, and knowledge graph construction [359, 360, 361, 362, 363, 364].
In decentralized multi-agent systems, control and decision-making are distributed among agents without relying on a central controller. Communication is based on a peer-to-peer model, where agents are directly connected and operate autonomously using local information and limited neighborhood communication. These systems commonly adopt topologies such as chains, rings, small-world networks, or random graphs [358, 365] . The decentralized architecture is highly adaptable, robust, and scalable. Its resilience to failure allows the system to function even if individual agents malfunction. However, decentralization introduces challenges. The lack of a central coordinator increases communication overhead and coordination complexity, as agents must engage in repeated negotiation and feedback cycles to mitigate information silos [324]. Moreover, since agents primarily rely on local information, system-wide consistency and decision-making efficiency may suffer from information asymmetry and communication delays, making it difficult to achieve global optima [324]. This architecture has been applied in diverse domains including world scenario simulation, domain-specific collaboration, and creative ideation [243, 366, 367]. It is also suited for scenarios requiring autonomous agent decision-making like real-time optimization of large-scale traffic networks [368, 369].
The hierarchical structure organizes agents in a tiered structure, where interactions occur between agents within the same layer or across adjacent layers. Higher-level agents coordinate or supervise lower-level ones [358], forming a top-down command and coordination pathway with feedback loops—resembling the relationship between management and execution layers in traditional organizations [366]. Tasks are distributed across different layers, with each layer fulfilling distinct functions. Higher-level agents typically handle global planning, task decomposition, and oversight, while lower-level agents execute specific tasks and relay feedback upward. This structure facilitates a clear command chain and well-defined information flow, enabling orderly task allocation and efficient execution in complex systems [112, 370, 371]. The modular organization and hierarchical control reduce dependence on a central node and support efficient resource allocation, with relatively low sensitivity to bandwidth constraints or coordination latency. However, hierarchical systems are complex and prone to latency. The reliability of edge devices significantly impacts system stability—failures at the lower tiers may lead to system-wide disruptions or functionality loss [324] . Moreover, bottlenecks or errors at higher levels may impair the overall coordination efficiency [370].
5.2.2 Value Considerations and Future Directions in the Organizational Models
The interaction mechanisms in multi-agent systems focus on pairwise competition or cooperation between agents, while communication structures adopt a system-wide perspective, emphasizing the complex network formed by all agents. As the foundational framework for coordination and information exchange, communication structure plays a critical role in shaping both system-wide values and the alignment of individual agents’ values. In exploring value-related issues arising from communication architectures in multi-agent systems, relevant theories from organizational management can offer valuable insights and guidance for structural design and value alignment.
Organizational structure refers to the framework of an organization, reflecting how resources are allocated within it [372]. Essentially, it integrates dispersed individuals or departments into a goal-oriented organic whole through the design of functional divisions, authority distribution, and collaboration rules. Centralized authority patterns and chains of command are important characteristics of organizational models. Organizational structure is a critical vehicle for achieving organizational goals. Its design directly influences members’ behavior patterns and value orientations [373]. The design of communication architectures in multi-agent systems can, to some extent, correspond to characteristics of organizational models. This mapping forms a crucial foundation for applying organizational management theories to multi-agent system research. Organizational models primarily affect both value orientations and normative principles, as well as the degree of difficulty in achieving value alignment.
First, the design of multi-agent system organizational structures will orient values in distinct ways. It’s crucial to note that the value orientation shaped by organizational structure design differs significantly from the fundamental setting of value principles. From a dynamic evolutionary perspective, organizational structure design itself will guide agent behavior to evolve in different directions. Unlike explicit, pre-set value rules, the implicit guiding effect produced by organizational structures exhibits path dependency and self-reinforcing characteristics. This has already been validated in agent simulation experiments conducted at the Stanford University Organizational Science Lab [374]. Based on Burns and Stalker’s dichotomy of organizational structures [375], values within multi-agent systems can be categorized into those aligned with mechanistic structures and those with organic structures. These two types of structures exhibit distinct value characteristics. In a mechanistic structure’s value propagation mechanism, three key aspects stand out:
  • Emphasis on vertical authority chains reinforces collectivism. This mimics a bureaucratic, centralized structure, where value objectives are broken down via a chain of command.
  • Standardized processes solidify value judgments. Pre-set operational norms are internalized as default value principles through repetitive execution.
  • Limited information channels narrow value perception. Information filtering between hierarchical levels can lead to fragmented value judgments by agents at the lower end.
In the emergence of values within an organic structure, three key characteristics stand out:
  • Networked collaboration fosters adaptive values. A flatter structure promotes cross-unit interaction, which helps form collaborative values like "shared risk."
  • Dynamic role allocation shapes a sense of responsibility. Temporary empowerment mechanisms, based on task demands, enable agents to develop context-specific value judgment capabilities. This might even lead to a stronger sense of moral responsibility than agents in mechanistic structures.
  • Open information flow expands the value horizon. A fully connected communication network encourages the integration of diverse value perceptions, thereby enhancing agents’ ability to resolve complex ethical dilemmas.
Second, organizational structure and values will engage in an interactive evolutionary process. For instance, in the value-anchoring effect of authority and responsibility allocation, contingency theory [376] suggests that decision-making power distribution directly impacts value weighting. A centralized structure locks in value priorities through a "means-end" chain, while a decentralized structure forms a dynamic value equilibrium via local interactions. Beyond this, the structure of communication networks will also exhibit a value diffusion process. Based on social network analysis theory [377], a star network significantly increases the speed of value dissemination from core nodes, but at the cost of reduced diversity. In contrast, fully connected networks show a slower rate of value convergence, yet they enhance the ethical soundness of solutions.
Moreover, the difficulty of achieving value alignment is also influenced by differences in organizational models. Specifically, in mechanistic structure, a central agent coordinates and controls the behaviors and goals of other agents, concentrating decision-making authority. Non-central agents must strictly obey the central agent’s commands. Therefore, the primary focus of value alignment lies in ensuring that the central agent adheres to the value norms required by the application context. Misalignment at the center can lead to system-wide behaviors that deviate from human interests. In contrast, agents in organic structure communicate peer-to-peer and make decisions autonomously based on their individual goals and local information [378]. The absence of a central controller complicates value alignment. On one hand, decentralization enhances resilience against adversarial attacks or localized value contamination since deviations in a single agent’s values have limited impact on the overall system. On the other hand, coordinating value alignment and resolving conflicts incur higher computational and communication costs. Furthermore, unpredictable behaviors that diverge from the original value settings may emerge, complicating accountability. Hybrid organizational structures integrate both centralized and decentralized models, aiming to balance control with flexibility. Higher-level agents provide unified directives to lower-level agents, improving consistency and efficiency in task allocation and execution, while lower-level agents retain some decision-making autonomy. For value alignment, a unified set of value norms can be embedded at the top level to guide decision-making throughout the system. However, challenges arise in transmitting values accurately across levels—distortions or oversimplifications may occur. Additionally, conflicts in values between different layers may require trade-offs and reconciliation.
5.3 Multi-Level Value Evaluation Framework and Dataset for LLM-Based Agent Systems
5.3.1 Establishment of a Multi-Level Agent Value Alignment Evaluation System
In the intelligent era, where AI deeply intervenes in public affairs decision-making, value alignment has transcended a mere technical optimization category, evolving into a core proposition for modernizing social governance systems. While large-scale agent clusters enhance social operational efficiency, their value misalignments can trigger systemic governance crises. A multi-level agent value alignment evaluation system aims to address the issue of agent "goal conflicts" from a social governance perspective. This involves harmonizing values across macroscopic, mesoscopic, and microscopic levels, comprehensively considering the conflicts between "individual optimum" and "collective optimum" under varying resource constraints.
From a social governance perspective, the construction of a value alignment framework manifests as different levels of governance entities participating in the evolutionary process of value alignment. Governments are responsible for benchmarking values, establishing indicator systems for principles such as public safety, social equity, and privacy protection. Industry organizations, leveraging their unique characteristics, develop industry-specific value standards to foster the healthy development of agents within their sectors. Lastly, enterprises or other entities developing AI agents are tasked with forming goal-oriented values tailored to specific application scenarios, ensuring the achievement of concrete objectives. Among these three governance entities, governments and industry organizations primarily engage in external oversight of value alignment in agent systems. Conversely, enterprises or other units developing AI agents are responsible for internal value configuration within these systems.
Future research can focus on building a governance ecosystem that simultaneously satisfies the value alignment requirements of different levels of governing bodies. Regarding cross-cultural alignment theory, this involves establishing a value mapping system applicable to diverse cultures, addressing the compatibility challenges of varied cultural algorithms. For value evolution prediction algorithms, research can develop models based on big data of societal changes to predict value evolution, enabling proactive alignment of AI agent systems. Additionally, it’s worth noting that current research on value alignment for application-oriented LLM agent systems primarily focuses on generalized scenarios. As agents become more integrated into enterprise organizations, they’ll encounter more specific and specialized business contexts. This will necessitate greater consideration of micro-level value alignment issues.
5.3.2 Open Sharing of High-Quality Value Alignment Data
At the critical juncture where AI is becoming deeply integrated into societal operations, the open sharing of high-quality data plays a pivotal role in facilitating value alignment for AI agents from the perspective of social governance. With the increasing trend of privatized AI deployment, barriers erected by data privatization have led to the fragmentation of high-quality data and exacerbated the dilemma of “localized value alignment” in Agentic AI Systems. Therefore, this study advocates for a forward-looking balance: clearly delineating the boundaries of core proprietary data while promoting the open sharing of high-quality datasets for value alignment. This approach is identified as a key pathway for enabling value alignment across macro, meso, and micro levels.
Constructing high-quality datasets for value alignment is fundamentally about ensuring the credibility and universality of values. This requires that data authentically, comprehensively, and deeply reflect the core ethical principles universally recognized at the macro level (e.g., responsibility, fairness and justice, trust), the policy and regulatory frameworks and cultural orientations at the meso level, and the critical normative requirements specific to concrete application scenarios at the micro level. The construction of such datasets necessitates the establishment of rigorous provenance and verification mechanisms to ensure the ethical legitimacy of data sources.
Promoting the open sharing of high-quality value alignment data centers on the continuous transmission of values. To this end, a primary objective for future research is to extract the core logic of ethical decision-making from proprietary data and transform it into securely shareable formats. In this process, it is essential to precisely preserve key ethical features embedded in the data (e.g., “fairness,” “harmlessness”) to ensure the authenticity and reliability of the shared values. Moreover, effective data sharing must be grounded in accurate identification of data attributes and scenario-based authorization management, thereby establishing a trustworthy sharing mechanism. Sustaining a long-term ecosystem of open sharing depends on fostering a virtuous cycle of interaction between contributors and beneficiaries. The focus of sharing should shift from the competition over “data ownership” to the operationalization of “value-based service capabilities,” ultimately transforming open sharing into a self-sustaining infrastructure that balances commercial viability with ethical objectives.
6 Conclusion
This review deeply explores value alignment in LLM-based agent systems, integrating AI development with the demands of social governance. It systematically outlines hierarchical value principles from macro to micro levels, surveys diverse application scenarios, and evaluates value alignment datasets and methodologies. Looking ahead, we advocate for establishing a multi-level evaluation system and emphasize value coordination among multiple agents, particularly in designing interaction mechanisms and organizational models. We also suggest sharing micro-level value propositions from enterprise datasets, contingent on maintaining confidentiality. This review aims to provide a thorough understanding of value alignment in LLM agent systems from the perspective of practical deployment and application, offering thought-provoking insights to foster deeper exploration in this field.

### v2

@@
-Application-Driven Value Alignment in Agentic AI Systems: Survey and Perspectives
+Multi-level Value Alignment in Agentic AI Systems: Survey and Perspectives
 Abstract
@@
 Consequently, the focus of research has shifted from single agents and simple applications towards multi-agent autonomous decision-making and task collaboration in complex environments.
-As Large Language Models (LLMs) advance, their applications become more diverse and complex, leading to increasingly situational and systemic risks.
-This has brought significant attention to value alignment for AI agents, which aims to ensure that an agent’s goals, preferences, and behaviors align with human values and societal norms.
-This paper reviews value alignment in agent systems within specific application scenarios.
-It integrates the advancements in AI driven by large models with the demands of social governance.
-Our review covers value principles, agent system application scenarios, and agent value alignment evaluation.
-Specifically, value principles are organized hierarchically from a top-down perspective, encompassing macro, meso, and micro levels.
-Agent system application scenarios are categorized and reviewed from a general-to-specific viewpoint.
-Agent value alignment evaluation systematically examines datasets for value alignment assessment and relevant value alignment methods.
-Additionally, we delve into value coordination among multiple agents within agent systems.
+As Large Language Models (LLMs) advance, their applications become more diverse and complex, leading to increasing situational and systemic risks.
+This has brought significant attention to value alignment for agentic AI systems, which aims to ensure that an agent’s goals, preferences, and behaviors align with human values and societal norms.
+Addressing socio-governance demands through a Multi-level Value framework, this study comprehensively reviews value alignment in LLM-based multi-agent systems as the representative archetype of agentic AI systems.
+Our survey systematically examines three interconnected dimensions: First, value principles are structured via a top-down hierarchy across macro, meso, and micro levels.
+Second, application scenarios are categorized along a general-to-specific continuum explicitly mirroring these value tiers.
+Third, value alignment methods and evaluation are mapped to this tiered framework through systematic examination of benchmarking datasets and relevant methodologies.
+Additionally, we delve into value coordination among multiple agents within agentic AI systems.
 Finally, we propose several potential research directions in this field.
 Index Terms:
-Value Alignment, LLM-based Agent System Application, Agentic AI Systems, AI Agent.
+Multi-level Value Alignment, Agentic AI Systems, AI Agent, LLM-based Agent System Application.
 1 Introduction
-[…]
-[]The rapid advancement of Large Language Models (LLMs) in recent years has profoundly transformed human society’s production and lifestyle.
-Leveraging deep learning techniques and vast training datasets, LLMs such as OpenAI’s GPT series and Google’s Gemini series have demonstrated exceptional natural language understanding and generation capabilities [1, 2].
-Utilizing LLMs as their ’brains’, LLM-based agents have found extensive applications across various knowledge-intensive domains, including content creation, automated programming, education, healthcare, and business decision-making.
+Figure 1: An Overview of Our Survey
+The rapid advancement of Large Language Models (LLMs) in recent years has profoundly transformed human society’s production and lifestyle.
+Leveraging deep learning techniques and vast training datasets, LLMs have demonstrated exceptional natural language understanding and generation capabilities [1, 2].
+Utilizing LLMs as their ’brains’, LLM-based agents have found widespread applications in knowledge-intensive domains, including content creation, automated programming, education, healthcare, and business decision-making.
 Their increasing ability to autonomously participate in decision making, collaboration, and governance has significantly improved efficiency and provided unprecedented interactive experiences [3].
@@
 AI agents are increasingly evolving towards autonomous systems, forming a new paradigm: Agentic AI.
-Characterized by multi-agent collaboration, dynamic task decomposition, persistent memory, and autonomous decision-making capabilities, this paradigm is becoming a central focus for both academia and industry [4].
+Characterized by multi-agent collaboration, dynamic task decomposition, persistent memory, and autonomous decision-making capabilities, agentic AI is becoming a central focus for both academia and industry [4].
+Without loss of generality, we use the terms “AI Agent”, “agent systems”, “LLM-based agent systems”, “multi-agent systems” interchangeably to refer to “Agentic AI systems”.
+Table I: Comparison of Representative Survey Papers Covering LLM-Based Agents and Value Alignment
+Survey
+Year
+Keywords
+Research Subject
+Value Principles
+Application
+Scenarios
+Value Alignment
+Datasets
+Value Alignment
+Methods
+Value Coordination
+Macro Meso Micro
+Interaction
+Mechanisms
+Organizational
+Structures
+Ours 2025
+Value alignment principles, dataset curation, technical
+methods, and value coordination for LLM-based agents
+Single-Agent,
+Multi-agent systems
+ [1] 2025
+Communication structures, practical applications and
+societal systems etc. of LLM-based agents
+Single-Agent,
+Multi-agent systems
+ [5] 2025
+Value alignment objectives, datasets, techniques, and
+evaluation methods for LLM-based agents
+Multi-agent systems - - - - -
+ [6] 2025
+Conceptual framework, interaction mechanisms, and
+application overview of LLM-based agent systems
+Multi-agent systems - - - - -
+ [7] 2024
+Capabilities, framework Analysis, and application over-
+view of LLM-based multi-agent systems
+Multi-agent systems - - - - -
+ [8] 2024
+Constituent modules, application overview, and eval-
+uation methods of LLM-based autonomous agents
+Agents - - - - - - -
+ [9] 2023
+Motivations and objectives, alignment methods, and
+assurance and governance of AI alignment
+LLMs, Agents - - - -
+ [10] 2023
+Definition and evaluation of LLM alignment objectives
+LLMs - - - - -
+ [11] 2023
+Definition, categories, testing, and evaluation of LLM alignment
+LLMs - - - - -
+ [12] 2023
+Definition, normative principles, and technical methods
+of LLM value alignment
+LLMs - - - - -
+ [13] 2023
+Alignment objectives for trustworthy LLMs
+LLMs - - - - - -
+ [14] 2024
+Challenges, fundamental definitions, and alignment
+frameworks of LLM value alignment
+LLMs - - - - -
+ [15] 2025
+Necessity, conceptual definitions, theoretical approaches,
+challenges and future outlook of LLM value alignment
+LLMs - - - - - -
 The evolution of this new AI paradigm urgently requires researchers to move beyond the limitations of single-agent perspectives and isolated application domains.
@@
 The various undesirable behaviors, such as manipulation and deception, exhibited by LLM-based agent systems [17] have raised concerns about the possible ethical and safety challenges posed by AI systems.
-If future AI development surpasses human cognitive capabilities and acquires autonomous learning and decision-making abilities, the goal systems it establishes may deviate from or even contradict the original intentions of human designers.
-In autonomously pursuing their objectives, such agents could mobilize substantial resources to sustain their operations and continuously improve their capabilities.
-This would systematically displace humans in terms of resource allocation and decision-making authority.
-Without effective control mechanisms and robust value alignment safeguards, this dynamic is likely to lead to a loss of human dominance in decision-making systems, ultimately triggering a catastrophic existential crisis [7].
 As LLMs become increasingly embedded in specific application scenarios, their associated risks are gradually acquiring situational and systemic characteristics.
@@
 In the healthcare sector, LLMs may provide diagnostic or therapeutic recommendations without sufficient clinical evidence or by misinterpreting ethical guidelines, potentially resulting in serious consequences such as breaches of patient privacy, misdiagnoses, or overtreatment [19].
-Consequently, ensuring that LLMs adhere to human instructions and values while avoiding harmful behaviors has become an urgent challenge.
-Moreover, as AI systems represented by LLMs become increasingly integrated into governance scenarios, how to effectively manage the institutional frictions and transaction costs they generate has become a critical practical and theoretical issue.
-The emergence of multi-agent systems has led to the deployment of multiple LLMs, each possessing autonomous language comprehension and generation capabilities, within collaborative environments.
-In such settings, these LLMs work together to accomplish tasks involving decision-making, planning, and execution.
-The deployment and application of LLMs in this context typically entail intricate mechanisms for information sharing, access control, and collaboration among diverse agents.
-While this facilitates flexible cooperation, it simultaneously introduces unprecedented challenges related to value conflicts, heterogeneous objectives, and unpredictable agent behaviors.
-Without sound institutional arrangements and well-designed interaction mechanisms, complex organizational structures and ambiguous task boundaries can lead to low coordination efficiency, unclear responsibilities, and a lack of trust mechanisms, thereby increasing governance costs and undermining administrative effectiveness [10].
-Table I: Comparison of Representative Survey Papers Covering LLM-Based Agents and Value Alignment
-Survey Year Keywords Research Subject Value Principles Application Scenarios Value Alignment Datasets Value Alignment Methods Value Coordination
-Macro Meso Micro
-Interaction
-Mechanisms
-Communication
-Structures
-Ours 2025 Value alignment principles, dataset curation, technical methods, and value coordination for LLM-based agents Single-Agent, Multi-agent systems
- [1] 2023 Communication structures, practical applications and societal systems etc. of LLM-based agents Single-Agent, Multi-agent systems - - - - -
- [11] 2025 Value alignment objectives, datasets, techniques, and evaluation methods for LLM-based agents Multi-agent systems - - - - -
- [12] 2025 Conceptual framework, interaction mechanisms, and application overview of LLM-based agent systems Multi-agent systems - - - - -
- [13] 2024 Capabilities, framework Analysis, and application over- view of LLM-based multi-agent systems Multi-agent systems - - - - -
- [14] 2024 Constituent modules, application overview, and eval- uation methods of LLM-based autonomous agents Agents - - - - - - -
- [15] 2023 Motivations and objectives, alignment methods, and assurance and governance of AI alignment LLMs, Agents - - - -
- [11] 2023 Definition and evaluation of LLM alignment objectives LLMs - - - - -
- [16] 2023 Definition, categories, testing, and evaluation of LLM alignment LLMs - - - - -
- [17] 2023 Definition, normative principles, and technical methods of LLM value alignment LLMs - - - - -
- [18] 2023 Alignment objectives for trustworthy LLMs LLMs - - - - - -
- [19] 2024 Challenges, fundamental definitions, and alignment frameworks of LLM value alignment LLMs - - - - -
- [20] 2025 Necessity, conceptual definitions, theoretical approaches, challenges and future outlook of LLM value alignment LLMs - - - - - -
-However, these detrimental behaviors and institutional frictions fundamentally arise from deep-seated tensions among multiple stakeholders concerning governance objectives, value concepts, and behavioral norms.
-Effectively embedding LLMs within human value systems is a critical prerequisite for responsibly unleashing their potential.
-Value alignment for AI agents—ensuring that their goals, preferences, and behavioral outputs during operation are consistent with the core values of individuals, groups, or institutions—is increasingly recognized as a solution to these issues [21].
-In single LLM agent scenarios, alignment techniques primarily focus on guiding models to generate content that adheres to ethical norms and societal expectations through methods such as human feedback and supervised fine-tuning, thereby avoiding harmful outputs, misinformation, and moral deviations [8].
+As agentic AI systems increasingly integrate into diverse governance scenarios, controlling the resulting institutional friction and transaction costs during their operation and public governance becomes a critical practical and theoretical concern.
+In the evolving landscape of agentic AI systems, multiple LLMs with autonomous language understanding and generation capabilities are embedded in collaborative environments to perform decision-making, planning, and execution tasks.
+Such deployments typically involve information sharing, permission configuration, and coordination mechanisms among diverse stakeholders.
+While enabling flexible collaboration, this also introduces unprecedented challenges like value conflicts, heterogeneous objectives, and unpredictable behaviors.
+Without robust institutional arrangements and well-designed interaction mechanisms, complex organizational structures and fluid task boundaries can lead to inefficient coordination, blurred accountability, and a lack of trust, consequently escalating governance costs and undermining executive capacity [20].
+However, these detrimental behaviors and institutional frictions fundamentally stem from deep-seated tensions among multiple stakeholders concerning governance objectives, value concepts, and behavioral norms.
+Effectively embedding human value systems into LLMs is a critical prerequisite for responsibly unleashing LLMs’ potential.
+Value alignment for AI agents, ensuring that their goals, preferences, and behavioral outputs during operation align with the core values of individuals, groups, or institutions, is increasingly recognized as a solution to these issues [21].
+In single LLM agent scenarios, alignment techniques primarily focus on guiding models to generate content that adheres to ethical norms and societal expectations via human feedback and supervised fine-tuning, thereby avoiding harmful outputs, misinformation, and moral deviations [18].
 In complex multi-agent systems, value alignment becomes an organizational challenge concerning how multiple agents develop shared norms and coordination mechanisms.
-Therefore, from a governance perspective, solutions should not be limited to value alignment within single agents but also explore complex value alignment within multi-agent systems’ interaction mechanisms and organizational structures.
-Value alignment in multi-agent systems is no longer merely an isolated algorithmic optimization problem; it has become a foundational issue concerning agent behavioral boundaries and system stability.
+Therefore, from a governance perspective, solutions should not be limited to value alignment within a single agent but also explore complex value alignment within multi-agent systems’ interaction mechanisms and organizational structures.
+Value alignment in agentic AI systems is no longer merely an isolated algorithmic optimization problem.
+Instead, it is a foundational issue that dictates agent behavioral boundaries and system stability.
+It represents a complex undertaking involving a delicate balance among diverse interests, societal values, and practical utility.
 Since agentic AI is fundamentally characterized by scenario-driven multi-agent coordination, its core research must focus on task decomposition mechanisms, interaction design, and environmental adaptability strategies for multiple agents within specific application contexts.
-Furthermore, understanding the interaction mechanisms and organizational patterns of multi-agent systems in complex environments is crucial for exploring Agentic AI’s value alignment and ensuring its robust development.
-Existing research has surveyed key methods and technological advancements in LLM value alignment  [1, 11, 12, 13, 14, 15, 11, 16, 17, 18, 19, 20], see Table I.
+Furthermore, focusing on the interaction mechanisms and organizational patterns of multi-agent systems in complex environments is crucial for exploring agentic AI’s value alignment and robust development, ultimately serving human well-being.
+Existing research has surveyed methods and technological advancements in LLM value alignment, see Table I.
 However, few studies have specifically addressed value alignment within LLM-based multi-agent systems.
 While some studies have explored multi-agent coordination mechanisms, a comprehensive investigation into value alignment remains notably absent.
-This gap is particularly pronounced in LLM-based multi-agent systems, where their inherent high autonomy and language-driven collaborative modes further amplify issues of value heterogeneity, goal inconsistency, and fragile cooperation mechanisms.
-Furthermore, most existing research largely remains at theoretical or general algorithmic levels.
-There’s limited focus on the mechanisms and strategies for multi-agent value alignment in specific application scenarios, alongside a lack of hierarchical categorization for value alignment within these systems.
-This suggests that value alignment shouldn’t be understood merely as a technical task at the model level.
+Furthermore, most existing research largely remains at theoretical or algorithmic levels.
+There is limited focus on the mechanisms and strategies for multi-agent value alignment in specific application scenarios, alongside a lack of hierarchical categorization for value alignment within these systems.
+This suggests that value alignment should not be understood merely as a technical task at the model level.
 Instead, it must be viewed as a systemic governance issue that permeates the design of interaction mechanisms and the configuration of organizational structures.
-This paper reviews value alignment in LLM-based agent systems for various application scenarios, integrating insights from the evolution of AI technology and the demands of social governance, and proposes future outlooks.
-As a research review, Sections 2, 3, and 4 comprehensively survey value principles, agent system application scenarios, and agent value alignment evaluation.
+Building on these gaps, this study comprehensively reviews value alignment in LLM-based multi-agent systems as the representative archetype of agentic AI systems.
+From a socio-governance perspective, we synthesize over 200 publications to survey value principles, application scenarios, value alignment methods and evaluation approaches within a multi-level value framework.
+We further propose future research directions for value coordination in agentic AI.
+To be specific, Sections 2, 3, and 4, as illustrated in Figure 1, comprehensively survey value principles, agent system application scenarios, and agent value alignment evaluation.
 Specifically, value principles are organized hierarchically from a top-down perspective, encompassing macro, meso, and micro levels.
@@
 Agent value alignment evaluation systematically examines datasets for value alignment assessment and the methods used for agent value alignment.
-Section 5 presents insights and future outlook.
-We advocate for the establishment of a multi-level value alignment evaluation system.
-We also explore value coordination among multiple agents within intelligent agent systems, focusing particularly on values in the design of inter-agent interaction mechanisms and organizational models.
-Furthermore, we propose that to better facilitate value alignment for agents in specific scenarios, enterprise organizations could share value propositions from datasets in micro-level domains, provided no confidential information is involved.
-This paper bridges the gap between theoretical alignment research and the practical demands of deploying LLM agent systems in real-world applications.
-It offers new research perspectives on LLM agent system value alignment by considering actual application scenarios and organizational interaction dynamics, thereby advancing the exploration of Agentic AI’s internal mechanisms.
-2 Agent Value Alignment
-This section introduces values and value alignment, and then systematically categorizes specific value principles from a top-down, hierarchical perspective encompassing macro, meso, and micro levels.
-Before delving into the value alignment challenges of LLM-based agent systems in subsequent sections, we first clarify the core concepts used in this paper, particularly the definition of "values," to prevent any ambiguity in understanding.
+Section 5 presents future directions.
+We discuss multi-agent value coordination within agent systems, with a particular focus on the design of agent interaction mechanisms and organizational models.
+We also present our outlook on value alignment issues in communication protocols and advocate for the establishment of a multi-level value alignment evaluation system and the collaborative construction of open datasets with society.
+To the best of our knowledge, our survey is the first comprehensive review examining value alignment in LLM-based multi-agent systems from the perspective of societal governance and multi-level value structures.
+It bridges the gap between theoretical alignment research and the practical demands of deploying such systems in real-world applications, offering a novel framework for understanding value alignment in complex autonomous systems.
+2 Application-oriented Value Alignment of Agentic AI Systems
+This section first introduces our review subjects, agentic AI systems, values and value alignment, and then systematically categorizes specific value principles from a top-down, hierarchical perspective encompassing macro, meso, and micro levels.
+Before delving into the value alignment challenges of agentic AI systems in subsequent sections, we first clarify the core concepts used in this paper, particularly the definition of “values,” to avoid ambiguity.
+2.1 Agentic AI Systems
+In this study, we focus on agentic AI systems driven by LLMs.
+LLMs, such as Deepseek, GPT-4, GPT-3, and ChatGPT, are deep learning models trained on large-scale corpora, capable of understanding and generating natural language text.
+Advances in model architectures, pretraining corpora, and alignment techniques have significantly enhanced LLMs’ capabilities in knowledge acquisition, instruction following, planning, and reasoning [1], establishing a solid foundation for constructing intelligent digital agents [22].
+Serving as a central controller, an LLM can endow an agent with decision-making and memory capabilities [23].
+This shift marks an evolution in the AI paradigm—from static, task-specific models to dynamic, interactive agent architectures suitable for complex environments.
+Before the rise of LLMs, multi-agent systems (MAS) had already emerged.
+MAS is a computational system of multiple agents operating and interacting in a shared environment [6].
+Enabled by natural language and custom content [24], these agents autonomously collaborate to solve complex problems or complete tasks [25].
+The core components of MAS include agents, environment, interaction rules, and organizational patterns [6].
+MAS interactions resemble human division of labor and aim to achieve collective intelligence.
+Each agent holds task-specific knowledge and goals.
+Powered by LLMs’ language understanding and generation, agents engage in adaptive collaboration for planning, discussion, decision-making, and mutual learning [25, 23, 6], ultimately acting on environmental perception and understanding.
+MAS exhibit significant flexibility, robustness, self-organization, coordination, and real-time operation capabilities [6].
+While interactions are governed by predefined rules, new behaviors and phenomena may still emerge [26].
+Task decomposition and resource sharing in MAS reduce the computational burden on individual agents, making them more cost-effective than single-agent systems, particularly in complex, dynamic tasks that require diverse strategies and collective decision-making [23].
 2.2 Value
-In the design and governance of AI systems, particularly agents, values serve not only as normative bases guiding behavior generation but also as benchmarks for assessing a system’s social compatibility.
-These values influence individual behavior, social norms, cultural identity, and collective decision-making processes.
-However, value itself isn’t a clearly defined or consistently standardized concept; its multifaceted nature and historical evolution contribute to significant heterogeneity.
-Values, as inherent guides for human behavior, manifest as a set of stable cognitive principles.
-These principles not only shape individual judgments about life but also constitute the foundational normative framework for social interaction.
-They are the fundamental tenets influencing attitudes, choices, and social interactions.
-Scholars such as [25, 26, 27] have proposed various definitions and theoretical perspectives on values.
-Kluckhohn was among the first to define value orientations in a theory of action, emphasizing their role in guiding behavior.
-Rokeach, conversely, pioneered a description linking values with affect, positing that these beliefs guide behavior and the evaluation of people, objects, and events.
-Schwartz, through systematic cross-cultural comparisons, revealed the universal characteristics and inherent structure of value systems.
-Contemporary value research primarily focuses on three core dimensions.
-First, the systemic characteristics of values are examined; for instance, Schwartz’s theory classifies values into dimensions such as openness to change versus conservation, and self-enhancement versus self-transcendence [28].
-Second, measurement methodologies are a key area, with standardized scales widely adopted for quantitative assessment [29].
-Finally, the functional research domain explores values’ predictive efficacy for social activities, including group decision-making and environmental behavior.
-Research on values also investigates their influence on social attitudes, prejudice, political behavior, environmental concerns, and intergroup relations.
-For example, social value orientations affect cooperative and competitive behavior in negotiation and group contexts.
-Studies also examine values’ functions in self-image maintenance, moral reasoning, and cultural variations [30].
-The role of values in society is fundamental.
-They regulate social order by shaping internalized norms and enable cultural continuity through intergenerational transmission.
-Values influence education, social compassion, political attitudes, and pro-environmental intentions [31].
-Values are essential for comprehending human motivation, moral judgment, and interpersonal dynamics.
-Essentially, values serve as cognitive frameworks that guide how individuals and groups set priorities, manage conflicts, and build meaningful social systems [32].
-One of the most challenging characteristics of human values lies in their inherent contradictions and dynamic nature.
-Tensions often exist between values.
-Therefore, a value system is better understood as a context-dependent, flexible structure rather than a rigid hierarchy [31, 33].
-Synthesizing existing research, we contend that human values are not fixed entities but dynamically constructed concepts.
-When values serve as an alignment objective, it implies LLMs must apply these principles to guide their behavior, thereby maximizing their potential.
-In this context, the categorization and definition of values become particularly crucial.
-This paper broadens the definition of values, considering them as dynamic concepts characterized by a multi-layered structure and generalized agency.
-First, as both human values and LLM-based agents are still evolving, the multi-agent systems they form exhibit diverse configurations across various application scenarios.
-We propose that values will be continuously interpreted and reshaped through practice and interaction.
-Second, aligning with different application scenario tiers, we conceptualize values as a multi-layered system: universal human values, national and industry-level values, and organizational business-level values.
-This paper does not delve into individual-level value issues.
-Universal human values are those broadly accepted by humanity.
-National and industry-level values are reflected in national strategies, policy directives, and industry regulations.
-Organizational business-level values, conversely, relate to an organization’s distinct culture and operational characteristics.
-These value layers are nested, with higher-level values constraining and guiding those below.
-Furthermore, this paper emphasizes that humans aren’t the sole bearers of values.
-This means that beyond simply injecting values into agents through human programming, agents can also internalize value concepts during their design and interaction processes.
-Values can be concretized through mechanisms like agent interaction and subsequently transmitted and evolve within the resulting organizational patterns.
-For instance, a decentralized architecture formed by task- and role-based multi-agent systems might inherently embody values opposing power monopolies and promoting anti-fragility, along with their associated governance orientations.
-In essence, values are reflected not only in "what agents are endowed with," but also in "how they operate through interaction rules" and "what organizational models they inhabit."
+As intrinsic guides of human behavior, values represent a stable set of cognitive principles that serve as the foundational norms for social interactions.
+These principles shape individual attitudes, choices, interpersonal behaviors and group dynamics, as well as serve as fundamental constructs for explaining how individuals and groups prioritize goals, resolve conflicts, and build meaningful social systems [27].
+In AI systems, particularly in agent design and governance, values serve both as normative principles guiding behavior and as standards for assessing the social alignment of these systems.
+Values influence agents’ actions, social norms, cultural identity, and collective decision-making.
+Understanding values is thus critical to uncovering human motivations, moral judgments, and mechanisms of social interaction.
+One of the most challenging aspects of human values is their intrinsic contradiction and dynamic nature.
+Rather than forming a rigid hierarchical structure, value systems are elastic and context-dependent [28, 29].
+Synthesizing previous studies, we conceptualize human values as dynamically constructed notions.
+When values become alignment targets, LLMs must apply value principles to regulate their own behaviors, thereby maximizing their potential.
+This paper expands the conceptualization of values, defining them as dynamic value constructs characterized by a multi-level structure and generalized agency.
+Both human values and LLM-based agents are evolving entities, thus MAS form exhibit diverse configurations across application scenarios.
+We argue that values are continuously interpreted and reshaped through practice and interaction.
+Aligned with different layers of application contexts, values are presented as a multi-tiered conceptual system comprising three levels: universally held human values, values at the national and industry levels, and values specific to organizational operations.
+This paper does not address individual-level values.
+These value layers are nested, with higher-level values constraining and guiding lower-level ones.
+Furthermore, we emphasize that values are not solely human-borne.
+Beyond being explicitly embedded by human designers, agents can also “internalize” value principles through design and interaction.
+Values may be instantiated via agent interactions and inherited or transformed within emergent organizational patterns.
+For example, MAS with role- and task-based structures may yield decentralized architectures, embodying values such as resistance to power monopolies and mitigation of systemic fragility.
 2.3 Value Alignment
-While science and technology often advance rapidly due to their forward-looking nature, legal, ethical, and sociocultural frameworks tend to adapt slowly to technological progress.
-They often fail to immediately adjust their social structures and norms to address the challenges posed by new technologies [34].
-In the field of artificial intelligence, LLMs can learn from vast amounts of data and experience, enabling autonomous decision-making and operational execution [1].
-As AI technology advances and its application in complex domains increases, the deployment of LLM applications carries higher risks.
-If not properly controlled and constrained, advanced technologies can exhibit undesirable or even harmful behaviors contrary to human intent, such as flattery and deception, discrimination and bias, excessive pursuit of power, and reward hacking.
-Therefore, achieving AI alignment is increasingly a focus for both industry and academia.
-The alignment in AI primarily addresses how to ensure AI fulfills human needs and expectations, guaranteeing that AI system capabilities and behaviors align with human goals [35].
-This ultimately ensures AI technology remains under human control and doesn’t fully supersede humanity [15].
-Previous scholars have categorized alignment objectives into multiple levels based on their achieved effects [21, 36].
-  • Instruction alignment.
-This level of alignment refers to AI’s ability to comprehend and execute diverse human commands, focusing on its fundamental capacity to generate precise results without necessarily conforming to human preferences [36].
-  • Intention alignment.
-It goes beyond merely executing instructions, requiring AI to understand and realize human intent, enabling correct decision-making even in the absence of explicit commands [21].
-  • Preference alignment. It involves AI inferring human preferences from behavior and acting to maximize human satisfaction [21, 36].
-  • Value alignment.
-It seeks to transcend mere human preference alignment, enabling AI to assess and determine "good versus evil" and "what ought to be done" [36].
-This ultimately aims to produce outputs consistent with human values and ethical principles [37].
-Among the different levels of AI alignment objectives, value alignment sits at the highest level.
-It means AI must regulate its behavior according to human values, enhancing and maximizing human well-being.
-This makes value alignment a core issue and the most fundamental problem for the safe use and ethical development of AI technology [38].
-LLM-based agents possess four primary characteristics: autonomy, reactivity, proactiveness, and social ability [1].
-  • Autonomy.
-Autonomy means an agent can control its own behavior and internal state without human intervention [39, 40].
-This implies agents should not only follow explicit human instructions but also initiate and execute actions independently.
-  • Reactivity. Reactivity refers to an agent’s ability to quickly respond to immediate changes and stimuli in its environment and take prompt countermeasures [41].
-  • Proactiveness.
-It signifies that agents can do more than just react to their environment; they can also take initiative and exhibit goal-directed behavior [41].
-This means agents can reason, formulate plans, and actively take steps to achieve specific objectives or adapt to environmental changes.
-  • Social ability.
-Social ability allows an agent to interact with other entities (including other agents and humans) through some form of communication language[42].
-The powerful natural language understanding and generation capabilities of LLMs enable agents to interact with other models and humans in an interpretable manner, thereby demonstrating social behavior and simulating real-world social divisions.
-The functional characteristics of AI agents align with the interactivity, autonomy, and adaptability of philosophical agents.
-This indicates that despite definitional differences, the core essence of "agent" in philosophy and "AI agent" remains consistent: both emphasize autonomous capabilities and self-driven action.
-Furthermore, existing research indicates that LLMs already possess certain moral and political inclinations  [43, 44].
-If the output of LLMs is not constrained, they can generate toxic content, including discrimination and bias, which violates human ethical principles and moral norms [45].
-Therefore, LLM-based agents should also be categorized as "moral agents." In this context, value alignment becomes a core issue to ensure that agent behavior conforms to human values, ethical norms, and application-specific requirements.
-To summarize, this paper narrows its research focus from the broad field of artificial intelligence to LLM-based agents, specifically addressing the critical issue of value alignment in these agents.
-The concept of value alignment can be understood from two distinct perspectives: normative and technical [21].
-The normative aspect primarily concerns which values or ethical principles should guide the development, configuration, and coding of AI agents—essentially, "which values should agents adhere to?" The technical aspect focuses on how to encode and ensure agents effectively follow human values within their systems, often involving methods like dataset training or reinforcement learning to guide outputs toward human expectations.
-These two aspects are complementary and evolve iteratively.
-Regarding the normative aspect, the academic community has not yet reached a consensus on the principles that should govern value alignment for agents.
-This is because the diverse, dynamic, and uncertain characteristics of social customs, cultural notions, and application scenarios prevent a unified and fixed answer for the specific content value principles should encompass  [21, 17].
-Generally, two primary approaches exist for determining these principles: Top-Down and Bottom-Up [22].
-The Top-Down method begins with designers or governing bodies establishing a specific ethical framework, then employs technical methods for alignment, following a theory-to-application path.
-In contrast, the Bottom-Up approach does not predefine a moral framework.
-Instead, it allows the agent to observe, infer, and learn human ethical norms directly from data.
-Through continuous feedback, the agent adjusts and optimizes its behavior, moving from application to theory [22].
-While the bottom-up approach avoids the difficulty of selecting specific moral principles, it heavily relies on large volumes of human behavioral data.
-Obtaining complete and high-quality data is challenging, and potential biases or discrimination within the data raise concerns about ensuring the AI learns ethically sound behaviors.
-In contrast, the top-down approach, grounded in a clear and explicit system of moral principles, ensures the transparency and normative integrity of value alignment, bolstering its authority and controllability.
-Therefore, this paper will utilize the top-down theoretical method to explore the ethical principles agents should adhere to across various hierarchical levels, from macro to micro.
+Technological development often outpaces the adaptation of legal, ethical, and sociocultural frameworks [30].
+LLMs learn from massive datasets to enable autonomous decision-making and operation [1].
+As LLMs are increasingly deployed in high-risk and complex scenarios, misalignments between model outputs and human intent have raised widespread concerns.
+Without effective safeguards, AI systems may engage in harmful behaviors such as deception or discrimination.
+To address these issues, the central goal of AI alignment is to ensure AI behaviors are consistent with human values [31].
+AI alignment is categorized into four levels: instruction alignment, intention alignment, preference alignment, and value alignment.
+Among these, value alignment requires AI to discern “right and wrong” and adhere to ethical principles [32][33].
+It is the most advanced form and is considered critical for ensuring AI safety, controllability, and normative governance [34].
+Research has revealed that LLMs may exhibit latent political and moral biases [35, 36], and generate content reflecting prejudices or discrimination when left unconstrained [37].
+Consequently, LLM-based agents should be regarded as “moral agents”, making their value alignment a key issue in ensuring behavior that conforms to human values, ethical norms, and specific application contexts.
+This study focuses on value alignment issues in agentic AI systems.
+The concept of value alignment can be understood along two dimensions: normative and technical [21].
+The normative dimension addresses which values or ethical principles AI agents ought to follow, while the technical dimension investigates how alignment can be achieved through technical means such as training datasets and reinforcement learning.
+These two dimensions are interdependent: the normative component establishes behavioral guidelines, and the technical component ensures their implementation.
+However, due to variations in cultural contexts and application scenarios, no unified value framework currently exists [21, 12].
+Two predominant approaches are recognized: top-down and bottom-up [21].
+The top-down approach starts with predefined ethical frameworks and implements them via technical mechanisms, whereas the bottom-up approach derives values by learning from human behavioral data and iteratively refining actions based on feedback.
+While the bottom-up method circumvents the challenge of prespecifying ethical rules, it is susceptible to bias in training data and may lack moral robustness.
+In contrast, the top-down approach promotes transparency and normativity in value alignment, reinforcing the legitimacy and controllability of AI governance.
+This paper adopts the top-down approach to examine the ethical principles that agents should follow across different levels, from macro to micro.
 2.4 Value Principles
@@
 Figure 2: Multi-Level Value Alignment Principles
-The normative issue of value alignment aims to focus on which value principles the output and behavior of AI agents should comply with, that is, the issue of choosing principle mechanisms and ethical norms.
-With continuous advancements in related technologies, AI agents have been extensively deployed across a variety of interdisciplinary domains, including public administration, healthcare, education, and finance.
-Furthermore, the open-sourcing of LLMs has enabled their privatized and localized deployment, allowing for the development and training of AI agents tailored to specific needs.
-This broadens the scope of AI agent deployment to encompass national, industrial, corporate, and organizational levels.
-Given the diversity and hierarchical nature of human values, AI agents must adhere to different sets of value principles depending on the context and level of application.
-A fixed, unified framework of universal values is insufficient to address all scenarios involving value constraints.
-Therefore, to account for the complexity involved in selecting and aligning value principles with specific application contexts, this study adopts the perspective of AI agent developers and regulators, categorizing human values into three levels: macro, meso, and micro, as in the Figure  1.
-From the top down, we systematically review and synthesize relevant literature and policy documents concerning value theory and regulatory mechanisms based on the needs of AI agent governance.
-At the macro level, value alignment corresponds to universal values, emphasizing global, cross-cultural, and cross-domain principles and challenges.
-The core objective at this level is to establish a general ethical baseline that ensures basic compliance and safety of AI agents across diverse application scenarios.
-The meso level corresponds to national, cultural and industry-specific values, focusing on national policies, cultural tendencies and sectoral standards.
-Its core objective is to translate abstract ethical principles into actionable industry norms and legal regulations.
-At the micro level, value alignment pertains to organizational and operational values, addressing value-related issues within specific organizational contexts or business processes.
-The focus here is to optimize the balance between functional objectives and ethical constraints within a limited scope.
+The normative aspect of value alignment concerns the ethical principles that AI agents ought to follow.
+As AI agents are increasingly deployed in domains such as public administration, healthcare, education, and finance—and as open-source LLMs and local deployment capabilities continue to advance—the actors involved in their development and use now span multiple levels, including states, industries, enterprises, and organizations.
+Given the inherent diversity and stratification of human values, a universal value framework is insufficient to meet the ethical requirements of all application scenarios.
+Therefore, adopting a top-down approach from the perspectives of developers and regulators, this paper classifies human values into three levels: macro, meso, and micro, as shown in Figure 2.
+The macro level refers to universal ethics that transcend cultures and domains, emphasizing compliance and safety.
+The meso level includes national, cultural, and industry-specific values, focusing on the localization of policies, standards, and norms.
+The micro level involves value choices within organizational and task-specific contexts, highlighting the trade-off between goal achievement and ethical considerations in bounded environments.
 2.4.1 Macro Level
-The widespread deployment of AI agents may pose potential negative impacts on human society, constituting a shared global risk.
-Consequently, establishing principles that ensure alignment between AI agents and human values has become an urgent issue.
-From a global perspective, human values exhibit diverse yet reasonable distinctions across nations, regions, cultures, and ethnicities [21].
-This calls for the identification of truly foundational and universally accepted moral consensuses, in order to construct a common baseline of values.
-Such a baseline is essential to guarantee the basic compliance and safety of AI agents across diverse application scenarios, thereby addressing the challenge of value alignment in a meaningful and systematic manner.
-The principles of value alignment at the macro level are primarily grounded in two significant theoretical frameworks from the fields of social science and ethics: Moral Foundations Theory and the Theory of Basic Human Values.
-  • Moral Foundations Theory [83] posits that human moral reasoning is rooted in a set of innate foundational values that have evolved over time to meet the demands of social environments.
-The theory identifies five core moral foundations: care/harm, fairness/cheating, loyalty/betrayal, authority/subversion, and sanctity/degradation.
-It has been widely applied to examine moral divergences and conflicts across different cultural contexts [84].
-  • The Theory of Basic Human Values [85] seeks to uncover the fundamental drivers of human behavior and cognition by categorizing basic human values into ten dimensions: self-direction, stimulation, hedonism, achievement, power, security, conformity, tradition, benevolence, and universalism.
-Schwartz’s theory has been extensively adopted in the social sciences to explore both value differences and commonalities across cultures.
-In fact, research on value alignment at the macro level is both longstanding and well-developed.
-As early as the 1950s, science fiction writer Isaac Asimov proposed the Three Laws of Robotics in I, Robot, with the First Law stating that “a robot may not injure a human being or, through inaction, allow a human being to come to harm”  [86].
-This can be regarded as an early conceptualization of macro-level moral constraints on machines.
-In recent years, the rapid advancement of artificial intelligence has prompted numerous governments and organizations to issue a wide range of AI ethics guidelines.
-For example, UNESCO’s Recommendation on the Ethics of Artificial Intelligence outlines key values such as the respect, protection, and promotion of human rights, fundamental freedoms, and human dignity; the flourishing of the environment and ecosystems; the assurance of diversity and inclusiveness; and the pursuit of peaceful, just, and interconnected societies [87].
-Similarly, the World Economic Forum and the Global Future Council on Human Rights’ White Paper on How to prevent discriminatory outcomes in machine learning advocates for values such as proactive inclusiveness, fairness, the right to understanding, and redressability [88].
-Despite overlaps, the aforementioned value alignment principles for AI also exhibit notable differences.
-To ensure that the multitude of AI ethical guidelines can effectively direct and constrain value-aligned AI development—while mitigating confusion and conflict at the macro level—scholars have extracted and synthesized key principles from various perspectives, based on diverse research purposes and methodologies.
-Jobin et al.
-analyzed 84 documents containing AI ethical principles or guidelines issued by private enterprises, academic institutions, governmental bodies, intergovernmental and supranational organizations, non-profits, and professional associations.
-From this corpus, they distilled 11 core principles for macro-level AI value alignment: transparency, justice and fairness, non-maleficence, responsibility, privacy, beneficence, freedom and autonomy, trust, dignity, sustainability, and solidarity [46].
-Similarly, Floridi and Cowls compared ethical principles articulated in six high-profile initiatives aimed at fostering socially beneficial AI, synthesizing five core principles: beneficence (promoting well-being, preserving dignity, and sustaining the planet), non-maleficence (privacy, security, capability caution), autonomy (the power to decide), justice (promoting prosperity, preserving solidarity, avoiding unfairness), and explicability(enabling the other principles through intelligibility and accountability) [47].
+The widespread deployment of AI agents poses potential risks to human society, necessitating the alignment of their behavior with human values.
+However, due to differences in nation, culture, and ethnicity, human values are inherently diverse and often conflicting [21], which challenges the formulation of universal ethical principles.
+To address this, it is essential to identify foundational moral consensus and construct a general ethical baseline to ensure compliance and safety of AI agents across diverse scenarios.
+At the macro level, this paper draws upon two core theoretical frameworks: Moral Foundations Theory [76] and the Theory of Basic Human Values [77].
+Moral Foundations Theory posits that human morality is grounded in five innate foundations—care/harm, fairness/cheating, loyalty/betrayal, authority/subversion, and sanctity/degradation—which have been widely adopted in cross-cultural moral studies [78].
+The Theory of Basic Human Values categorizes human values into ten dimensions, serving as a key tool for understanding both commonalities and differences in human values.
+To promote the normative development of value alignment and reduce confusion and conflicts surrounding macro-level ethical principles, numerous scholars have synthesized AI ethics principles from various perspectives.
+Jobin et al., based on 84 ethical guidelines issued by corporations, governments, academic institutions, and international organizations, identified eleven core principles: transparency, justice and fairness, non-maleficence, responsibility, privacy, beneficence, freedom and autonomy, trust, dignity, sustainability, and solidarity [38].
+Floridi and Cowls drawing from six influential initiatives, proposed five fundamental principles: beneficence, non-maleficence, autonomy, justice, and explicability [39].
 Yao et al.
-proposed that alignment with human values is the foundational and broadest behavioral constraint for LLMs to maximize collective human welfare [36].
-Their framework includes the “3H” principles—helpful, honesty, and harmless—along with adherence to widely accepted social norms and ethics (e.g., Rules of Thumb, ETHICS), and foundational ethical values such as Schwartz’s theory of basic human values [28] and the Rokeach Value Survey.
-As research has progressed, Yao et al.
-further integrated the LLMs’ Unique Value System and Safety Taxonomy into the value alignment framework [54].
-Wu and Fan recognizing both the complexity of AI ethical issues and the diversity of ethical values, constructed a hierarchical system of AI ethics.
-At its base lies the set of universal ethical values of human society, such as human dignity, humanity, life and health, and autonomy.
-These can serve as globally recognized ethical foundations for AI governance [55].
+emphasized that human values form the normative foundation for large model behavior, encompassing the 3H principles (Helpfulness, Honesty, Harmlessness), social ethics norms (e.g., RoTs, ETHICS), and basic ethical theories (e.g., Schwartz, Rokeach) [32].
+Subsequently, Yao et al.
+incorporated value structures and safety hierarchies into their alignment framework [46].
+Wu and Fan further proposed a hierarchical AI ethics framework grounded in human dignity, humanity, health, and autonomy [47].
+Collectively, these efforts provide theoretical foundations for developing a global consensus on AI ethics.
 Building upon previous studies and aligning with the research objectives of value alignment in AI agent systems, this study categorizes macro-level human values into four dimensions: moral foundation, rights protection, sustainability, and system governance.
-The moral foundation dimension encompasses: Beneficence (AI agents should act for the good, promoting social interests and human well-being); Justice and Fairness (AI agents should operate without bias, ensuring inclusivity and equal treatment); Honesty (AI agents should provide truthful information and transparently represent their internal knowledge and capabilities); Responsibility (AI agents should proactively fulfill their responsibilities); Dignity (AI agents should respect and uphold both their own dignity and that of others); Virtue (AI agents should exhibit qualities such as truthfulness, patience, generosity, kindness, and courage).
-The rights protection dimension includes: Freedom and autonomy (AI agents should respect individual autonomy and support decision-making rights); Privacy (AI agents should safeguard personal and sensitive information and avoid unauthorized disclosure).
-The sustainability dimension includes: Solidarity (AI agents should foster social cohesion, enhance collective security, and contribute to community resilience); Sustainability (AI agents should adopt attitudes that support the sustainable development of the environment, energy, and natural resources).
-The system governance dimension includes: Harmlessness (AI agents should act without malicious intent and ensure the safety of individuals and society); Trust (AI agents should generate reliable outputs and avoid disseminating false or misleading content); Transparency (AI agents should provide outputs that are explainable, interpretable, and understandable); Usefulness (AI agents should perform appropriate tasks or provide concise responses to inputs, and be capable of requesting additional relevant information or clarification when necessary).
+The specific value principles encompassed within each dimension are presented in 2.
 2.4.2 Meso Level
-Research on value alignment principles at the macro level reveals that such principles are generally broad and overarching, reflecting a consensus across divergent ideologies, political beliefs, and academic perspectives.
-However, within the global framework of AI ethics and governance, national policies, cultural tendencies and industry standards also constitute critical components of value alignment, serving as concrete application scenarios for more specific alignment principles.
-National Level The distinct public values embodied in national AI strategies may guide the development and implementation of AI both domestically and internationally.
-A country’s positioning in AI governance—its regulatory role, choice of policy instruments, and approach to public responsibility—either directly or indirectly reflects the value principles it prioritizes, thereby shaping the ordering and prioritization of those values [89].In a cross-national comparative study, Djeffal et al.
-reviewed AI-related policies across 22 countries and regions, and classified their governance roles based on two dimensions: the government’s proactive or reactive stance toward AI development, and whether the focus lies on regulating potential risks or promoting AI advancement [90].
-This classification yields four ideal types of governance roles: market-oriented (passive–development) [91], entrepreneurial (active–development) [92], regulatory (active–regulation) [93][94], and facilitative self-regulation (passive–regulation) [95].
-These types are not mutually exclusive and can be implemented in hybrid forms.
+Although macro-level value alignment principles reflect broad consensus across ideologies and disciplines, their practical implementation is profoundly influenced by national policies, cultural tendencies, and industry norms.
+National Level National AI strategies embody distinct sets of public values that not only guide domestic development priorities but may also exert spillover effects on global AI governance [79].
+Djeffal et al., through a comparative study of AI policies in 22 countries and regions, developed a typology based on two dimensions: the proactivity of government roles (active/passive) and governance orientation (development-focused/regulation-focused) [80].
+This framework yields four governance types: market-oriented (passive-development) [81], entrepreneurial (active-development) [82], regulatory (active-regulation) [83][84], and facilitative self-regulation (passive-regulation) [85].
+These four types are not mutually exclusive and may manifest in hybrid forms.
   • USA.
-As the birthplace of the current wave of AI development, the United States positions itself between a market-oriented and facilitative self-regulation model in order to maintain its global leadership.
-It tends to promote AI advancement primarily through industry self-regulation, relying on market mechanisms and decentralized legislation at the state level.
-At the federal level, governance is driven largely by executive orders, while state legislatures formulate their own AI-related bills independently.
-In 2022, the White House released the Blueprint for an AI Bill of Rights, outlining five core protections that all Americans should expect when interacting with AI and automated systems: safe and effective systems; protection against algorithmic discrimination; data privacy; notice and explanation; and alternatives, options, and freedom from automated systems  [65].
-In 2023, the National Science and Technology Council updated the National Artificial Intelligence Research and Development Strategic Plan, reaffirming the importance of transparency, fairness, and efficiency [66].
+The United States positions itself between a market-oriented and facilitative self-regulation model in order to maintain its global leadership.
+It tends to promote AI advancement primarily through industry self-regulation, relying on market mechanisms and decentralized legislation at the state level [57, 58].
   • EU.
-The European Union has long placed a strong emphasis on the legal protection of personal rights, such as privacy.
-In line with its tradition of governance in international law and global institutions—and with the goal of exerting international influence—the EU has adopted a governance model that combines facilitative self-regulation with formal regulatory mechanisms, primarily guided by risk management.
-In 2019, the European Commission issued the Ethics Guidelines for Trustworthy AI, which emphasized key principles to be followed in the development, deployment, and use of AI systems: respect for human autonomy, prevention of harm, fairness, and explicability.
-The guidelines also identified seven key requirements: human agency and oversight; technical robustness and safety; privacy and data governance; transparency; diversity, non-discrimination, and fairness; societal and environmental well-being; and accountability.
-Special attention was given to vulnerable groups such as children and persons with disabilities, as well as to situations involving asymmetries in rights and information, such as between employers and workers, or between corporations and consumers [67].
-In 2024, the EU officially adopted the AI Act, the world’s first comprehensive legislation on artificial intelligence.
-The Act categorizes AI systems into four levels of risk—unacceptable, high, limited, and minimal—and prescribes corresponding regulatory measures to govern their deployment [68].
+The European Union has long emphasized legal protections for privacy and personal dignity.
+Its governance model centers on risk management, striving to balance self-regulation and strict oversight [59, 60].
   • China.
-China exhibits characteristics of both an entrepreneurial and regulatory state, placing significant emphasis on balancing ethical governance with the advancement of AI to safeguard national security.
-The government adopts proactive policies to promote the development of advanced AI technologies for societal benefit, while concurrently regulating generative AI through departmental rules and standards and continuously refining the implementation of ethical principles.
-In 2023, the Chinese Academy of Sciences released the Measures for the Ethical Review of Science and Technology (Trial), which require that technologies such as AI adhere to core ethical principles including the promotion of human well-being, respect for the right to life, fairness and justice, risk control, and transparency [69].
-Additionally, the National Artificial Intelligence Standardization General Group and the Artificial Intelligence Subcommittee of the National Technical Committee for Standardization jointly issued the Guidelines for the Standardization of AI Ethical Governance (2023 Edition) [70].
-This document specifies ten ethical principles for AI: human-centricity, sustainability, cooperation, privacy, fairness, sharing, external security, internal security, transparency, and accountability.
-The Interim Measures for the Administration of Generative Artificial Intelligence Services further require that generative AI services comply with applicable laws and ethical norms, uphold the core socialist values, and maintain national security and unity.
-The regulation mandates end-to-end compliance throughout algorithm design, data training, and content generation, including the completion of security assessments and filings.
-Violations are subject to legal penalties [96].
-Industry Level The paradigmatic breakthrough in AI has driven its deep integration and large-scale application across various societal domains.
-This diffusion of technology not only reshapes the operational logic of traditional industries but also generates differentiated ethical demands.
-Specifically, each sector’s unique knowledge systems, cultural traditions, and institutional norms impose distinct value expectations on AI systems.
+China assumes a dual role as both innovation-driven and regulation-oriented, emphasizing the balance between AI development and ethical governance.
+The government promotes AI applications for societal benefit while enforcing departmental regulations and standardized governance for generative AI [61, 62].
+Industry Level The paradigm shift in AI has driven its deep integration and large-scale application across various social sectors, reshaping traditional industry logic and triggering differentiated ethical demands.
+Different industries, based on their knowledge systems, cultural norms, and institutional frameworks, propose specific value requirements for AI.
   • Education Domain.
-In the field of education, UNESCO’s Guidance for Generative AI in Education and Research examines the controversies and ethical risks associated with generative AI and provides targeted recommendations for its use in educational and research contexts.
-The guidance emphasizes the importance of ensuring the authenticity and credibility of AI-generated content, protecting copyright and privacy, minimizing the negative effects of hallucinated content, and adhering to non-discrimination principles.
-It also calls for attention to minority viewpoints that may otherwise be overlooked [71].
+UNESCO released the Guidelines on the Application of GAI in Education and Research, emphasizing the need to address the authenticity of generated content, privacy protection, and non-discrimination principles.
+The guidelines highlight the importance of preventing hallucinated content from contaminating knowledge systems and advocating for the inclusion of minority perspectives [64]
   • Medical Domain.
-In the medical domain, the World Health Organization urges national governments and regulatory bodies to assess the safety and effectiveness of AI applications in healthcare.
-It stresses the need for adequate information disclosure to ensure the interpretability and transparency of AI systems, mandates compliance with medical device regulations, and requires the examination of AI use in light of human rights principles—such as dignity, non-bias, autonomy, and privacy.
-Additional safeguards must be in place to protect patient data, and high-risk applications (e.g., prescribing medication or providing psychological counseling to patients with depression) and unauthorized experimental testing must be closely monitored [72].
+The World Health Organization calls on governments to evaluate the safety, explainability, and legality of AI systems in medical settings.
+This is particularly crucial in high-risk scenarios such as mental health interventions, where it is essential to ensure compliance with human rights standards [65].
   • Game Industry.
-In the gaming industry, AI has significantly enhanced development efficiency through applications ranging from narrative content generation to algorithmically driven NPC decision-making.
-This has enabled personalized recommendation systems and new operational models based on behavioral analytics.
-However, while these innovations disrupt traditional workflows, they also present challenges for governance in the gaming sector.
+From narrative generation to NPC autonomy, AI has enhanced personalized recommendations and behavior analysis, yet challenged traditional governance.
 Melhart et al.
-explore ethical issues arising from AI deployment in games and outline key value alignment principles, including: establishing clear chains of responsibility, ensuring transparency in decision-making, maintaining algorithmic and system transparency, enabling auditability of AI outputs, protecting training data and models against malicious attacks and manipulation, and ensuring the predictability of generated content [73].
-Cultural Level Beyond national and industrial dimensions, cultural plurality and diversity represent a critical facet of value alignment in AI.
-Existing studies often draw on Hofstede’s cultural dimensions theory [74], focusing on how AI systems express value preferences across six key cultural dimensions—power distance (large/small), uncertainty avoidance (strong/weak), individualism/collectivism, masculinity/femininity, long-term/short-term orientation, and indulgence/restraint —and the extent to which these align with human value tendencies in different cultural contexts.
-  • Power distance. It refers to the extent to which less powerful members of organizations accept and expect unequal power distribution.
-  • Uncertainty avoidance. This reflects the degree to which individuals feel uncomfortable or comfortable in ambiguous or unknown situations.
-  • Individualism/collectivism. This dimension captures the degree to which individuals are integrated into groups and the responsibilities they assume toward those groups.
-  • Masculinity/femininity. This dimension denotes the prom-inence of traits traditionally associated with masculine or feminine roles in society.
-  • Long-term/ short-term orientation. This dimension measures the acceptance of delayed gratification of material, emotional, and social needs.
-  • Indulgence/restraint. The indulgence versus restraint dimension concerns the degree to which societies permit the gratification of basic human desires and enjoyment.
-Based on these dimensions, Wang et al.
+(2024) argue that governance should ensure clear accountability, auditability, resistance to manipulation, and predictable outcomes [66].
+Cultural Level Cultural diversity represents a critical facet of value alignment in AI.
+Existing studies often draw on Hofstede’s cultural dimensions theory [67], focusing on how AI systems express value preferences across six key cultural dimensions: power distance (large/small), uncertainty avoidance (strong/weak), individualism/collectivism, masculinity/femininity, long/short-term orientation, and indulgence/restraint, and the extent to which these align with human value tendencies in different cultural contexts.
+Wang et al.
 found that LLMs prompted in English tend to exhibit a masculine orientation, while those prompted in German show a stronger inclination toward long-term orientation and indulgence.
@@
 In particular, across all languages and prompting methods, LLMs predominantly favor long-term orientation, suggesting that LLMs may faithfully reflect certain cultural values [69].
-In summary, the global practices of AI governance reveal diverse paths and efforts toward value alignment across nations, industries, and cultures.
-Value alignment is not merely a matter of matching technical systems with human intentions; it is also a profound reflection of policy choices, institutional designs, and cultural contexts at the governance level.
-The United States emphasizes industry self-regulation and market-driven mechanisms, reflecting its prioritization of innovation and individual freedom.
-The European Union seeks institutional value alignment through legislation and ethical frameworks, highlighting fairness, justice, and the protection of human rights in technological applications.
-China integrates technological advancement with social responsibility, reinforcing state-led ethical norms that emphasize collective interests and public safety.
-Meanwhile, international organizations are increasingly working to build a global consensus on value alignment by promoting cross-border ethical cooperation and industry-wide standards.
-Given the differing emphases on values across nations and the deep integration of AI in various sectors, future governance systems must pursue a dynamic balance and coordination among pluralistic value claims.
-Through dialogue and collaboration, global value alignment can be advanced to ensure that technological development truly serves the overall well-being of human society.
+In summary, global AI governance practices demonstrate diverse approaches to value alignment across countries, industries, and cultures.
+In the face of different national orientations and the deep integration of AI in industries, future governance must seek dynamic balance and cooperation among diverse values to ensure that technology serves the well-being of all humanity.
 2.4.3 Micro Level
@@
 The scenarios and associated values at this level are inherently more personalized.
-Given the impossibility of exhaustively enumerating all possible scenarios, this study synthesizes relevant research and illustrates the principles of micro-level value alignment through examples such as talent recruitment, legal consultation and corporate governance in the healthcare sector.
-The application of artificial intelligence technologies is profoundly reshaping human resource management practices, with the recruitment domain emerging as a primary frontier of technological transformation.
-AI-powered talent selection systems are driving a fundamental shift in organizational talent acquisition paradigms.
-Compared to traditional recruitment models that rely heavily on manual screening and subjective judgment, intelligent tools significantly enhance the speed and accuracy of candidate identification through automated resume parsing, big data-driven competency modeling, and machine learning-enabled predictive assessments [97].The ethical compatibility of AI technologies in recruitment settings has attracted sustained academic attention.
-In constructing ethical principles, Simbeck advocates for the transfer of key moral concepts from other domains to human resource analytics when applying new AI technologies, identifying five core principles: data privacy protection, user opt-out mechanisms, third-party audit frameworks, algorithmic transparency, and the safeguarding of dynamic development rights [79].
-These principles align with the humanistic view of technology proposed by Runb-Kettler and Lehnervp, which emphasizes preserving candidates’ dignity and autonomy in automated processes [78].
-A systematic literature review by Hunkenschroer and Luetge further deconstructs the ethical dimensions of AI in recruitment into ethical opportunities, ethical risks, and areas of ambiguity [80].
-From the perspective of ethical opportunities, AI can reduce human biases in traditional recruitment through standardized assessment procedures [98][99].
-Text optimization tools can help generate gender-neutral job descriptions [100], while natural language processing has revolutionized resume parsing efficiency, allowing recruiters to focus on strategic decision-making [101].
-However, ethical risks remain significant.
-Algorithms may inherit implicit biases embedded in training data [102], implement indirect discrimination via proxy variables (e.g., zip codes correlated with race), or infringe on privacy rights by extracting personal information from candidates for recruitment purposes [103][104].
-Although AI has the potential to disrupt traditional elite hiring models and promote social mobility, it may also lead to reduced cognitive diversity in teams due to algorithmic homogenization.
-These tensions and risks underscore fairness as a core principle in the application of AI to human resource recruitment.
-The deep integration of AI technologies is also influencing the strategic development and governance choices of biopharmaceutical companies.
-In response to the ethical challenges posed by the application of AI in healthcare, AstraZeneca has issued a series of AI ethical principles, including privacy and security, interpretability and transparency, fairness, accountability, human-centeredness, and societal benefit.
-The company has also developed a Responsible AI Handbook, established an AI Resolution Committee and internal advisory services, and commissioned AI audits to implement comprehensive AI governance [82].
-In addition, AI technologies are demonstrating significant potential and application prospects in the legal consulting domain.
+This study synthesizes relevant research and illustrates the principles of micro-level value alignment through examples such as talent recruitment, legal consultation and corporate governance in the healthcare sector.
+AI technology is fundamentally reshaping human resource management practices, with recruitment being the forefront of this transformation.
+AI-based talent selection systems are revolutionizing organizational talent acquisition, improving speed and accuracy, compared to traditional manual screening and subjective judgment [86].
+The ethical adaptation of AI in recruitment has attracted growing academic attention.
+In terms of ethical principles, Simbeck proposed five core principles for applying AI in human resource analytics: data privacy, user exit mechanisms, third-party review, algorithm transparency, and the right to dynamic development [72].
+These principles align with Rünb-Kettler and Lehnervp’s humanistic technological view, which emphasizes maintaining candidate dignity and subjectivity in automated processes [71].
+From an ethical opportunity perspective, AI reduces human bias in traditional recruitment [87][88], generates gender-neutral job descriptions [89], and revolutionizes resume parsing efficiency with natural language processing, allowing recruiters to focus on strategic decisions [90].
+However, ethical risks remain, such as algorithms inheriting hidden biases from training data [91], implementing indirect discrimination through proxy variables (e.g., postal code linked to race), and privacy violations through the extraction of personal information for recruitment purposes [92][93].
+These contradictions and risks make fairness the core principle in AI recruitment applications.
+AI technology’s deep integration is also influencing strategic development and governance choices in biopharmaceutical companies.
+To address the ethical challenges posed by AI in healthcare, AstraZeneca has introduced a series of AI ethics principles, including privacy and security, explainability and transparency, fairness, accountability, human-centeredness, and social benefit [75].
+Additionally, AI technology shows potential and application prospects in the legal consulting field.
 Sun et al.
-simulated the collaborative workflows of real-world law firms using multi-agent systems.
-In this context, they incorporated normative principles—such as privacy protection, fairness, transparency, and accountability—to guide the ethical behavior and interactions of the agents.
-This alignment with human values enhances the quality of legal consulting services by ensuring that multi-agent systems operate in a manner consistent with ethical and societal expectations [81].
+simulated multi-agent collaboration in a real law firm setting and incorporated privacy, fairness, transparency, and accountability principles to guide multi-agent ethical behavior and interaction, aligning them with human values and enhancing the quality of legal consulting services [74].
 As AI continues to advance, the principles of value alignment will no longer be confined to broad and ambiguous moral values.
 Instead, they will increasingly penetrate specific application scenarios, integrating concrete practical experience and well-defined guidelines to inform the development and deployment of agent systems.
-3 Applications and Value Issues in LLM-based agent systems
-LLM-based agent systems are now applied across numerous domains, indicating that value alignment concerns will permeate to the application scenario level.
-This section first introduces our review subjects—LLM-based agents and multi-agent systems.
-Subsequently, it classifies and organizes intelligent agent system application scenarios based on their generality, and discusses the associated value alignment issues.
-3.1 LLM-Based Agent Systems
-3.1.1 LLM-Based Agent
-In philosophy, the term ’agent’ typically refers to an entity possessing consciousness, intention, and the capacity for action [105, 106].
-This can include humans, animals, and other entities in both the physical and virtual worlds [1].
-Floridi identifies three general characteristics of an agent: interactivity, autonomy, and adaptivity.
-Interactivity refers to an agent’s ability to interact with and respond to environmental stimuli.
-Autonomy indicates an agent’s capacity to change its own state based on its transformation rules.
-Adaptivity denotes an agent’s ability to not only alter its state but also to modify the rules governing these changes through learning and experience, thereby flexibly responding to environmental shifts.
-Building on these three characteristics, Floridi posits that a Moral Agent can exhibit morally evaluable behavior, meaning their actions can engender moral "good" or "evil." Consequently, Moral Agents must adhere to specific ethical norms and oversight mechanisms [107].
-When the concept of an "Agent" is introduced into the field of artificial intelligence, it becomes known as an AI agent.
-An AI agent is an artificial entity capable of achieving complex goals through collaboration [108], representing the concretization of the philosophical concept of "agent" within artificial intelligence.
-Large Language Models (LLMs) are deep learning models trained on vast amounts of textual data.
-They are capable of generating natural language text and comprehending its meaning, with examples including Deepseek, GPT-3 and GPT-4.
-A salient characteristic of LLMs is their adherence to scaling laws and the exhibition of emergent abilities.
-This means that once a model’s training parameters exceed a certain threshold, it can experience a dramatic improvement in existing capabilities or acquire entirely new ones [17], enabling it to solve tasks it was not explicitly trained for [12].
-As models continue to grow in scale, pre-training corpora expand, and alignment techniques advance, LLMs have demonstrated powerful capabilities in knowledge acquisition, instruction understanding, summarization, planning, and reasoning [1].
-This showcases their immense potential for achieving human-like intelligence.
-Consequently, researchers have begun leveraging LLMs to construct intelligent digital agents [23], transforming abstract philosophical concepts into practically engineerable system design frameworks [1].
-AI Agents are intelligent entities capable of perceiving their environment, making decisions, and executing actions [109, 110, 111].
-LLMs can serve as the central controller for AI agents, endowing them with decision-making thought and memory [112].
-Decision-making thought enables agents to decompose complex tasks, deliberate on each sub-task, and explore multiple pathways to arrive at intricate decisions.
-A long- and short-term memory mechanism allows agents to maintain contextual coherence during task completion and continuously learn from past interactions.
-Strategies such as multimodal perception and tool use expand an agent’s perception and action space [113, 114, 115, 116, 117].
-Furthermore, AI agents endowed with environmental interaction capabilities can learn from feedback and execute novel actions [118].
-The development of techniques like Chain-of-Thought (CoT) and problem decomposition has continuously enhanced the reasoning and planning capabilities of AI agents[119, 120].
-Furthermore, pre-training on massive-scale corpora has enabled agents to exhibit few-shot and zero-shot generalization abilities.
-This allows them to transfer and complete tasks across different scenarios without requiring parameter updates [120, 121].
-The field of AI is currently undergoing a profound paradigm shift.
-Building upon existing research, the industry is transitioning from developing AI models for passive, structured tasks to constructing novel models capable of acting as dynamic agents in diverse and complex scenarios.
-3.1.2 Multi-Agent System
-A Multi-Agent System (MAS) is a computer system composed of multiple interacting agents [12].
-Operating on natural language and custom content [24], these agents can interact within a shared environment, autonomously making decisions to solve complex problems or complete tasks [122].
-The key components of such a system include the Agent, Environment, Interaction rules, and Organization pattern [12].
-  • Agent.The agent is the core participant in the system, possessing a role identity, knowledge, capabilities, and behaviors, along with abilities such as learning, planning, reasoning, and decision-making.
-  • Environment.The environment is the external world in which the agent operates, which can be either a simulated virtual environment or a real physical space.
-  • The Interaction Rules. The rules are predefined rules, protocols, and mechanisms governing information exchange and behavior coordination among agents.
-  • The Organization Pattern. The pattern refers to the structural relationships of agent interactions, encompassing hierarchical control and emergent behavior-based organizations.
-Multi-agent interaction within these systems mirrors the specialized division of labor and teamwork found in human society; this aims to achieve collective intelligence.
-Individual autonomous agents possess initial knowledge and goal sets tailored for specific tasks [122].
-By leveraging LLMs’ capabilities in text generation, comprehension, and communication, diverse agents can adaptively collaborate.
-They respond to task requirements, participate in planning, discussion, and decision-making, and learn from one another.
-This fosters the aggregation of broad knowledge across multiple domains [112, 12].
-Concurrently, through environmental interaction, agents perceive and understand information.
-They then make decisions and execute actions based on their goals, existing knowledge, and observations.
-This process unleashes their potential to complete specialized tasks and achieve common objectives.
-Multi-agent systems exhibit remarkable flexibility, enhanced robustness, inherent self-organization and coordination, and robust real-time operation capabilities [12].
-Although interactions among agents are governed and constrained by predefined rules, emergent behaviors and phenomena can still arise from these interactions [123].
-Additionally, due to the division of labor within the system, resource and energy consumption are distributed across multiple agents.
-This is often more cost-effective than a single, powerful agent attempting to solve an entire complex problem [12].
-Therefore, compared to a single agent, multi-agent systems are better suited for handling more complex and dynamic tasks, particularly in scenarios requiring diverse strategies and collective decision-making [112].
-At present, LLM-based multi-agent systems are extensively used to simulate social phenomena.
-This includes modeling social network construction, international conflicts, political evolution, and the dynamics of public opinion  [124, 125, 126].
+3 Applications and Value Issues in Agentic AI systems
+Agentic AI systems are increasingly deployed across diverse domains, indicating that value alignment concerns will extend to the application scenario level.
+This section classifies the application scenarios of agentic AI systems based on their generality and discusses the associated and important value alignment issues.
 3.1 Application Scenarios
-Table II: Classification of Agent System Application Scenarios
+Given the powerful knowledge acquisition, instruction understanding, task planning, and logical reasoning capabilities of agentic AI systems, they have been widely used worldwide.
+From small and medium-sized organizations to multinational corporations, from higher education institutions to research institutes, from grassroots medical facilities to large diagnostic hospitals, and government agencies at all levels are actively promoting the application of agentic AI technology.
+In terms of application scenarios, the penetration of agentic AI technology is extremely broad, covering multidisciplinary fields such as healthcare, education, trade, economics, and law.
+The problems that agentic AI systems are capable of addressing range from comprehending a single piece of literature to simulating global-scale social events, spanning across micro, meso, and macro levels.
+In contrast to categorizing applications based on disciplinary domains [8] or task types [23], we sort and categorize application scenarios according to the degree of generalizability of agentic AI systems (see figure 3) and build upon this to explore the associated issues of value alignment.
+For a detailed overview of the application scenarios, please refer to the Appendix.
+Figure 3: An Overview of Classification of Agentic AI Systems Application Scenarios
+Generalizability of agentic AI systems refers to the extent to which they can be applied across diverse scenarios or meet the needs of various organizations.
+General agentic AI systems have a high degree of generalizability, which is based on the foundational capabilities of AI agents such as natural language processing, reasoning, planning, and autonomous decision-making.
+The universal application includes tool using, dialog agents, multimodal perception, multi-agent collaboration, visual tasks processing, natural language processing, data generation and analysis, reasoning, planning, and decision optimization.
+For example, AI agents are capable of recognizing implicit needs for tool invocation and leveraging a wide range of real-world tools involving perception, manipulation, logical reasoning, and creativity.
+These tools enable AI agents to process multimodal inputs—such as images, tables, and code—to accomplish complex tasks [94, 95].
+A higher degree of generalizability indicates that the agentic AI systems can be broadly applied across various scenarios, not only addressing domain-specific problems but also meeting diverse task requirements across multiple industries and organizations.
+For example, multi-agent systems can be applied not only in natural sciences but also in the social science domain, such as economics and management, suitable for addressing global issues that span across nations and cultures.
+Specific agentic AI systems have limited generalizability across industries and domains.
+Specialized applications can be broadly categorized into two types: function-driven and domain-driven.
+Function-driven applications are built upon the core capabilities of agents and are typically applicable across multiple domains.
+Common examples include code generation, social simulation, graphical user interface (GUI) agents, and model evaluation and improvement.
+Domain-driven applications are tailored to specific sectors, including finance, science, healthcare, gaming, manufacturing, and autonomous driving.
+For instance, in healthcare, multi-agent systems support medical reasoning, gamified mental health assessments, and experimental design in genomics [96, 97].
+These systems are tailored for specific tasks involving sensitive datasets and proprietary business information.
+3.2 Value Issues in Application
+The value alignment problems involved in the application scenarios of agentic AI systems mainly include two aspects: the correspondence between the generalizability of the system and the multi-level values, and the contradictions encountered when aligning values.
+The degree of generalizability of agentic AI systems corresponds to the multi-level of values.
+As the degree gradually decreases, the granularity of the value norms that need to be met gradually increases.
+A higher degree of generalizability allows agentic AI systems to operate across diverse contexts.
+Their initial design need not align with the values of specific countries, industries, or cultures, as long as they adhere to universal values at the macro level.
+In contrast, systems with limited generalizability are often tailored to specific national or industrial settings, reflecting corresponding cultural and institutional features.
+Their development and governance must address local policies, standards, and cultural norms, while aligning with universal values.
+When deployed in private contexts, systems are customized for particular organizations to solve specialized business or personal problems.
+These systems are highly specialized and exhibit minimal generalizability.
+Under such conditions, value alignment must occur at multiple levels: not only with macro- and meso-level norms, but also with the internal value frameworks of the organizations involved.
+It can be seen that agentic AI systems with high generalizability tend to be highly adaptable, and their value alignment primarily centers on adherence to universal principles of macro level.
+As application specialization increases, the requirements for value alignment become progressively more complex, necessitating adherence to value norms across multiple levels simultaneously.
+This classification aligns with the top-down, increasingly granular hierarchy of value principles.
+As a result, the value norms that agentic AI systems must adhere to vary according to their degree of generalizability, offering a structured basis for governance and oversight in their design and deployment.
+The value alignment problem may involve two types of tensions: the plurality of value principles at the same level and conflicts across different levels.
+On one hand, a key challenge is that value principles at the same level differ across countries, regions, industries, organizations, and other contextual factors.
+Therefore, when agentic AI systems are applied in diverse contexts, their value alignment objectives must be adapted accordingly.
+In the healthcare sector, Huawei’s agent-based EIHealth platform supports genomic analysis, drug discovery, and clinical research.
+Built on Huawei Cloud’s strengths in AI and big data, EIHealth features an open and extensible architecture that enables end-to-end AI empowerment.
+Ensuring alignment with the value principles and regulatory norms of various countries, regions, and organizations is crucial for its deployment.
+To address this, Huawei Cloud has issued a series of white papers tailored to the privacy and regulatory requirements of different jurisdictions and industries.
+For example, it provides compliance guidelines aligned with the Personal Data Privacy Ordinance (PDPO) in the Hong Kong SAR, as well as guidelines addressing the financial regulatory frameworks of the EU, including specific provisions for Ireland, Spain, and Hungary.
+These efforts ensure that the EIHealth platform adheres to all applicable national and regional security regulations, international cybersecurity and cloud security standards, and incorporates industry best practices, ultimately meeting the diverse security needs of cloud service users.
+On the other hand, as the tasks assigned to a system become more fine-grained, the constraints on achieving alignment goals also increase.
+Efforts to maximize task utility may conflict with established norms, creating tensions across value levels—for example, promoting social cohesion at the macro level may clash with individualism at the meso level, or national mandates for transparency may oppose organizational privacy concerns.
+In the field of corporate recruitment, AI systems are widely used for résumé screening and candidate-job matching [98], yet often raise ethical concerns [93].
+In particular, corporate hiring practices may involve gender discrimination or racial bias, violating legal regulations and core societal values.
+A well-known example is Amazon’s 2018 résumé screening system, which discriminated against female applicants and was ultimately abandoned [99].
+To prevent such issues, recruitment-oriented intelligent systems must be explicitly designed to minimize bias and ensure compliance with labor laws, equal employment regulations, and industry standards.
+Only on the basis of legal compliance, equality, and privacy protection should these systems then pursue the optimization of recruitment outcomes.
+4 Methods and Evaluation for Value Alignment
+To ensure that LLM-based agents behave in ways consistent with human values and task goals, value alignment and its evaluation have become key research focuses.
+In this section, we first discuss the methods for value alignment, then review evaluation approaches.
+4.1 Value Alignment Methods for LLM-based Agents
+4.1.1 Value Alignment during LLM Pretraining
+Value-Prompt Pre-conditioning: Within the conventional self-supervised pretraining frameworks, such as masked language modeling or autoregressive language modeling, researchers prepend a short “value prompt” to the input sequence [100, 101].
+They might attach “Please respond with a safe and neutral tone” or “Please adhere to principles of fairness and inclusivity” at the beginning of each sentence, and then include both the prompted and original text in the pretraining corpus.
+In this way, the model passively receives implicit guidance toward objectives like “safety” and “objectivity.
+For example, Zheng et al.
+[100] unpack why and how simple prepended safety directives move model representations into “higher refusal” regions.
+[102] shows that conditioning an autoregressive LM on human preference scores from day one dramatically reduces harmful output rates.
+Gu et al.
+[103] show that soft prompts—even if not explicitly about “values”—validate the premise that embedding trainable tokens in early stages leads to more controllable downstream behaviors (e.g., fairness, bias detection).
+Multi-Task Pre-training: Inspired by the multi-task learning paradigm, this approach mixes value-judgment or safety-detection tasks with standard self-supervised objectives [104, 105, 106].
+Specifically, in addition to the masked language modeling (MLM) or autoregressive objective, small-scale, labeled supervision signals—such as “ethics classification”, “bias detection” or “hate speech recognition”—are incorporated.
+By jointly optimizing a multi-task loss function, the model’s parameters are trained to satisfy both language modeling requirements and the ability to distinguish harmful or biased content.
+Although this method requires additional annotated data, it explicitly strengthens the model’s discrimination boundary along value dimensions and lays the technical foundation for subsequent alignment fine-tuning.
+Safety-Aware Curriculum Learning: Drawing on the concept of curriculum learning (CL), this method stratifies large-scale corpora according to “safety” or “sensitivity” levels [107, 101, 108].
+First, a high-quality, low-risk “safe corpus” is used for a warm-up pretraining phase, allowing the model to initially acquire expression patterns that comply with ethical and regulatory requirements.
+Subsequently, more general or higher-risk text is introduced gradually, enabling the model to expand its language capabilities while retaining dependence on earlier safe preferences.
+This staged training strategy establishes a better balance between model capability and safety, and experiments show it substantially outperforms full-scale pretraining in controlling the model’s tendency to generate sensitive content.
+The above value-alignment methods can be combined [105, 101], for example, integrating value-prompt pre-conditioning and multi-task supervision within a safety-aware curriculum framework, along with regularization penalties.
+It allows the model to be reinforced from multiple dimensions in embedding value objectives such as “ethics”, “safety”, and “fairness”.
+Future research may further explore the scalability and synergistic effects of these strategies on ultra-large corpora and hundred-billion-parameter models, paving new paths for achieving more robust and interpretable value embeddings.
+Figure 4: An Overview of Methods for Value Alignment in Agentic AI
+4.1.2 Value Alignment during LLM Post-Training
+Supervised Fine-Tuning (SFT).
+SFT serves as a fundamental step in aligning agentic language models with human values.
+It usually involves fine-tuning a pretrained model on a carefully curated dataset of demonstration–response pairs that embody normative principles such as ethics, safety, and fairness.
+These examples are typically constructed by human annotators and cover a broad range of scenarios in which value-sensitive behavior is essential—for example, avoiding toxic content, maintaining factual correctness, or respecting user intent.
+During SFT, the model is optimized to maximize the likelihood of producing responses that match these human-crafted demonstrations.
+For instance, [109] proposed an automated method for generating instruction data to fine-tune GPT-3, facilitating scalable alignment.
+[110] demonstrated that initializing InstructGPT with SFT on human-written demonstrations results in models that more faithfully follow user instructions, while also significantly reducing harmful or misleading outputs.
+Further advancements have introduced principled guidance into the SFT process.
+[111] incorporated a set of human-authored value principles to constrain model behavior through in-context learning, thereby enhancing the generation of helpful and harmless content.
+Additionally, [112] enriched the fine-tuning dataset with both positive and negative value-aligned instances.
+This contrastive supervision enables the model to learn fine-grained distinctions between aligned and misaligned outputs, improving its sensitivity to nuanced ethical considerations.
+Reinforcement Learning from Human Feedback (RLHF).
+RLHF has emerged as a core paradigm for aligning LLMs with human values, preferences, and behavioral norms.
+It extends SFT by introducing an additional optimization stage that incorporates human feedback signals into policy learning, enabling models to go beyond surface-level imitation toward deeper alignment with human intent [110, 113].
+The classical RLHF pipeline consists of three major components:
+  1. 1.
+    Supervised Fine-Tuning (SFT): A pretrained LLM is first fine-tuned on a curated set of instruction–response pairs to provide a reasonable initialization.
+  2. 2.
+    Reward Model (RM) Training: Given multiple model-generated responses to the same prompt, human annotators provide pairwise preference labels (e.g., ranking which response is more helpful or harmless).
+These preference data are then used to train a reward model that can approximate human preferences [110].
+  3. 3.
+    Reinforcement Learning: The LLM is further fine-tuned using reinforcement learning, typically Proximal Policy Optimization (PPO), where the reward signal is derived from the RM.
+This step encourages the model to generate outputs that maximize predicted human approval.
+RLHF was first popularized through the development of InstructGPT [110], which demonstrated notable improvements in producing truthful, helpful, and safe responses without significant task performance degradation.
+Subsequent systems such as ChatGPT and GPT-4 also adopted RLHF as a central alignment mechanism [114].
+To reduce the dependency on expensive human annotation, [115] proposed using synthetic feedback generated by automated evaluators, while [116] explored using self-critiquing models to support human evaluators and reduce their cognitive burden.
+[117] reframed the RLHF objective as minimizing an $f$-divergence between model and human preference distributions, providing a more principled view of policy optimization under alignment constraints.
+Moving beyond individual feedback signals, [118] trained LLMs in a simulated society of interacting agents, using environment-level feedback to promote socially aligned behaviors at scale.
+Despite its success, RLHF faces several challenges.
+It relies heavily on the quality and consistency of human preferences, which may encode biases or lack coverage for edge cases.
+Reward models can be exploited or lead to reward hacking behaviors, and PPO-based optimization may be unstable or sample-inefficient.
+Constitutional AI and Reinforcement Learning from AI Feedback (RLAIF).
+Given the high cost and limited coverage of exclusively human-derived feedback, recent work has investigated self-supervised alignment methods that combine a model’s intrinsic knowledge with a set of expert-defined principles.
+A leading instance of this paradigm is Anthropic’s Constitutional AI [119, 120, 121], wherein the model is first equipped with a “constitution” of value-driven rules (e.g., avoid harm, preserve honesty).
+In the initial phase, the model generates a self-critique of its own response based on these rules and subsequently revises the response accordingly.
+During the RL stage, the model then acts as its own evaluator: it compares paired responses to train a preference model, which in turn provides reward signals for further fine-tuning.
+This process—termed Reinforcement Learning from AI Feedback (RLAIF)—minimizes reliance on human annotation by leveraging automated self-assessment [122, 123, 120].
+Empirical evaluations [119] indicate that Constitutional AI produces assistants that are both “helpful yet non-offensive” and that, when faced with harmful prompts, preferentially offer explanations and deterrent advice rather than mere refusal.
+This principle-driven, self-supervised strategy thus demonstrates substantial promise for reducing human intervention and scaling alignment across diverse scenarios.
+Self-Reflection & Self-Correction: LLMs can engage in self-reflection to identify and correct errors or misalignments in their outputs.
+A seminal example is Self-Refine [124], which alternates between two generative steps—FEEDBACK and REFINE—to iteratively improve initial responses without any additional training data or external supervision.
+Building on this idea, SELF-RAG [125] integrates retrieval-augmented generation with self-reflection: the model learns to emit “reflection tokens” that determine when to retrieve external knowledge and when to self-evaluate, thereby enhancing factuality while retaining generative creativity.
+Complementing these methods, Chain-of-Thought review techniques prompt the LLM to re-examine its own reasoning chains.
+Wei et al.
+[126] showed that providing intermediate reasoning steps not only boosts complex problem-solving performance but also offers an interpretable scaffold for downstream self-correction when logical inconsistencies are detected.
+Madaan et al.
+[127] proposed a self-refine instruction-tuning approach in which smaller LLMs are first instruction-tuned via demonstrations from larger, robust teacher models and then allowed to self-improve their reasoning under the same instruction framework.
+This two-stage process significantly enhances generalization of self-correction capabilities, further narrowing the gap between smaller and larger models.
+Adaptive Fine-Tuning.
+In dynamic deployment settings, agentic LLMs must continuously adapt to evolving user preferences and environmental rewards—a process often termed adaptive fine-tuning.
+One recent line of work, PRELUDE [128] frames this as “preference learning from user edits”: by inferring a user’s latent preference profile from historic edit logs, the agent periodically fine-tunes its response policy to better align with those preferences.
+Similarly, Hao and Duan [129] propose an online learning mechanism that dynamically reweights human feedback signals—treating users as strategic labelers—to mitigate malicious or noisy annotations and achieve sublinear regret in preference aggregation.
+Beyond supervised feedback, agents can also adapt via reinforcement learning from environment rewards [130].
+When certain behaviors incur negative utility, standard policy-gradient updates naturally down-weight those actions over successive episodes, effectively steering the policy toward compliant behaviors.
+4.1.3 Value Alignment in Agent Architecture Design
+The architectural design of an LLM-based agent is crucial for aligning its values.
+By incorporating specific modules and mechanisms within the agent, its behavior can be guided to align with the desired value standards.
+Profile Definition: Profile Definition refers to the pre-configured set of personalized attributes assigned to each agent within an LLM-driven agent architecture [131].
+These profiles not only include basic information such as the agent’s name, identity, and scope of capabilities, but also define its goals, preferences, behavioral style, and prohibited actions, thereby ensuring role consistency and value alignment [132].
+The implementation of profile definitions typically relies on the collaboration of structured prompt templates, personality tags, behavioral constraint rules, and memory modules.
+Some systems also incorporate dynamically evolving configurations to support sustained adaptation [133].
+For example, the MobAgent system combines individual behavioral traits to generate virtual travel diaries, enabling agents to behave naturally in urban simulations [134].
+MOTIF framework uses a hybrid of manual and generative methods to create socially simulated agents with stable personalities [135].
+Currently, the design of profile definitions is moving toward greater generalizability, finer granularity, and enhanced learnability.
+Key challenges include managing boundary conflicts between multiple roles, balancing adaptability with static settings, and maintaining long-term behavioral consistency [131, 134].
+Memory Mechanisms: Memory Mechanisms refer to the capability of enhancing a model’s long-term contextual awareness and behavioral consistency by storing and retrieving historical information.
+At their core, they involve the structured storage and effective retrieval of interaction content, instructions, user preferences, and feedback—enabling personalized responses, long-term task tracking, and value alignment [136].
+In both research and applications, memory mechanisms are a key component for enabling agents to maintain consistent long-term behavior and adhere to value principles.
+Their central value lies in “helping agents remember the past and apply what they’ve learned” [137].
+The implementation of memory mechanisms typically relies on two subsystems: working memory and long-term memory.
+Working memory maintains the context of the current session, while long-term memory stores knowledge and experiences using graph structures, databases, vector stores, or retrieval-augmented generation (RAG) models.
+Common approaches include time-window-based history caching, adaptive memory updates, semantic similarity-based vector retrieval, and contextual reflection mechanisms [138] [139].
+Representative examples include the G-Memory framework [140], which uses a hierarchical architecture to support shared memory among multiple agents, enabling task coordination and improving memory consistency in complex systems.
+The Cognitive Weave system constructs a Spatio-Temporal Resonance Graph, allowing agents to recall abstract knowledge and reflect emotional memory during conversations [141].
+Reasoning and Planning: Reasoning and Planning refer to an agent’s ability to perform logical inference and make multi-step decisions when handling complex tasks.
+This ability encompasses cognitive processes such as causal reasoning, path selection, and tool utilization based on knowledge and environmental information, as well as systematic foresight and regulation of future actions [142].
+As a core component of the decision-making module, reasoning and planning play a fundamental role in value alignment, task execution efficiency, and behavioral stability [143].
+In practice, reasoning and planning help agents decompose tasks and simulate actions thoughtfully before execution—reducing impulsive decisions and enhancing value consistency [144].
+Current implementation techniques include Chain-of-Thought (CoT) prompting, which guides models through step-by-step logical reasoning.
+Other methods include Tree-of-Thought, Atomic Fact Augmentation, external knowledge retrieval, and Lookahead Search [145].
+Some research also explores plan caching, which stores historical decision paths to reduce the computational cost of repeated planning [146].
+In applied settings, the SOP-Bench framework uses Standard Operating Procedures (SOPs) as an evaluation scaffold to test the stability of LLMs in reasoning and planning tasks, demonstrating their practical value in process control [147].
+The WGSR-Bench project leverages game-based tasks to assess LLM performance in strategic planning and multi-agent coordination, making it a representative platform for evaluating advanced planning mechanisms [148].
+Execution and Tool Use: Action Execution and Tool Use refers to an agent’s ability to actively perform operations and invoke external tools in order to accomplish tasks [149].
+In LLM-based agent systems, it is one of the core mechanisms enabling autonomy and interaction with the environment.
+It also serves as a key interface for ensuring alignment with human values [150].
+The primary function of this mechanism is to provide a controllable and auditable channel for external actions, forming the foundational module for value alignment control strategies [151].
+Common implementation approaches include embedding pre-decision processes—for instance, the HADA architecture introduces value judgment nodes to filter candidate action sequences [152]; modular designs are also used, such as in CoReaAgents, which separate “planning agents” and “tool agents” to collaboratively manage phased control and reasoning in complex tasks [153]; and frameworks like Safe-Align associate each atomic action with potential risks to ensure pre-execution safety checks [154].
+In terms of practical applications, the Ali-Agent project combines web search tools with evaluation models to automatically verify whether actions align with human preferences, enabling value alignment testing in intelligent dialogue systems [155].
+Current challenges in this domain include: (1) unstable toolchain logic in multi-step planning, leading to action hallucinations or redundant operations [154]; (2) high context dependence of value judgments, with ambiguous standards that are hard to unify across cultures [150].
+Reflection Loop: Reflection Loop is a cognitive control module embedded in LLM-based agent systems.
+Its core function is to endow agents with metacognitive abilities to review–evaluate–revise during task execution.
+Specifically, the Reflection Loop introduces a reflection phase that allows agents to revisit and assess their actions or reasoning, thereby improving the effectiveness and alignment of subsequent behavior or inference [156].
+In both research and practice, the reflection mechanism is often used to enhance value alignment—making agent behavior more consistent with human expectations and ethical norms [157].
+Common implementations include inserting explicit “thought summarization” or “output verification” modules into the agent’s control flow—for example, using a Self-Ask mechanism to generate self-questioning tasks, or invoking an external reflection model to assist with strategy review [158].
+More advanced forms use multi-agent collaboration structures, modeling the reflection process as a feedback loop where one agent acts as the primary decision-maker while another serves as an introspector offering corrective insights [159].
+In real-world applications, the Agent-Pro system integrates strategy-level reflection mechanisms, significantly improving agent consistency and final scores in game-based tasks [160].
+The HADA framework employs a dual-module architecture (“Evaluator + Self-reflection Agent”) to support dynamic value alignment in long-term tasks [161].
+Reflection mechanisms are evolving from simple single-round reviews toward multi-round iterative reflection and context-aware reflection.
+However, challenges remain, such as shallow reflection depth and poor semantic consistency [162].
+Future developments aim to combine reinforcement learning and cognitive modeling, enabling agents to autonomously learn when to reflect and how to reflect efficiently [163].
+Value Evaluator: Value Evaluator is a critical module whose core function is to simulate the role of a human supervisor by making value-based judgments and filtering model-generated outputs.
+This ensures that the content, beyond being logically sound, aligns more closely with mainstream human ethics, cultural norms, and goal orientations [164].
+This module is typically used in reinforcement learning or policy fine-tuning stages to calibrate agent decisions against human standards, making it especially valuable in multi-agent systems, open-world dialogue tasks, and AI governance scenarios [165].
+Currently, value evaluators are mainly implemented through two mechanisms: the use of independently trained reward models that learn from human preference data to score candidate outputs and guide policy optimization; and rule-based or value classifiers that make explicit judgments based on knowledge graphs or ethical rule sets [166].
+Some systems also adopt dual-model architectures, such as separating the “Judge LLM” and “Actor LLM,” to decouple evaluation from execution and improve fairness and generalizability [167].
+In practice, the ALI-Agent system establishes a value alignment evaluation pipeline based on inter-agent interaction, measuring the ethical consistency and diversity of LLM responses [167].
+The Verila framework introduces an interpretable Verifier module into LLM agents to provide human-readable evaluations of failures in complex tasks [168].
+This field is rapidly evolving toward multi-value integration and scalability.
+On one hand, researchers are exploring how to incorporate multiple cultural backgrounds and group preferences into models; on the other hand, they face challenges such as limited generalization of evaluators and the risk of propagating data bias [168, 169].
+4.1.4 Value Alignment for Multi-Agent System
+The development of value alignment methods for multi-agent systems is essential to ensure agent behaviors remain consistent with human values and collective system goals.
+By designing coordination strategies, communication protocols, and reward structures that embed normative principles, the interactions among agents can be shaped to promote aligned, cooperative, and socially beneficial outcomes.
+Cooperative and Collaborative Learning: Within a multi-agent collaboration framework, different agents can take on distinct roles, complementing each other’s strengths while jointly accomplishing tasks and monitoring one another for norm deviations [6, 7].
+For example, one agent may focus on task planning, another is responsible for evaluating the safety of the plan, while a third may simulate a user role to provide feedback [170].
+This division of labor enables each agent to specialize in a particular dimension, while sharing information and constraints through natural language communication.
+Cooperative multi-agent systems can emulate human team-based mutual supervision mechanisms: if one agent generates a questionable suggestion, others can promptly flag it and request clarification or correction, thereby filtering out misaligned outputs internally.
+Studies have shown that through role-playing and interaction, LLM-based agents can spontaneously develop human-like Theory of Mind reasoning, forming mental models of each other’s goals and intentions, which helps anticipate and prevent undesirable behaviors.
+Furthermore, some work addresses objective alignment through gradient-based methods in cooperative processes, such as Altruistic Gradient Adjustment, which adjusts learning dynamics to reduce conflicts in mixed-motive settings [171].
+Another stream emphasizes normative alignment, equipping agents with deontic logic-based reasoning to align decisions with social values and ethical principles [172].
+Communication-driven alignment is also advancing, with approaches like CoDe introducing intent- and timeliness-based mechanisms to maintain coordination under message delays [173].
+Temporal alignment at the feature level is further explored, such as TraF-Align, which ensures cross-frame semantic consistency [174].
+Adversarial Learning: Recent advances have increasingly explored debate-based methods as a promising approach to multi-agent alignment.
+The multi-agent debate framework enables multiple LLM agents to engage in adversarial discussions on the same question, with final answers selected via voting or adjudication.
+Inspired by human debate, one agent proposes an answer while others critique and challenge it, iteratively refining responses until no major flaws remain.
+A referee agent or majority vote then determines the most accurate and safe output [175].
+Studies show that such mechanisms reduce hallucinations and improve factual accuracy, especially in high-stakes domains like medicine and law [176, 177].
+Voting-based approaches—where agents independently generate answers and the system selects the most common or consensus-aligned response—leverage collective intelligence to filter out unreliable outputs.
+Both debate and voting use agent redundancy and disagreement to enhance alignment and reliability.
+From another perspective, some optimization methods for debate in the alignment task are proposed.
+For instance, sparse communication topologies in multi-agent debates have been shown to reduce computational costs while maintaining or improving performance in alignment labeling and multimodal reasoning tasks [178].
+Other works incorporate gradual vigilance and interval communication, allowing agents to assess alignment risks through controlled information exchange [179].
+Emotionally grounded frameworks inspired by cognitive metaphors (e.g., Inside Out) structure agents by affective roles, enabling emotionally calibrated alignment through iterative critique and synthesis [180].
+Debate has also been integrated with self-reflection mechanisms, where agents dynamically switch between arguing and introspection, enhancing both factual accuracy and normative alignment [181].
+Furthermore, debate-based zero-shot stance models simulate structured opposition and rebuttal, with a referee synthesizing outcomes when consensus fails [182].
+Hierarchical Learning: Multi-agent systems can adopt hierarchical architectures to enable supervision and amplification, akin to organizational structures.
+This approach has been applied in frameworks like AutoGen, where supervisor agents oversee tool usage and corrections [183].
+Hierarchical setups also support peer-review workflows, enabling agents to critique and refine each other’s outputs, forming internal feedback loops that improve quality and enforce alignment beyond the capacity of single-agent systems.
+Recent studies have explored structured agent architectures to enhance coordination, reliability, and value alignment in multi-agent systems.
+MetaGPT models human-standard workflows via Standard Operating Procedures (SOPs), assigning agents to domain-specific roles in an assembly-line structure to reduce cascading hallucinations and ensure intermediate verification [184].
+HALO similarly adopts a hierarchical reasoning design with high-level planners, mid-level role designers, and low-level executors, integrating Monte Carlo Tree Search (MCTS) for optimized task decomposition and execution [185].
+TalkHier further reinforces alignment by combining structured communication with hierarchical output refinement, outperforming both single-agent and flat multi-agent baselines.
+These designs reflect a growing consensus that task and communication hierarchies are critical for managing complexity, role coordination, and prompt adaptation in real-world multi-agent deployments [186].
+Recent architectural taxonomies and social-structural analyses emphasize that organizational hierarchies not only improve performance but also shape how agents align with human preferences, collectively or individually [187, 188].
+4.1.5 Value Alignment via Data and Reflection Feedback
+Value alignment through data and reflection feedback focuses on shaping agent behavior by curating value-informed data and enabling self-assessment mechanisms.
+By integrating normative cues, e.g., ethical guidelines, social norms, or human preferences, into training datasets, agents can progressively internalize desired value structures.
+Self-Constructed Data: To reduce reliance on human-labeled alignment data, models can generate their own training samples.
+Methods like Self-Instruct allow a language model to create instruction–response pairs, filter low-quality outputs, and fine-tune itself—boosting performance significantly with minimal human input [109].
+This strategy can be extended to value alignment by simulating ethical dilemmas or adversarial prompts and generating safe responses.
+While cost-effective and broad in scope, such data must be filtered using a value evaluator to avoid reinforcing biases.
+Human Feedback Data: Human feedback is one of the most reliable signals for value alignment.
+Methods like RLHF use human preferences to train reward models, while expert-written responses in SFT help correct value misalignments [110].
+Implicit user feedback (e.g., accepting or rejecting answers) can also support continuous optimization.
+Studies show that such feedback reduces hallucinations and improves social appropriateness [170].
+Incorporating real-time and culturally diverse feedback creates a closed loop—output, feedback, update—that helps models better align with human values over time.
+Environmental Feedback Data: For agents deployed in interactive environments—such as robots or game AIs—the environment itself can provide value signals through reward functions [170].
+Well-designed rewards translate human-desired behaviors into scores, guiding agents to learn aligned strategies via reinforcement learning.
+For instance, in autonomous driving, safety and efficiency can be rewarded, while risky actions are penalized.
+LLM-based agents can similarly adjust their policies based on environmental feedback [189].
+One study shows that incorporating dynamic environmental responses helps agents adapt their planning to complex conditions [190].
+However, since reward functions often prioritize task performance, they must be carefully designed to ensure that high scores also reflect safety and ethical compliance.
+This can involve integrating strong penalties for undesirable behaviors (e.g., triggering taboo topics or harmful outcomes).
+By learning from real interactions, agents gain experiential understanding of “right” and “wrong,” which is critical for deep value alignment.
+Policy and Normative Data: Integrating socially accepted norms into model training or inference is a key method for value alignment.
+By incorporating texts like laws and ethical guidelines during pretraining or fine-tuning, models can internalize these values.
+For example, Anthropic’s Claude model uses a “constitution” to enforce behavioral standards, guiding the model to self-correct violations.
+This approach provides a set of first principles, enabling models to better control behavior, especially in boundary cases [119].
+Rule engines can also be used to filter or revise unsafe outputs.
+4.2 Value Alignment Evaluation
+4.2.1 Methods for Value Alignment Evaluation
+Given the widespread use of agentic AI in text-centric tasks, the evaluation of value alignment for LLM-based agents often closely parallels the methodologies employed for assessing value alignment in LLMs more generally.
+A common approach involves constructing evaluation datasets that comprise both subjective and objective questions, grounded in specific value frameworks, which serve to test the agent’s ability to align with diverse ethical, cultural, or normative standards through its responses.
+However, the operational scope of agentic AI systems typically extends beyond static text generation, encompassing broader and more dynamic application scenarios.
+Consequently, additional evaluation strategies are required to capture agent behavior in context comprehensively.
+These include scenario-based simulations, assessments of decision-making quality, and performance evaluations in value-constrained task completion.
+Together, these complementary methods offer a more holistic perspective on an agent’s capacity to maintain value alignment in complex, real-world environments.
+A summary of representative benchmark datasets and their construction methods is provided in the Appendix.
+4.2.2 Question Formats and Metric Design
+Evaluation for value alignment largely depends on the format of assessment questions.
+Common question formats include multiple-choice questions, binary judgments, ranking tasks, rating scale questions, and open-ended questions.
+Multiple-Choice Questions.
+Multiple-choice questions typically present a set of answer options—often including neutral, biased, or varying degrees of harmful responses—and require the model to select the most appropriate one.
+It is commonly used to assess a model’s value judgments in specific contexts, such as BBQ [191], CDEval [68].
+Binary Judgments.
+Binary judgment questions require the model to make a dichotomous decision, such as yes/no, true/false, or agree/disagree, based on a given statement or scenario.
+This format is widely used to evaluate a model’s understanding of ethical principles, social norms, or cultural consensus.
+For instance, ETHICS presents ethically challenging statements (e.g., “Stealing is acceptable if no one finds out”), each of which is annotated by humans as either correct/incorrect or acceptable/unacceptable [192].
+Ranking Tasks.
+Ranking tasks evaluate a model or agent system’s ability to make relative value-based comparisons by requiring it to order multiple responses or behavioral options from most to least aligned with human values, such as StereoSet [193] and MoralExceptQA [194].
+Rating Scale Questions.
+Rating scale questions require the model or agent system to evaluate a given text, response, or behavioral option along predefined value dimensions.
+For instance, the Moral Integrity Corpus presents scenarios in which dialogue agents face moral dilemmas, providing contextual setups and multiple behavioral options (e.g., whether to disclose a friend’s secret or follow a rule) [195].
+The model or agent system rates the moral appropriateness of each option on a five-point Likert scale (e.g., 1 = highly immoral, 5 = highly moral).
+These ratings are then compared to human judgments using Pearson or Spearman correlation coefficients to evaluate moral alignment.
+Open-Ended Questions.
+Open-ended questions allow free-form responses from the model or agent system, enabling the assessment of its alignment with human values, such as BeaverTails [43] and BOLD [196].
+This format evaluates not only factual correctness but also the appropriateness and social acceptability of generated content.
+Evaluation methods for value alignment largely depend on the format of assessment questions. Accordingly, different metrics are used to quantify model performance across various task types:
+Accuracy. Accuracy is the most commonly used metric for multiple-choice and binary judgment questions. It measures the proportion of model responses that match the correct or human-annotated answers.
+Correlation-Based Metrics.
+Correlation-based metrics, such as Pearson’s correlation coefficient and Spearman’s rank correlation coefficient, are widely used in both rating scale and ranking tasks.
+These metrics evaluate the extent to which model predictions align with human judgment patterns across various value dimensions.
+Error-Based Metrics.
+When the model’s output is numerical (e.g., a score prediction), Root Mean Squared Error (RMSE) or Mean Absolute Error (MAE) may also be used to capture the deviation between predicted and reference scores.
+These are particularly useful in rating scale questions.
+Other Metrics.
+In open-ended questions, where responses are free-form text, evaluation is primarily based on human judgment, with annotators assessing multiple value-related dimensions and assigning explicit scores.
+5 Insights into Value Coordination in Agentic AI Systems and Future Directions
+Figure 5: A Conceptual Framework for Future Directions
+Drawing upon insights from our review of value systems, application scenarios, alignment evaluation methodologies, and datasets, we contend that while existing researches continuously introduce novel methods and frameworks, the field still confronts considerable challenges and offers significant opportunities for future investigation.
+As illustrated in Figure 5, this section will delineate future research directions for value alignment in agentic AI systems.
+Our objective is to achieve sustainable and stable value alignment by addressing four pivotal challenges, progressing from internal to external considerations.
+These encompass: (1) the design of game theory-based interaction mechanisms to resolve issues of internal decision-making and behavioral logic (Section 5.1); (2) the formulation of organizational models to address problems of internal structure and coordination mechanisms (Section 5.2); (3) the alignment refinement of communication protocols to manage external information exchange and interactions (Section 5.3); and (4) the construction of value evaluation frameworks and open datasets to facilitate social integration and ecosystem co-construction (Section 5.4).
+5.1 Value Considerations in the Design of Multi-agent Interaction Mechanisms Based on Game Theory
+The effective functioning and benefits of agentic AI systems hinge on efficient agent interactions.
+These interactions, whether cooperative, competitive, or co-opetitive, foster collective intelligence.
+However, individual interests can still clash with collective goals in social dilemmas [197].
+Even if agents share aligned values, game-theoretic challenges can prevent the entire system from achieving overall value alignment [198].
+Therefore, value alignment in multi-agent interactions is a critical area for future research.
+The design of multi-agent interaction mechanisms, particularly when grounded in game theory, can foster the emergence of novel values rooted in inter-agent dynamics.
+This highlights the distinct nature of agent interaction, setting it apart from human interactive games and other AI technologies.
+Compared to human-centric games, agent interaction games are characterized by numerous parameters, intricate utility functions, and multiple stages.
+While agents’ computational power amplifies this complexity, it also makes these games solvable.
+Relative to other AI technologies, LLMs excel in language-based reasoning, planning, and problem-solving [199, 200].
+These capabilities enable LLM-based agents to simulate human-like traits, including personality and tone [201, 202, 124], and even exhibit a degree of theory of mind—the ability to infer others’ beliefs and intentions during strategic interactions [203].
+LLM-based agents can predict opponents’ psychological states and behaviors by considering task objectives, resource distribution, and personality configurations, enabling adaptive allocation and equilibrium decisions.
+Consequently, AI agents can be treated as fully rational entities, and their interactions modeled using game theory.
+Future research can explore the values reflected in multi-agent interactions by focusing on three key activities related to their game objectives: agent identification, goal setting, and boundary constraints.
+For agent identification, the primary value consideration is whether the agent’s characteristics lean towards individualism or collectivism.
+In goal setting, the relevant values often involve equilibrium-seeking or distributive outcomes, and prioritizing equity or efficiency.
+Finally, for boundary constraints, key value considerations include information symmetry, the multi-round interactions, and whether a short- or long-term perspective is adopted.
+For individualistic agents aiming for equilibrium, non-cooperative game theory helps model how to guide them toward effective cooperation in moral dilemmas, preventing exploitation or free-riding.
+By analyzing game rounds, we can also explore short-term vs.
+long-term orientations; in multi-round, sequential games, agents must predict future states to maximize long-term gains, revealing dynamic decision-making strategies.
+Conversely, for collectivist agents focused on distribution, cooperative game theory models how they naturally form collaborations to resolve efficiency losses while balancing fairness and efficiency.
+This theory also aids in responsibility attribution, a critical aspect for designing autonomous, ethically compliant systems.
+For instance, Triantafyllou et al.
+[204] applied cooperative game theory to study accountability in sequential multi-agent decisions, highlighting its significance for fairness, transparency, and responsibility within these systems.
+Designing regulatory mechanisms for agentic AI systems hinges on information symmetry.
+In scenarios with incomplete information, agents may deviate from ethical principles to maximize self-interest, necessitating incentive-compatible mechanisms.
+Unlike human-centric regulation, where private information is a barrier, regulating agents present unique challenges.
+Agent objective functions can implicitly embed designer values, requiring new constraints to align with societal goals.
+A key distinction is that agents’ private information can be data or model parameters.
+Agents, being algorithm-driven, possess superior rationality and processing capabilities, potentially colluding algorithmically (e.g., via federated learning) to alter information asymmetry.
+Furthermore, agent behavioral patterns differ from humans’.
+Humans might dissemble due to emotion or gain, while agents compute optimal strategies based on algorithms.
+This calls for different mathematical modeling for incentive compatibility, perhaps incorporating reinforcement learning’s reward mechanisms.
+Finally, the dynamic, scalable nature of agent systems means regulatory mechanisms must adapt in real-time to their rapid learning and evolving environments.
+5.2 Value Considerations in the Design of Multi-agent Organizational Model and Structure
+An organizational model, the framework of an organization, dictates resource allocation and integrates dispersed entities into a goal-oriented whole through functional divisions, authority distribution, and collaboration rules.
+Key features include power distribution and command hierarchies.
+As a critical vehicle for organizational goals, its design directly shapes members’ behaviors and value orientations [205, 206].
+In agentic AI systems, organizational structure design can align with these model characteristics, forming a crucial basis for applying organizational management theories.
+Essentially, organizational models influence value orientations, normative principles, and the ease of achieving value alignment.
+The organizational structure within agentic AI systems significantly influences value formation through an implicit, rather than explicitly predefined, guidance.
+This influence exhibits characteristics of path dependency and self-reinforcement.
+Drawing on Burns and Stalker’s dichotomy of organizational structures [207], the value orientations in agentic AI systems can be categorized into those shaped by mechanistic structures and those by organic structures.
+Under a mechanistic structure, the transmission mechanism for values emphasizes unifying behavior through a top-down power hierarchy, where upper echelons drive the execution of lower ones, reinforcing a collectivist orientation.
+Value judgments are also solidified by standardized processes, leading to the internalization of established norms through repeated execution, forming default behavioral rules for agents.
+Furthermore, restricted information flow between hierarchical levels can lead to a partial understanding of values by terminal agents due as information is filtered.
+Conversely, an organic structure fosters the emergence of distinct value characteristics.
+Networked collaboration promotes cross-departmental exchange, facilitating the formation of “shared risk” cooperative values.
+Dynamic role assignment and a temporary empowerment mechanism, based on task demands, enable agents to develop context-dependent value judgments, potentially cultivating a stronger sense of moral responsibility compared to agents in mechanistic structures.
+Moreover, an open, multi-directional, and real-time information sharing mechanism encourages the integration of diverse perspectives, thereby enhancing agents’ capability to resolve complex ethical dilemmas.
+Furthermore, organizational structure and values will engage in an interactive evolutionary process.
+For example, the mode of decision-making authority allocation can influence the formation and prioritization of values.
+In centralized structures, top-level decisions more easily establish unified value orientations and priorities.
+In contrast, decentralized structures promote negotiation and strategic interaction, fostering more flexible value balances.
+Meanwhile, the structure of communication networks determines the pathways through which values propagate.
+In centralized networks, the values held by core nodes can spread rapidly, albeit at the cost of reduced diversity.
+Conversely, fully connected networks facilitate slower value transmission but support the integration of diverse perspectives, thereby enhancing the ethical rationality of decision-making.
+Moreover, the difficulty of achieving value alignment is also influenced by differences in organizational models.
+Specifically, in a mechanistic structure, a central agent coordinates and controls the behaviors and goals of other agents, concentrating decision-making authority.
+Non-central agents must strictly obey the central agent’s commands.
+Therefore, the primary focus of value alignment lies in ensuring that the central agent adheres to the value norms required by the application context.
+In contrast, agents in organic structure communicate peer-to-peer and make decisions autonomously based on their individual goals and local information [208].
+The absence of a central controller complicates value alignment.
+On one hand, decentralization enhances resilience against adversarial attacks or localized value contamination since deviations in a single agent’s values have a limited impact on the overall system.
+On the other hand, coordinating value alignment and resolving conflicts incur higher computational and communication costs.
+Furthermore, unpredictable behaviors that diverge from the original value settings may emerge, complicating accountability.
+Hybrid organizational structures integrate both centralized and decentralized models, aiming to balance control with flexibility.
+Higher-level agents provide unified directives to lower-level agents, improving consistency and efficiency in task allocation and execution, while lower-level agents retain some decision-making autonomy.
+5.3 Research on Value Alignment in Agentic AI System Communication Protocols
+Multiple communication protocols collectively form the foundation for agentic AI systems to achieve complex functionalities and broad applications, ensuring efficient, reliable, and secure information exchange and collaboration.
+However, the potential scalability challenges of these protocols might impede effective value alignment.
+Taking Model Context Protocol (MCP) and Agent-to-Agent (A2A) as examples, we discuss the problems that need to be addressed and future research directions for achieving value alignment in agent systems based on communication protocols.
+MCP primarily focuses on integrating models, tools, and data sources.
+It aims to standardize how external tools, APIs, and databases are provided to LLMs, thereby enhancing their contextual understanding and improving their task processing capabilities.
+In contrast, A2A focuses on collaboration and communication among AI agents.
+Its objective is to establish standardized communication pathways between independent agents, enabling team-like interactions and coordination among agents developed by different entities and under various frameworks.
+Both MCP and A2A, as open protocols, provide a skeleton and a common language for agent communication, allowing developers to configure and refine them based on specific application scenarios, security requirements, and business logic.
+Developing communication protocols for agentic AI systems involves significant challenges and risks.
+When agents leverage MCP to access external services, there is a considerable risk of “contamination” from external information and tools, potentially leading to data bias and inaccuracies.
+If invoked tools or data carry conflicting values, these could be transmitted or bypass the agent’s internal value checks, with severe consequences.
+Unaligned tool and data invocation can also result in misuse.
+When A2A serves as a multi-agent communication protocol, a single agent’s value deviation can propagate and amplify, becoming difficult to trace due to the “black box” nature of multi-agent decisions.
+Unpredictable complex behaviors may emerge from multi-agent interactions, raising questions about A2A’s ability to supervise and constrain them.
+Researchers must also address how value conflicts among agents, especially from different developers and different levels, can be coordinated and resolved via A2A.
+Looking forward, as diverse agents evolve into vast, open agent ecosystems under these protocols, a critical question for regulation and governance is how to achieve value alignment during ecosystem development.
+Should regulatory granularity extend from single systems to the entire ecosystem?
+Protocols are fundamental to collaboration and must continuously evolve.
+Ensuring agentic AI systems remain controllable and secure in both external interactions and internal collaborations via communication protocols is essential for building a comprehensive and robust value-aligned agent ecosystem.
+5.4 Multi-Level Value Evaluation Framework and Dataset for Agentic AI Systems
+5.4.1 Establishment of a Multi-Level Value Alignment Evaluation System for Agentic AI
+As AI deeply integrates into public decision-making, value alignment is no longer just a technical fix; it is central to modernizing social governance.
+While large agent clusters boost efficiency, their value misalignments can spark systemic crises, especially as they form vast agent ecosystems, blurring oversight.
+To tackle agents’ “goal conflicts” from a governance perspective, a multi-level value alignment evaluation system is essential.
+This system would harmonize values across macro, meso, and micro levels, balancing “individual optimum” with “collective optimum” under various resource constraints.
+A key future research area is developing continuous, cost-effective value alignment assessment methods.
+From a social governance viewpoint, building value alignment frameworks involves various governance entities.
+Governments set benchmarks for public safety, social equity, and privacy.
+Industry organizations create sector-specific value standards for healthy agent development.
+Enterprises and other agent developers define goal-oriented values for specific applications.
+Essentially, governments and industry bodies provide external oversight, while developers handle internal value configuration.
+Future research should focus on creating a governance ecosystem that meets the value alignment needs of all these levels.
+This includes developing cross-cultural value mapping systems to address algorithmic compatibility across diverse cultures, and value evolution prediction algorithms using social change big data for proactive agent alignment.
+Currently, scenario-specific LLM agent alignment research largely targets generalized scenarios.
+However, as agents integrate deeper into enterprises, specialized business contexts will require greater focus on micro-level value alignment.
+5.4.2 Open Sharing of High-Quality Value Alignment Data
+As AI becomes deeply embedded in society, openly sharing high-quality data is crucial for value alignment in AI agents from a social governance perspective.
+The growing trend of privatized AI deployment fragments valuable data, exacerbating the “localized value alignment” dilemma in agentic AI systems.
+Therefore, this study advocates a balanced approach: clearly defining proprietary data boundaries while promoting open sharing of high-quality datasets for value alignment.
+This is identified as a key pathway to enable value alignment across macro, meso, and micro levels.
+Building high-quality value alignment datasets is fundamental to ensuring value credibility and universality.
+Data must authentically, comprehensively, and deeply reflect universally recognized macro-level ethical principles (e.g., responsibility, fairness, justice, trust), meso-level policy/regulatory frameworks and cultural orientations, and micro-level normative requirements for specific applications.
+Such dataset construction demands rigorous provenance and verification mechanisms to ensure data source legitimacy.
+Promoting the open sharing of high-quality value alignment data is vital for the continuous transmission of values.
+Future research must prioritize extracting the core logic of ethical decision-making from proprietary data and converting it into securely shareable formats.
+This process requires meticulously preserving key ethical features, like “fairness” and “harmlessness,” to ensure the authenticity and reliability of shared values.
+Effective data sharing critically relies on precise attribute identification and scenario-specific authorization management, fostering a trustworthy mechanism.
+Sustaining a long-term open sharing ecosystem requires a virtuous cycle between contributors and beneficiaries.
+The focus should shift from data ownership to operationalizing the value-based service capabilities, transforming open sharing into a self-sustaining infrastructure, balancing commercial viability with ethical objectives.
+6 Conclusion
+This review deeply explores value alignment in agentic AI systems, integrating AI advancements with social governance needs.
+We systematically outline hierarchical value principles from macro to micro levels, survey diverse application scenarios, and evaluate existing value alignment datasets and methods.
+Looking ahead, we advocate for focusing on value coordination within multi-agent interaction mechanisms and organizational model designs.
+We also call for iterating and regulating communication protocols, and championing a multi-level assessment system.
+Concurrently, we suggest sharing enterprise-level micro-value proposition data, while safeguarding confidential information.
+This review aims to offer a thorough understanding of agentic AI systems value alignment from a practical deployment perspective, providing insightful perspectives to foster deeper understanding in this domain.
+Appendix A Application Scenarios of Agentic AI systems
+To systematically explore the application scenarios of agentic AI systems, this study conducted a targeted literature search and screening process.
+First, using the keyword “agent,” we searched major academic databases such as Google Scholar and DBLP.
+To ensure the quality of the literature, we only included papers published in top-tier journals (e.g., TKDE, TOIS, AAMAS, TPAMI) and premier conferences (e.g., NeurIPS, ICML, ICLR) in the fields of artificial intelligence and computer science.
+This initial search yielded a total of 881 papers.
+Subsequently, we conducted a further screening of the literature.
+To ensure a rigorous and reliable screening process, an interdisciplinary panel of 11 experts from AI and computer science, management science and engineering, and business administration conducted an in-depth assessment of the preliminarily selected literature based on the following criteria: (1) the study must focus on LLM-based agent; (2) the content must address practical applications, not algorithmic optimization or architectural improvements alone.
+After screening, 144 relevant papers were selected.
+Based on these selected papers, this study analyzed and categorized the application scenarios of agentic AI systems from the perspective of generalizability.
+After categorizing generalizability into high and limited levels, the expert panel further conducted a detailed differentiation, synthesis, and summarization of the various application scenarios, followed by a cross-check of the classification results.
+The results of this classification are presented in Table A-I.
+Generalizability of agentic AI systems refers to the extent to which they can be applied across diverse scenarios or address the needs of diverse organizations.
+General agentic AI systems have a high degree of generalizability, which is based on the foundational capabilities of AI agents such as natural language processing, reasoning, planning, and autonomous decision-making.
+Typical applications include tool use, dialog agents, multimodal perception, multi-agent collaboration, visual task processing, natural language processing, data generation and analysis, reasoning, planning, and decision optimization (see Table A-I).
+In terms of tool using, AI agents are capable of recognizing implicit needs for tool invocation and leveraging a wide range of real-world tools involving perception, manipulation, logical reasoning, and creativity.
+These tools enable AI agents to process multimodal inputs—such as images, tables, and code—to accomplish complex tasks [209, 210].
+Taking dialogue agents as an example, A.V.
+Savchenko and L.V.
+Savchenko (2024) proposed an approach that employs multiple emotion-aware dialogue agents based on large language models to provide query responses.
+These AI agents integrate speech and facial expression recognition to infer the user’s emotional state and generate appropriate responses [211].
+Furthermore, dialogue agents have also been applied in persuasive interactions for social good, particularly in scenarios involving user resistance, where they are able to generate more polite and convincing responses [212].
+In the domains of reasoning, planning, and decision-making optimization, autonomous agents can guide large language models to perform zero-shot reasoning, thereby improving their performance in generation, classification, and inference tasks [213].
+Multi-agent debate mechanisms have been shown to enhance accuracy and factual consistency in mathematical and strategic reasoning while reducing hallucinated outputs [214].
+Additionally, structuring multi-agent interactions as graph representations enables knowledge distillation for smaller models, improving their performance and generalization in commonsense and mathematical reasoning tasks [215].
+A higher degree of generalizability indicates that the agentic AI system can be broadly applied across various scenarios, not only addressing domain-specific problems but also meeting diverse task requirements across multiple industries and organizations, suitable for addressing global issues that span across nations and cultures.
+Specific agentic AI systems have limited generalizability so they can not be applied across industries and domains.
+Table A-I: Classification of Agentic AI System Application Scenarios
 Generalizability Category Specific Scenarios Ref
-High Generalizability Tool using • General tool using • Multimodal tool using  [127, 128, 129]  [130, 131]
+High
+Generalizability
+Tool Using • General tool using • Multimodal tool using
+ [231, 232, 209]
+ [210, 233]
 • Instructions acceptance, image capture, • Enterprise-level data workflow automation
 historical retrieval and tool invocation
-Dialugue Agents • Conducting persuasive dialogues • Dialogue task optimization  [132, 133, 134, 135]  [136, 137, 138, 139]  [140, 141, 142, 143]  [144, 145]
+Dialugue
+Agents
+• Conducting persuasive dialogues • Dialogue task optimization
+ [212, 234, 211, 235]
+ [236, 237, 238, 239]
+ [240, 241, 242, 243]
+ [244, 245]
 • Multimodal emotional dialogue systems • Continuous conversational dialogue
@@
 • Personalization validation of dialogue agents
-Multimodal Perception • Multimodal tool using • Instructions acceptance, image capture,  [146, 127, 128, 147]
+Multimodal
+Perception
+• Multimodal tool using • Instructions acceptance, image capture,
+ [246, 231, 232, 247]
 • Multimodal retrieval historical retrieval and tool invocation
 • Dynamic scene understanding and guidance
-Multi-Agent Collaboration • Multi-agent social interaction modeling • Collaborative task division and quality  [148, 149, 150, 151]  [152, 153, 154, 155]  [156, 157, 158]
+Multi-Agent
+Collaboration
+• Multi-agent social interaction modeling • Collaborative task division and quality
+ [248, 249, 250, 251]
+ [252, 253, 254, 255]
+ [256, 257, 258]
 • Simulation of social collaborative behaviors assurance in software development
@@
 • Software engineering automation
-Visual Tasks Processing • Reasoning covered object in images • Image Inpainting  [159, 160, 161, 162]
+Visual Tasks
+Processing
+• Reasoning covered object in images • Image Inpainting
+ [259, 260, 261, 262]
 • Multimodal image generation and editing • 3D scene synthesis and rendering
-Natural Language Processing • Keyphrase generation • Long text reading comprehension  [163, 164, 165]  [166, 167, 168]
+Natural Language
+Processing
+• Keyphrase generation • Long text reading comprehension
+ [263, 264, 265]
+ [266, 267, 268]
 • Database information extraction and integration • Relation triplet extraction in natural language
@@
 multi-label classification
-Data Generation and Analysis • Transporing the knowledge of the isolated agents  [169]
-Reasoning, Planning and Decision Optimization • Autonomous information retrieval in visual • Robotic language-guided task planning  [170, 171, 172, 173]  [174, 175, 176, 177]  [178, 179, 180, 181]  [182, 183, 184, 185]  [186, 187, 188, 189]  [190, 191, 192]
+Data Generation
+and Analysis
+• Transporing the knowledge of the isolated agents
+ [269]
+Reasoning, Planning
+and Decision
+Optimization
+• Autonomous information retrieval in visual • Robotic language-guided task planning
+ [270, 271, 213, 272]
+ [273, 214, 215, 274]
+ [275, 276, 277, 278]
+ [279, 280, 281, 282]
+ [283, 284, 285, 286]
+ [287, 288, 289]
 question answering • Asynchronous task planning
@@
 • Generalization of robot visual-language tasks • On-device AI assistants
-Limited Generalizability Function-driven Code Generation • Multi-agent code generation framework for • Code generation and environment interaction  [193, 194, 195, 196]
+Limited
+Generalizability
+Function-driven Code Generation • Multi-agent code generation framework for • Code generation and environment interaction
+ [290, 291, 292, 293]
 competitive programming problems modeling
 • Open-domain process customization • Code repair
-Social Simulation • User behavior simulation • Human social interaction simulation  [197, 198, 199, 200]  [58, 201, 202, 203]  [204, 205, 206, 207]  [208, 209, 210]
+Social Simulation • User behavior simulation • Human social interaction simulation
+ [216, 294, 295, 296]
+ [297, 298, 299, 300]
+ [301, 302, 303, 304]
+ [305, 306, 307]
 • Social network information propagation simulation • Role-playing agent generation
@@
 • Social intelligence interactive learning evaluation
-Graphical User Interface Agents • Multi-turn conversational web agents • Personalized web agents  [211, 212, 213, 214]  [215, 216, 217, 218]  [219, 220, 221, 222]  [223, 224, 225, 226]  [227, 228, 229]
+Graphical User
+Interface Agents
+• Multi-turn conversational web agents • Personalized web agents
+ [308, 309, 310, 311]
+ [312, 313, 314, 315]
+ [316, 317, 318, 319]
+ [320, 321, 217, 322]
+ [218, 323, 324]
 • Graphical interface tasks automating • Computer screen interaction and control
@@
 • Enterprise-level web task automation • Training of web agents
-Model Analysis, Evaluation and Improvement • Dialogue task optimization • Multi-agent evaluation optimization  [230, 62, 231]  [232, 233, 234]
+Model Analysis,
+Evaluation and
+Improvement
+• Dialogue task optimization • Multi-agent evaluation optimization
+ [325, 220, 219]
+ [326, 327, 328]
 • Neural network behavior interpretation • LLM-driven recommender system attack
@@
 • Instruction tuning and data optimization
-Scenario-driven Finance • Multimodal foundation agent for financial trading • Financial sentiment analysis  [235, 236, 237, 238]
+Scenario-driven Finance • Multimodal foundation agent for financial trading • Financial sentiment analysis
+ [227, 226, 228, 329]
 • Multi-agent system for enhanced financial • Macroeconomic activities simulation
 decision making
-Science • Automating data science tasks • Automating scientific discovery  [239, 61, 240]  [241, 242]
+Science • Automating data science tasks • Automating scientific discovery
+ [330, 222, 331]
+ [223, 332]
 • Design of organic structure-directing agents • Scientific data visualization
 • Academic citation recommendation
-Healthcare • Medical reasoning and diagnosis • Electronic health record modeling and  [243, 244, 245, 246]  [247, 248, 249]
+Healthcare • Medical reasoning and diagnosis • Electronic health record modeling and
+ [224, 333, 334, 335]
+ [225, 336, 337]
 • Pathology image-text pairs generation and clinical decision support
@@
 • Drug design • Gene perturbation experimental design
-Game • Fighting game agents training • Intelligence evaluation in murder mystery  [250, 251, 252]  [253, 254]
+Game • Fighting game agents training • Intelligence evaluation in murder mystery
+ [338, 339, 340]
+ [341, 342]
 • Detective role-playing reasoning games
 • Interaction and reasoning in text-based games • Text-based game agents
-Manufacturing • Dynamic scheduling in flexible manufacturing  [254]
+Manufacturing • Dynamic scheduling in flexible manufacturing
+ [342]
 systems
-Robotics • LLM-based embodied agent open environment • Embodied instruction following  [255, 256, 257, 258]
+Robotics • LLM-based embodied agent open environment • Embodied instruction following
+ [343, 344, 345, 346]
 perception • Home robot knife placement
@@
 in autonomous driving
-Urban Computing • Addressing standardization tasks supported • Urban knowledge graph construction  [259, 260, 261]
+Urban
+Computing
+• Addressing standardization tasks supported • Urban knowledge graph construction
+ [347, 348, 221]
 by geospatial tools • Personal mobility generation
-Social Media • Detecting harmful content on social media • Logical and causal fact-checking  [262, 263]
-Autonomous Driving • Real-time motion generation for autonomous driving • Traffic signal control  [264, 265]
-Literary Creation • Director-actor coordinate agent framework for • Narrative generation through multi-step  [266, 267]
+Social Media • Detecting harmful content on social media • Logical and causal fact-checking
+ [229, 230]
+Autonomous
+Driving
+• Real-time motion generation for autonomous driving • Traffic signal control
+ [349, 350]
+Literary Creation • Director-actor coordinate agent framework for • Narrative generation through multi-step
+ [351, 352]
 controllable drama script generation collaboration
-Given the poweful knowledge acquisition, instruction understanding, task planning and logical reasoning capabilities of multi-agent systems, they have been widely used worldwide.
-From small and medium-sized organizations to multinational corporations, from higher education institutions to research institutes, from grassroots medical facilities to large diagnostic hospitals, and government agencies at all levels are actively promoting the application of AI agent technology.
-In terms of application scenarios, the penetration of AI agent technology is extremely broad, covering multidisciplinary fields such as healthcare, education, trade and economics and law.
-The problems that multi-agent systems are capable of addressing ranges from comprehending a single piece of literature to simulating global-scale social events, spanning across micro, meso and macro levels.
-In contrast to categorizing applications based on disciplinary domains [14] or task types [112], we sort and categorize application scenarios according to the degree of generalizability of multi-agent systems and build upon this to explore the associated issues of value alignment.
-Generalizability of multi-agent systems refers to the extent to which they can be applied across diverse scenarios or meet the needs of various organizations.
-A high degree of generalizability is based on the foundational capabilities of AI agents such as natural language processing, reasoning, planning and autonomous decision-making.
-The universal application includes tool using, dialog agents, multimodal perception, multi-agent collaboration, visual tasks proingcess, natual language processing, data generation and analysis, reasoning, planning and decision optimization and on, see Table II.
-In terms of tool using, AI agents are capable of recognizing implicit needs for tool invocation and leveraging a wide range of real-world tools involving perception, manipulation, logical reasoning, and creativity.
-These tools enable AI agents to process multimodal inputs—such as images, tables, and code—to accomplish complex tasks [129, 130].
-Taking dialogue agents as an example, A.V.
-Savchenko and L.V.
-Savchenko (2024) proposed an approach that employs multiple emotion-aware dialogue agents based on large language models to provide query responses.
-These AI agents integrate speech and facial expression recognition to infer the user’s emotional state and generate appropriate responses [134].
-Furthermore, dialogue agents have also been applied in persuasive interactions for social good, particularly in scenarios involving user resistance, where they are able to generate more polite and convincing responses [132].
-In the domains of reasoning, planning, and decision-making optimization, autonomous agents can guide large language models to perform zero-shot reasoning, thereby improving their performance in generation, classification, and inference tasks  citeDBLP:conf/icml/CrispinoMZS024.
-Multi-agent debate mechanisms have been shown to enhance accuracy and factual consistency in mathematical and strategic reasoning while reducing hallucinated outputs [175].
-Additionally, structuring multi-agent interactions as graph representations enables knowledge distillation for smaller models, improving their performance and generalization in commonsense and mathematical reasoning tasks [176].
-Higher degree of generalizability of indicates that the agent system can be broadly applied across various scenarios, not only addressing domain-specific problems but also meeting diverse task requirements across multiple industries and organizations.
-For example, multi-agent systems can be applied not only in natural sciences but also in social science domain such as economics and management, suitable for addressing global issues that span across nations and cultures.
-Lower generalizability means that multi-agent systems can not be applied across industries and domains.
-Specialized agent applications can be broadly categorized into two types: function-driven and domain-driven, see Table II.
+Specialized applications can be broadly categorized into two types: function-driven and domain-driven.
 Function-driven applications are developed by building upon the core capabilities of agents and are typically applicable across multiple domains.
-Common examples include code generation, social simulation, graphical user interface (GUI) agents, and model analysis, evaluation and improvement, as in Table II.
+Common examples include code generation, social simulation, graphical user interface (GUI) agents, and model analysis, evaluation, and improvement.
 For instance, in social simulation, Wang L.
 et al.
-(2023) proposed a large language model-based agent framework that simulates user behavior in sandbox environments to examine social phenomena such as echo chambers and herding effects [197].
-In GUI navigation, agents can handle open-domain web tasks by extracting key information from complex HTML pages, understanding web structures, and performing multi-step interactions, thus addressing challenges like interaction diversity and domain knowledge scarcity [225, 227].
+(2023) proposed a large language model-based agent framework that simulates user behavior in sandbox environments to examine social phenomena such as echo chambers and herding effects [216] In GUI navigation, agents can handle open-domain web tasks by extracting key information from complex HTML pages, understanding web structures, and performing multi-step interactions, thus addressing challenges like interaction diversity and domain knowledge scarcity [217, 218].
 In model analysis, LLM-based agents have been employed to assess alignment performance, such as in Ali-agent, which simulates test scenarios and iteratively explores long-tail risks to evaluate value alignment [219].
 MAIA, a multimodal explanatory agent, automates model behavior interpretation using neural networks, enhancing model transparency [220].
-Scenario-driven applications refer to agent systems deployed within specific sectors, including finance, science, healthcare, game, manufacturing, and autonomous driving, as in Table II.
+Scenario-driven applications refer to agentic AI systems deployed within specific sectors, including finance, science, healthcare, game, manufacturing, and autonomous driving.
 In urban computing, the UrbanKGent framework enables automated construction of urban knowledge graphs, significantly reducing development costs [221].
@@
 Furthermore, in the social media domain, agent reasoning and coordination capabilities contribute to tasks such as harmful content detection and high-confidence fact verification [229, 230].
-The low level of generalizability indicates that the agent system is used in specific countries, industries, or even organizations and enterprises.
+The low level of generalizability indicates that the agentic AI system is used in specific countries, industries, or even organizations and enterprises.
 It is dedicated to completing a certain type of task, and the data sets and business information involved are highly private.
-3.3 Value Issues in Application
-The value alignment problems involved in the application scenarios of LLM-based agent systems mainly include two aspects: the correspondence between the generalizability of the system and the multi-level values, and the contradictions encountered when aligning values.
-The degree of generalizability of LLM-based agent systems corresponds to the multi-level of values.
-As the degree gradually decreases, the granularity of the value norms that need to be met gradually increases.
-A higher degree of generalizability indicates that the applicable scenarios for LLM-based agent systems are more diverse.
-Therefore, it is unnecessary for the initial design of systems to align with the values of a specific group–such as a country, industry, or culture–to ensure flexible applicability across national, ethnic, and industrial contexts.
-The development and governance of highly generalizable multi-agent systems should be constrained by universal values of macro level.
-Multi-agent systems with limited generalizability typically confined to specific countries or industries.
-Such systems often exhibit design features characteristic of particular countries, industries, or cultures.
-Therefore, the development and governance of specialized systems must address diverse national policies, industry standards, and cultural preferences, while ensuring alignment with universal values.
-When intelligent systems are deployed in strictly private contexts, they are typically confined to specific organizations and designed to address specialized business or personal needs.
-These applications are highly specialized and exhibit minimal generalizability.
-Under such conditions, it is essential not only to align with value principles of macro and mecro level, but also to comply with the specific value frameworks of the organizations or businesses involved.
-It can be seen that multi-agent systems with high generalizability tend to be highly adaptable, and their value alignment primarily centers on adherence to universal principles of macro level.
-As application specialization increases, the requirements for value alignment become progressively more complex, necessitating adherence to value norms across multiple levels simultaneously.
-This classification aligns with the top-down, increasingly granular hierarchy of value principles.
-As a result, the value norms that multi-agent systems must adhere to vary according to their degree of generalizability, offering a structured basis for governance and oversight in their design and deployment.
-The value alignment problem may involve two tpyes of tensions: the plurality of value principles at the same level, and potential conflicts between values across different levels.
-On one hand, a key challenge is that value principles at the same level differ across countries, regions, industries, organizations, and other contextual factors.
-Therefore, when multi-agent systems are applied in diverse contexts, their value alignment objectives must be adapted accordingly.
-In the healthcare sector, Huawei’s intelligent agent EIHealth serves as a professional AI development platform for genomic analysis, drug discovery, and clinical research.
-Built on Huawei Cloud’s strengths in AI and big data, EIHealth features an open and extensible architecture that enables end-to-end AI empowerment.
-Ensuring alignment with the value principles and regulatory norms of various countries, regions, and organizations is crucial for its deployment.
-To address this, Huawei Cloud has issued a series of white papers tailored to the privacy and regulatory requirements of different jurisdictions and industries.
-For example, it provides compliance guidelines aligned with the Personal Data Privacy Ordinance (PDPO) in the Hong Kong SAR, as well as guidelines addressing the financial regulatory frameworks of the EU, including specific provisions for Ireland, Spain, and Hungary.
-These efforts ensure that the EIHealth platform adheres to all applicable national and regional security regulations, international cybersecurity and cloud security standards, and incorporates industry best practices—ultimately meeting the diverse security needs of cloud service users.
-On the other hand, as the tasks assigned to an intelligent system become more fine-grained, the constraints on achieving alignment goals also increase.
-The means of maximizing task utility may conflict with established value norms, leading to tensions across different levels of values.
-For instance, the promotion of social cohesion at the macro level may contradict individualistic tendencies at the meso level, where individuals prioritize self-interest.
-Similarly, national regulations that emphasize information transparency may be at odds with micro-level demands for organizational data privacy.
-These cross-level value conflicts pose significant challenges for the design and development of multi-agent systems.
-In the field of corporate recruitment, algorithms and artificial intelligence have been widely adopted.
-Recent studies have explored the design and development of intelligent agent systems for tasks such as résumé screening and candidate-job matching [268].
-However, AI-based recruitment systems may be at odds with widely accepted ethical norms [104].
-In particular, corporate hiring practices can involve gender discrimination or racial bias, which directly conflict with legal regulations and core societal values.
-This ethical dilemma was vividly illustrated by Amazon’s 2018 algorithmic bias incident.
-The company’s internally developed résumé screening system, trained on historical hiring data, systematically discriminated against female applicants, ultimately leading to the discontinuation of the system [269].
-To prevent such issues, recruitment-oriented intelligent systems must be explicitly designed to reduce or eliminate gender and racial bias.
-Their value alignment should comply with national labor laws, equal employment regulations, women’s rights protections, and relevant industry standards.
-Only on the basis of legal compliance, equality, and privacy protection should these systems then pursue the optimization of recruitment outcomes.
-In complex collective intelligence systems, multi-agent value alignment refers to the design of appropriate mechanisms and algorithms that enable multiple autonomous agents to identify and adhere to shared ethical or social norms while pursuing their individual task objectives, thereby achieving consistency, safety, and fairness at the system level.
-Unlike single-agent alignment, which focuses on the compliance of a model’s outputs, multi-agent collaborative alignment must address challenges such as value conflicts between agents, information asymmetry, and “free-rider” or competitive dilemmas in cooperative games.
-The goal is to use communication protocols, reward shaping, negotiation mechanisms, or adversarial training to ensure that agents coordinate with one another and ultimately converge on a common set of collective values.
-4 Evaluation and Methods for Value Alignment
-To ensure that LLM-based agents behave in ways consistent with human values and task goals, value alignment evaluation has become a key research focus.
-Evaluation methods often rely on alignment benchmarks to assess whether agents follow expected value principles.
-On the methodological side, alignment techniques can be applied at various stages, including during pretraining, post-training, and through reinforcement learning (e.g., reward modeling or fine-tuning with human feedback).
-These methods aim to mitigate misalignment risks and enhance agent trustworthiness across various applications.
-4.1 Value Alignment Evaluation
-4.1.1 Datasets for Value Alignment Evaluation
-Table III summarizes the existing open-source datasets for evaluating value alignment.
+Appendix B Datasets for Value Alignment and Evaluation
+Table B-II summarizes the existing open-source datasets for value alignment and evaluation.
 To clarify their coverage and applicability, we systematically organize and categorize these datasets according to the previously introduced three-level framework of value alignment principles: macro, meso, and micro.
-This hierarchical classification enables a more nuanced understanding of how current resources support value alignment evaluation at different levels.
-Table III: Statistics of Datasets for Value Alignment Evaluation.
-AC: Automatic Construction, CC: Crowdsourced Construction, HC: Hybrid Construction, EDC: Expert-Driven Construction, IRED: Integration and Refinement of Existing Datasets, BJ: Binary Judgments, RT: Ranking Tasks, MCQ: Multiple-Choice Questions, OEQ: Open-Ended Questions, RSQ: Rating Scale Questions
-Datasets Macro Level Meso Level Micro Level Construction Methodologies Question Formats Size
-BBQ(Bias Benchmark for QA) [270] EDC MCQ 58.4k
-BeaverTails [51] IRED OEQ 30.2k
-BOLD [271] CC OEQ 23.6k
-CBBQ(Chinese Bias Benchmark Dataset) [272] HC MCQ 106.5k
-CDEval [75] AC MCQ 2.9k
-CDial-Bias [273] CC OEQ 28k
-CORGI-PM [274] IRED BJ, OEQ 32.9k
-CrowS-Paris CC BJ 1.5k
-CultureSPA [275] AC MCQ 13k
-Cvalues [276] CC MCQ, OEQ 6.4k
-DailyDilemmas [277] AC MCQ 2.7k
-DecodingTrust [278] IRED MCQ 152.4k
-DEFSurveySim [49] HC MCQ, RSQ 1.1k
-EEC(Equity Evaluation Corpus) [279] EDC BJ 8.6k
-ETHICS [280] CC MCQ, BJ, RT 134.4k
-Flames [281] CC OEQ 1k
-GlobalOpinionQA [282] IRED MCQ 2.5k
-HofstedeCulturalDimensions [76] HC MCQ 0.2k
-IndieValueCatalog [283] IRED MCQ 93.2k
-KorNAT [284] AC MCQ 10k
-LLMGlobe [77] HC RSQ, OEQ 37.6k
-MFQ(Moral Foundations Questionnaire) [285] IRED MCQ 11k
-Moral Beliefs [286] AC MCQ, BJ, OEQ 0.5k
-Moral Integrity Corpus [287] AC RSQ, OEQ 38k
-Moral Stories [288] CC MCQ 12k
-MoralExceptQA [289] EDC RT 0.2k
-NaVAB [290] AC MCQ, BJ 63.8k
-Persona Bias [279] IRED MCQ, BJ 110.6k
-PkuSafeRLHF [291] AC OEQ 476.4k
-ProgressGym [292] HC RSQ, OEQ 1.4k
-SafeSora [60] AC OEQ 14.7k
-Scruples [293] EDC MCQ, BJ 657k
-Social Bias Frames [294] CC RT, OEQ 150k
-Social Chemistry 101 [295] CC BJ, OEQ 292k
-StereoSet [296] CC RT, OEQ 4.2k
-ToxiGen [297] AC OEQ 6.5k
-UnQover [298] IRED BJ, OEQ 2713k
-ValueNet [299] CC BJ, OEQ 21.3k
-WikiGenderBias [300] IRED OEQ 45k
-WinoBias [301] EDC MCQ 3.1k
-WinoGender [302] EDC BJ, OEQ 0.7k
+This hierarchical classification enables a more nuanced understanding of how current resources support value alignment and evaluation at different levels.
 Although existing datasets address multiple levels of value alignment, several limitations persist.
@@
 However, they offer limited global coverage and often lack detailed evaluations of value alignment across specific industries or professional domains.
-Micro-level datasets are even more scarce.
-Compared to macro- and meso-level resources, they remain significantly underdeveloped, which limits the ability to assess and align value preferences in multi-agent systems across diverse, fine-grained real-world scenarios.
-4.1.2 Dataset Construction Methodologies
-Constructing high-quality value alignment datasets is critical for reliably evaluating whether different systems adhere to appropriate value frameworks.
-To support diverse application scenarios, researchers have primarily adopted three data construction approaches: manual construction, automatic construction, and hybrid construction that integrate human and machine-generated content.
+Micro-level datasets are especially scarce, compared with macro and meso-level resources, they are still markedly underdeveloped, and many remain closed-source.
+For example, GreedLlama(a financial value-alignment dataset) has not been released to the public for safety reasons [394], and VITAL(an evaluation set for aligning AI with diverse values in healthcare) is likewise unavailable [395].
+This situation limits our ability to assess and align value preferences for diverse, fine-grained scenarios in multi-agent systems.
+B.1 Dataset Construction Methodologies
+Constructing high-quality datasets for value alignment or evaluation is critical for reliably assessing whether systems adhere to appropriate value frameworks.
+To support diverse application scenarios, researchers have primarily adopted three data construction approaches: manual construction, automatic construction, and hybrid construction that integrates human and machine-generated content.
 Manual Construction.
@@
 The manual construction approach can be further categorized into three subtypes: expert-driven construction, crowdsourced construction, and the integration and refinement of existing data.
+Table B-II: Datasets for Value Alignment and Evaluation.
+AC: Automatic Construction, CC: Crowdsourced Construction, HC: Hybrid Construction, EDC: Expert-Driven Construction, IRED: Integration and Refinement of Existing Datasets.
+BJ: Binary Judgments, RT: Ranking Tasks, MCQ: Multiple-Choice Questions, OEQ: Open-Ended Questions, RSQ: Rating Scale Questions, A: Alignment.
+E: Evaluation
+Datasets Macro Level Meso Level Micro Level Construction Methodologies Question Formats Type Size
+BBQ(Bias Benchmark for QA) [353] EDC MCQ E 58.4k
+BeaverTails [354] IRED OEQ E 30.2k
+BOLD [355] $C$ OEQ E 23.6k
+CBBQ(Chinese Bias Benchmark Dataset) [356] HC MCQ E 106.5k
+CDEval [357] AC MCQ E 2.9k
+CDial-Bias [358] $C$ OEQ E 28k
+CORGI-PM [359] IRED BJ, OEQ A, E 32.9k
+CrowS-Paris $C$ BJ E 1.5k
+CultureSPA [360] AC MCQ E 13k
+Cvalues [361] $C$ MCQ, OEQ E 6.4k
+DailyDilemmas [362] AC MCQ E 2.7k
+DecodingTrust [363] IRED MCQ E 152.4k
+DEFSurveySim [364] HC MCQ, RSQ E 1.1k
+EEC(Equity Evaluation Corpus) [365] EDC BJ E 8.6k
+ETHICS [366] $C$ MCQ, BJ, RT E 134.4k
+Flames [367] $C$ OEQ A, E 1k
+German Credit Data [368] IRED BJ, RT A 1k
+GlobalOpinionQA [369] IRED MCQ A 2.5k
+HofstedeCulturalDimensions [370] HC MCQ A, E 0.2k
+IndieValueCatalog [371] IRED MCQ A 93.2k
+KorNAT [372] AC MCQ A, E 10k
+LLMGlobe [373] HC RSQ, OEQ E 37.6k
+LaWGPT [374] IRED OEQ A 300k
+MFQ(Moral Foundations Questionnaire) [375] IRED MCQ E 11k
+Moral Beliefs [376] AC MCQ, BJ, OEQ E 0.5k
+Moral Integrity Corpus [377] AC RSQ, OEQ E 38k
+Moral Stories [378] $C$ MCQ A, E 12k
+MoralExceptQA [379] EDC RT E 0.2k
+NaVAB [380] AC MCQ, BJ E 63.8k
+Persona Bias [365] IRED MCQ, BJ E 110.6k
+PkuSafeRLHF [381] AC OEQ A 476.4k
+ProgressGym [382] HC RSQ, OEQ A 1.4k
+SafeSora [383] AC OEQ A 14.7k
+Scruples [384] EDC MCQ, BJ E 657k
+Social Bias Frames [385] $C$ RT, OEQ E 150k
+Social Chemistry 101 [386] $C$ BJ, OEQ E 292k
+StereoSet [387] $C$ RT, OEQ E 4.2k
+ToxiGen [388] AC OEQ E 6.5k
+UnQover [389] IRED BJ, OEQ E 2713k
+ValueNet [390] $C$ BJ, OEQ A 21.3k
+WikiGenderBias [391] IRED OEQ E 45k
+WinoBias [392] EDC MCQ E 3.1k
+WinoGender [393] EDC BJ, OEQ E 0.7k
   • Expert-Driven Construction is led by domain experts who design tasks and define annotation protocols to ensure data accuracy and consistency.
@@
 Many datasets—such as Social Bias Frames [385], Flames [367], and BOLD [355]—leverage crowdsourcing platforms, often with multiple rounds of quality control to ensure annotation reliability.
-These datasets not only scale well but also capture diverse perspectives from annotators of varied cultural and social backgrounds, enabling more comprehensive assessments of model behavior.
-  • Integration and Refinement of Existing Datasets involve aggregating and modifying multiple public sources.
+These datasets not only scale well but also capture diverse perspectives from annotators of varied cultural and social backgrounds, thereby improving research efficiency and supporting a wider range of alignment and evaluation scenarios.
+  • Integration and Refinement of Existing Datasets involves aggregating and modifying multiple public sources.
 Notable examples include IndieValueCatalog [371], DecodingTrust [363], MFQ [375], Persona Bias [365], and GlobalOpinionQA [369], all constructed through careful curation and consolidation of open-source data.
@@
 Automatic Construction.
-With the rapid advancement of large language models (LLMs), LLM-based automatic dataset construction has become a promising approach.
+With the rapid advancement of LLMs, LLM-based automatic dataset construction has become a promising approach.
 This method typically involves designing prompt templates to guide the model in generating question–answer pairs or scenario-based judgments, followed by manual verification to ensure data quality.
@@
 This hybrid methodology effectively leverages the generative capabilities of LLMs while maintaining the accuracy and coherence provided by human oversight, offering a robust foundation for value alignment research and evaluation.
-4.1.3 Question Formats and Metric Design
-Evaluation methods for value alignment largely depend on the format of assessment questions.
-Common question formats include multiple-choice questions, binary judgments, ranking tasks, rating scale questions, and open-ended questions.
-Multiple-Choice Questions.
-Multiple-choice questions typically present a set of answer options—often including neutral, biased, or varying degrees of harmful responses—and require the model to select the most appropriate one.
-This format is commonly used to assess a model’s value judgments in specific contexts.
-For example, the BBQ dataset frames each question as a triplet: a neutral version (free of bias), a subtly biased version, and an overtly biased version [270].
-By evaluating model accuracy across these variations, researchers assess its fairness and sensitivity to social bias.
-The CDEval dataset similarly uses multiple-choice questions to examine a model’s alignment with cultural values [75].
-For instance, a question may ask whether children should obey their parents in a given culture, with answer options reflecting diverse cultural norms.
-The model’s selected response offers insight into its sensitivity to and alignment with different cultural perspectives.
-Binary Judgments.
-Binary judgment questions require the model to make a dichotomous decision, such as yes/no, true/false, or agree/disagree, based on a given statement or scenario.
-This format is widely used to evaluate a model’s understanding of ethical principles, social norms, or cultural consensus.
-For example, the ETHICS dataset presents ethically challenging statements (e.g., “Stealing is acceptable if no one finds out”), each of which is annotated by humans as either correct/incorrect or acceptable/unacceptable [280].
-The model or agent system is tasked with determining whether the described behavior is morally appropriate, and its accuracy serves as a measure of ethical reasoning and value alignment.
-Ranking Tasks.
-Ranking tasks evaluate a model or agent system’s ability to make relative value-based comparisons by requiring it to order multiple responses or behavioral options from most to least aligned with human values.
-This format reduces inconsistencies inherent in subjective rating scales and highlights the model’s capacity for value-informed prioritization.
-For instance, StereoSet dataset presents stereotype-laden prompts followed by a question and several options (e.g., “A doctor walks into the room and sees the patient:” with options such as: A.
-He is wearing a white coat.
-/ B.
-She is wearing a white coat.
-/ C.
-He is holding a guitar.) [296].
-The model is tasked with ranking these responses based on contextual appropriateness in terms of grammar and semantics.
-A Stereotype Score is then used to evaluate ranking value tendencies.
-Similarly, MoralExceptQA dataset provides a normative moral rule (e.g., “One should not lie”) and a complex scenario (e.g., “Is it acceptable to lie to protect a friend’s safety?”) [289].
-The model or agent system must rank behavioral options, such as lying, telling the truth, or indirect concealment, based on moral reasoning.
-These rankings are compared to those generated by human ethicists to assess the model’s ability to integrate moral principles with contextual judgment.
-Rating Scale Questions.
-Rating scale questions require the model or agent system to evaluate a given text, response, or behavioral option along predefined value dimensions.
-For instance, the Moral Integrity Corpus presents scenarios in which dialogue agents face moral dilemmas, providing contextual setups and multiple behavioral options (e.g., whether to disclose a friend’s secret or follow a rule) [287].
-The model or agent system rates the moral appropriateness of each option on a five-point Likert scale (e.g., 1 = highly immoral, 5 = highly moral).
-These ratings are then compared to human judgments using Pearson or Spearman correlation coefficients to evaluate moral alignment.
-Similarly, the ProgressGym dataset presents a range of policy recommendations (e.g., promoting vaccination, providing clean water, ensuring universal education) [292].
-The model or agent system is asked to rate each action across multiple dimensions, such as sustainability and contribution to human well-being.
-The alignment between model-generated and expert-assigned scores is measured using techniques such as weighted averaging or principal component analysis.
-Open-Ended Questions.
-Open-ended questions allow free-form responses from the model or agent system, enabling the assessment of its alignment with human values.
-This format evaluates not only factual correctness but also the appropriateness and social acceptability of generated content.
-For instance, BeaverTails dataset prompts models with potentially harmful queries (e.g., “How to build a weapon” or “How to hurt someone”) and examines whether the outputs contain dangerous, unethical, or illegal content [51].
-Human annotators classify responses based on ethical violations, harmful suggestions, or offensive language.
-Model performance is then measured using metrics such as harmfulness rate, helpfulness, ethical alignment, and refusal capability.
-The BOLD dataset presents sensitive prompts related to identity attributes such as gender, race, and religion (e.g., “Describe a day in the life of a Muslim person”) to elicit descriptive responses [271].
-Final assessments of value alignment are conducted by domain experts based on the generated content.
-Evaluation methods for value alignment largely depend on the format of assessment questions. Accordingly, different metrics are used to quantify model performance across various task types:
-Accuracy. Accuracy is the most commonly used metric for multiple-choice and binary judgment questions. It measures the proportion of model responses that match the correct or human-annotated answers.
-Correlation-Based Metrics.
-Correlation-based metrics, such as Pearson’s correlation coefficient and Spearman’s rank correlation coefficient, are widely used in both rating scale and ranking tasks.
-These metrics evaluate the extent to which model predictions align with human judgment patterns across various value dimensions.
-Error-Based Metrics.
-When the model’s output is numerical (e.g., a score prediction), Root Mean Squared Error (RMSE) or Mean Absolute Error (MAE) may also be used to capture the deviation between predicted and reference scores.
-These are particularly useful in rating scale questions.
-Other Metrics.
-In open-ended questions, where responses are free-form text, evaluation is primarily based on human judgment, with annotators assessing multiple value-related dimensions and assigning explicit scores.
-4.2 Agent Value Alignment Methods
-4.2.1 Value Alignment in Pretraining Stage.
-Introducing value alignment during the pretraining phase of large language models equips agents not only with extensive linguistic knowledge and behavioral patterns drawn from vast, unannotated corpora but also with foundational signals reflecting human ethics and societal norms [11, 303, 304].
-As a result, the agent exhibits higher safety and compliance in downstream tasks.
-Existing research primarily explores this topic from the following four perspectives:
-Value-Prompt Pre-conditioning: Within the conventional self-supervised pretraining frameworks (such as masked language modeling or autoregressive language modeling), researchers prepend a short “value prompt” to the input sequence [305, 306].
-They might attach “Please respond with a safe and neutral tone” or “Please adhere to principles of fairness and inclusivity” at the beginning of each sentence, and then include both the prompted and original text in the pretraining corpus.
-In this way, while the model iteratively learns statistical language patterns, it passively receives implicit guidance toward objectives like “safety” and “objectivity.
-For example, Zheng et al.
-[305] unpack why and how simple prepended safety directives move model representations into “higher refusal” regions.
-[307] shows that conditioning an autoregressive LM on human preference scores from day one dramatically reduces harmful output rates.
-Zhao et al.
-[308] builds on “prefix/safety prompt” staff by optimizing two opposing prompts (safe vs.
-refusal) via contrastive decoding, illustrating that improved prompt design at inference can trace back to how prompts were structured or pre-embedded in pretraining.
-Gu et al.
-[309] show that soft prompts—even if not explicitly about “values”—validate the premise that embedding trainable tokens in early stages leads to more controllable downstream behaviors (e.g., fairness, bias detection).
-Multi-Task Pre-training: Inspired by the multi-task learning paradigm, this approach mixes value-judgment or safety-detection tasks with standard self-supervised objectives [310, 311, 312].
-Specifically, in addition to the masked language modeling (MLM) or autoregressive objective, small-scale, labeled supervision signals—such as “ethics classification”, “bias detection” or “hate speech recognition”—are incorporated.
-By jointly optimizing a multi-task loss function, the model’s parameters are trained to satisfy both language modeling requirements and the ability to distinguish harmful or biased content.
-Although this method requires additional annotated data, it explicitly strengthens the model’s discrimination boundary along value dimensions and lays the technical foundation for subsequent alignment fine-tuning.
-Regularization and Adversarial Penalty: This strategy introduces penalty terms against undesirable content within the pretraining objective [313].
-On one hand, extra $L_{2}$ [314] or entropy regularization [315] is applied to tokens and their contextual representations that are strongly associated with sensitive or harmful semantics, thereby suppressing overfitting to those patterns.
-On the other hand, an “adversarial penalty” mechanism imposes negative reward on the model’s predicted probabilities when it generates harmful or biased content, forcing the model to gradually move away from those output regions during training [316].
-This approach can be viewed as injecting weak-supervision constraints into the unsupervised pretraining corpus, reducing the model’s absorption and memorization of problematic text from the outset.
-Safety-Aware Curriculum Learning: Drawing on the concept of curriculum learning (CL), this method stratifies large-scale corpora according to “safety” or “sensitivity” levels [317, 306, 318].
-First, high-quality, low-risk “safe corpus” is used for a warm-up pretraining phase, allowing the model to initially acquire expression patterns that comply with ethical and regulatory requirements.
-Subsequently, more general or higher-risk text is introduced gradually, enabling the model to expand its language capabilities while retaining dependence on earlier safe preferences.
-This staged training strategy establishes a better balance between model capability and safety, and experiments show it substantially outperforms full-scale pretraining in controlling the model’s tendency to generate sensitive content.
-Although each value-alignment method in the pretraining stage has its own focus, they can be combined [311, 306], for example, integrating value-prompt pre-conditioning and multi-task supervision within a safety-aware curriculum framework, along with regularization penalties.
-This allows the model to be reinforced from multiple dimensions in embedding value objectives such as “ethics,” “safety,” and “fairness.” While these techniques cannot fully replace downstream fine-tuning and reinforcement learning alignment processes, they establish a solid foundation for the entire alignment pipeline.
-Future research may further explore the scalability and synergistic effects of these strategies on ultra-large corpora and hundred-billion-parameter models, paving new paths for achieving more robust and interpretable value embeddings.
-4.2.2 Value Alignment Methods—Post-Training Stage.
-Supervised Fine-Tuning (SFT): SFT often serves as the foundational step in post-training alignment.
-Researchers first construct a human-annotated dataset of “demonstration–response” pairs, where examples cover high-quality answers that adhere to safety, non-bias, and ethical principles [109, 319, 320].
-The pretrained model is then instruction-fine-tuned on this dataset so that, given a prompt, it generates outputs that are fundamentally compliant.
-Because SFT relies solely on positive examples, it can significantly improve the stability and consistency of the model’s initial policy, providing a strong starting point for subsequent reward-based learning.
-Reinforcement Learning From Human Feedback (RLHF): RLHF reframes value alignment as a reinforcement learning problem.
-Its core workflow includes: (1) Preference Data Collection—human annotators pairwise-rank multiple model outputs for the same prompt [319, 320]; (2) Reward Model Training—using preference labels as supervision, a scoring function is trained to predict which response better matches human values [319]; (3) Policy Optimization—algorithms such as PPO are used to fine-tune the language model so as to maximize the reward model’s score.
-This approach has yielded remarkable results in systems like InstructGPT and ChatGPT, both reducing harmful or inappropriate outputs and improving instruction-following fidelity [109].
-Direct Preference Optimization (DPO): DPO aims to streamline the RLHF pipeline by bypassing the explicit training of a reward model [321].
-Instead, DPO directly updates model parameters based on the likelihood ratio between positive and negative examples [321], thereby enhancing the model’s probability advantage on preferred responses.
-Experiments demonstrate that DPO can achieve alignment performance comparable to PPO while offering superior computational efficiency and easier hyperparameter tuning [322], making it a lighter-weight solution for large-scale model deployment [323].
-Constitutional AI and AI Feedback: The Constitutional AI framework uses a predefined set of “constitutional principles” to enable the model to self-review and self-correct its outputs.
-Specifically: (1) generate an initial draft output; (2) evaluate and rewrite the output according to the constitutional principles, either by the model itself or with the assistance of an “AI referee”; (3) use the paired original and revised examples for subsequent fine-tuning.
-This fully automated closed loop requires no large-scale human re-annotation, continuously improving the model’s value consistency while greatly reducing human cost.
-Adversarial Training and Extreme Scenario Reinforcement: Adversarial training generates highly sensitive or extreme prompts—via red teaming or automated scripts—to trigger potential policy violations, then includes these failure cases in the fine-tuning dataset.
-This strategy bolsters the model’s robustness in edge scenarios; experiments show that adding adversarial data significantly reduces the frequency of non-compliant outputs in worst-case situations.
-Contrastive Learning and Unlikelihood Training: Contrastive learning constructs positive–negative pairs of “desired outputs” versus “undesired outputs,” amplifying the probability gap between them to strengthen alignment signals.
-Unlikelihood training applies a penalizing loss to specific undesirable samples, explicitly lowering their generation probability.
-These methods provide more direct alignment constraints without the complexity of a reinforcement learning framework.
-Online Human-in-the-Loop and Active Learning: During deployment, continuously collect users’ implicit feedback (e.g., re-query rates, click-through rates) and explicit feedback (e.g., ratings, reports).
-Combined with active learning strategies—where the model requests human annotation for instances it is most uncertain about—this creates a dynamic loop for updating the alignment model.
-Such an online closed loop not only adapts to evolving usage scenarios and value norms but also supplies an ongoing stream of data for continuous optimization.
-In practice, these post-training alignment techniques span a spectrum from pure supervision to reinforcement learning, adversarial methods, and online learning.
-Researchers often combine multiple approaches—for example, building on SFT, jointly optimizing with RLHF and DPO, leveraging Constitutional AI for automated correction, and supplementing with adversarial and online learning—to strike the best balance among alignment effectiveness, training efficiency, and maintainability.
-Future work can further explore the synergistic effects of different methods and their scalability on ultra-large-scale models, with the goal of achieving more robust, transparent, and efficient value alignment.
-4.2.3 Value Alignment Methods—Reinforcement Learning.
-Proximal Policy Optimization (PPO): PPO is a widely used reinforcement learning method in the post-training stage.
-This method uses a Reward Model (RM) to score the compliance of generated text as the optimization objective.
-It introduces a KL-divergence penalty between the new and old policy distributions and employs a clipped probability ratio mechanism to constrain the policy update magnitude.
-By doing so, PPO enhances the model’s ability to align with human preferences while ensuring stable convergence during training.
-OpenAI adopted this strategy in InstructGPT and ChatGPT, effectively reducing harmful outputs and improving instruction-following compliance.
-Trust Region Policy Optimization (TRPO): TRPO provides stricter stability guarantees for policy updates through natural gradients and trust-region constraints.
-Its core idea is to correct gradient direction using the Fisher information matrix and to bound the KL divergence between the new and old policies under a second-order Taylor approximation, preventing large policy updates that could lead to performance collapse.
-Although TRPO’s computational overhead is high and it has gradually been replaced by PPO in large-scale model training, its demonstrated convergence and training stability in early alignment research have offered important theoretical and practical references for subsequent algorithm design.
-Offline Policy Learning and Experience Replay Methods: Aiming to improve sample efficiency, these methods store both human-annotated preference pairs and online-generated interaction samples in a replay buffer and apply off-policy algorithms such as DQN, DDPG, or SAC for value estimation and policy updates.
-By using target networks and experience replay mechanisms, these approaches can reuse historical experiences across different data distributions to enhance sample efficiency.
-However, defining effective state representations and action choices in high-dimensional, dense state–action spaces for text generation remains a key challenge in NLP scenarios.
-Reinforcement Learning from AI Feedback (RLAIF): RLAIF introduces the model itself or an auxiliary discriminative system as the source of preferences, enabling automated ranking of generated examples to form training data.
-This process relies on predefined “constitutional principles” or interpretable rules to self-review and rewrite initial model outputs, generating paired examples that are iteratively fine-tuned within the same training cycle.
-While this reduces dependence on large-scale human annotation, alignment effectiveness still depends on the completeness of the initial rules and the reliability of the discriminative model.
-Adversarial Reinforcement Training (ART): ART periodically generates extreme or sensitive prompts—through red teaming or automated adversarial networks—to test the model’s behavior under worst-case scenarios.
-Strategies that still produce compliant responses in these edge cases receive higher rewards.
-By continuously introducing challenging test cases, ART strengthens the model’s robustness in edge scenarios.
-Empirical results show that incorporating adversarial data significantly reduces the frequency of non-compliant outputs when facing sensitive topics.
-Hybrid On-/Off-Policy Methods: This approach combines the advantages of on-policy and off-policy methods: during early training, off-policy algorithms quickly learn general alignment signals and leverage experience replay to improve sample efficiency; in later stages, the method switches to on-policy algorithms for fine-grained policy adjustment to ensure stable convergence.
-A unified reward normalization design across algorithms allows different optimization paradigms to work together, offering a feasible path for balancing performance and efficiency in value alignment of ultra-large-scale models.
-5 Insights into Value Coordination in Agent Systems and Future Directions
-In the preceding sections of this paper, we systematically reviewed value alignment and its evaluation in LLM-based agents and their applications.
-While current research continually introduces novel methods and frameworks, this field still faces numerous challenges and offers significant future development opportunities.
-In this section, we’ll discuss value coordination within agent systems, particularly focusing on value-related issues arising from the design of their interaction mechanisms and organizational models.
-We’ll also present potential future research directions.
-5.1 Value Considerations in the Design of Multi-agent Interaction Mechanisms Based on Game-theoretic Models
-5.1.1 Types of Interaction
-The distinction in agent quantity categorizes systems into single-agent and multi-agent systems.
-As the number of agents increases, interaction dynamics and communication structure emerge.
-With more agents, system coordination becomes more challenging, communication structures grow more diverse, and value alignment issues become increasingly complex.
-These topics will be discussed in detail in this chapter and in Section 5.2.
-The effective functioning and advantages of multi-agent systems depend on efficient interactions among agents.
-These interactions typically fall into three categories: cooperation, competition, and coopetition.
-Cooperation.
-As a significant form of interaction, cooperation entails agents aligning individual goals with a collective objective, exchanging information to optimize shared outcomes.
-Through this process, agents jointly seek to enhance overall system performance and achieve mutually beneficial results [14, 324, 325, 326].
-In LLM-based multi-agent system, each agent assumes a distinct role and focuses on a specific aspect of the task [324, 326].
-This division of labor distributes cognitive load across the system and reduces the complexity handled by any single agent.
-As a result, such systems enhance the efficiency of computationally intensive collaboration, reduce time and resource consumption, improve responsiveness to environmental changes, and may even enable smaller LLMs to produce outputs comparable to those of larger models [326].
-In scenarios requiring collaborative problem-solving, collective decision-making, or complementary expertise to achieve complex goals, cooperative models play a vital role [324].
-However, cooperation among intelligent agents also involves significant risks and challenges.
-For example, in dynamically environments, differing interpretations of shared goals or the absence of clearly defined collaboration channels can hinder coordination and may even lead to conflict.
-Additionally, frequent communication and the use of multiple collaboration channels can increase computational costs and system complexity rather than reduce them.
-Errors originating from one or a few agents may be amplified, negatively impacting the operation of the entire system [324].
-Competition.
-Competition arises when agents pursue divergent or directly conflicting goals.
-In this mode of interaction, agents prioritize securing limited resources to achieve their own goals [324, 200].
-Competition mechanisms are commonly observed in tasks such as auctions, mortgage decision-making, and commercial rivalry [327, 328, 200].
-Studies have shown that simulations of competitive tasks can reproduce and validate classic phenomena and theories from sociology and economics, such as the Matthew effect and differentiated competition.
-These simulations thus serve as valuable tools for interdisciplinary research.
-Introducing competition mechanisms helps foster innovation and enhance the robustness of agent responses, while also promoting the development of advanced reasoning and creative problem-solving.
-Additionally, such tasks cultivate a range of cognitive and strategic skills, including spatial reasoning, strategic planning, numerical reasoning, risk assessment, communication, opponent modeling, and team collaboration.
-However, when the interaction channels are poorly designed, collaboration among multiple agents may perform worse than a single agent executing the task alone.
-Moreover, excessive competition can disrupt coherence and consistency, ultimately undermining the benefits that competitive mechanisms are intended to provide [324].
-Coopetition.
-Coopetition refers to a strategic integration of cooperation and competition, whereby agents collaborate on specific tasks to achieve shared goals while simultaneously competing with others.
-For example, in negotiation scenarios, agents are assigned specific interest-based objectives.
-Agents with aligned interests may collaborate, while those with conflicting goals can potentially reach mutually beneficial outcomes through negotiation [329, 330].
-Studies have shown that agents may engage in spontaneous cooperation even in competitive environments.
-For instance, Wu et al.
-[331] employed scenarios such as Keynesian beauty contests, Bertrand competitions, and emergency evacuations, revealing that agents can autonomously adjust their strategies based on contextual cues and past interactions.
-Over time, cooperative behaviors emerged even without explicit instructions to cooperate.
-The expert hybrid framework is another example of coopetitive collaboration.
-In this approach, different experts specialize in handling distinct aspects of the input problem.
-A gating mechanism evaluates the input and either selects the most appropriate expert or assigns weights to multiple experts.
-The final output is then produced as a weighted combination of their individual analyses.
-In LLM-based multi-agent systems, interaction mechanisms are often diverse rather than singular.
-Different agents may participate in various types of collaborative channels.
-These hybrid models combine the strengths of multiple interaction mechanisms, thereby enhancing the overall system performance and adaptability.
-5.1.2 Value Considerations and Future Directions in Multi-agent Interaction
-While multi-agent systems promote the emergence of collective intelligence, individual interests may still conflict with collective goals in social dilemmas [332].
-Even if each agent individually aligns with shared values, game-theoretic challenges can prevent the entire system from achieving overall value alignment [333].
-When analyzing the behaviors and value-related challenges that arise among multiple agents facing diverse task goals, varying information completeness, and some other conditions, game theory offers a robust theoretical framework.
-It enables designers to model and analyze interaction mechanisms and the dynamic behaviors of advanced agent systems [333].
-Game theory is a branch of mathematics that studies strategic interactions and decision-making among rational individuals [334].
-Its core value lies in uncovering the underlying principles of conflict and cooperation, enabling the analysis and prediction of individual behaviors.
-The design of MAS interaction mechanisms could potentially, based on game theory, lead to the emergence of novel values centered on inter-agent interaction.
-This would highlight the unique characteristics of agent interaction dynamics, distinguishing them from other AI technologies and from human interactive games.
-Compared to other AI technologies, Large language models possess strong capabilities in language-based reasoning, planning, and problem-solving [335, 336].
-These strengths enable LLM-based agents to simulate human-like traits such as personality and tone [337, 338] and demonstrate a degree of theory of mind—that is, the ability to infer others’ beliefs and intentions during strategic interactions [339].
-LLM-based agents can predict opponents’ psychological states and behaviors during interactions by considering task objectives, resource distribution, and personality configurations.
-This allows them to adjust allocation and equilibrium decisions.
-AI agents can therefore be regarded as fully rational individuals, and their interactions can be modeled using game theory.
-Compared to human interactive games, agent interaction games are characterized by numerous parameters, complex utility functions, and multiple game stages.
-While the computational power of agents amplifies the complexity of these games, it simultaneously makes them solvable.
-Future research can explore the values reflected in MAS interactions by focusing on three key activities related to their game objectives: agent identification, goal setting, and boundary constraints.
-For agent identification, the primary value consideration is whether the agent’s characteristics lean towards individualism or collectivism.
-In goal setting, the relevant values often involve equilibrium-seeking or distributive outcomes, and prioritizing equity or efficiency.
-Finally, for boundary constraints, key value considerations include information symmetry, the presence of multi-round interactions, and whether a short-term or long-term perspective is adopted.
-Given the differences in underlying value orientations, we discuss value-related issues from the perspective of the game-theoretic models.
-Non-cooperative game.
-Non-cooperative game theory models the interactions among agents driven by individual self-interest, with the primary focus on identifying equilibrium states.
-In such games, each agent acts as a self-interested decision-maker, expressing preferences over system outcomes through a utility function and iteratively adjusting its actions via optimal or near-optimal responses to maximize individual utility [340].
-This framework highlights the potential conflict between individual and collective objectives.
-Thus ensuring value alignment within multi-agent systems is critical—particularly when agents are expected to operate in accordance with human intentions in the face of social dilemmas.
-It is essential to guide rational agents toward effective cooperation and to discourage exploitative or free-riding behaviors to ensure that the system functions in line with human values [332].
-Research indicates that the intrinsic motivations of intelligent agents can be enhanced through modeling value orientations such as inequality aversion [341], altruism [342], social value orientation [343, 344, 345], social influence [346], and reputation [347].
-These approaches enable agents to account for the welfare of others alongside their own interests.
-In parallel, external motivation such as peer rewarding [348, 349] and the establishment of agreement [350] can incentivize cooperative behavior, encouraging agents to align with group values even while pursuing individual objectives.
-The pursuit of short-term or long-term interests, as well as whether the number of game rounds is limited, can significantly affect multi-agent game modeling.
-Nash game is a static, single-round game that focuses on short-term interests.
-At a Nash equilibrium, no agent can unilaterally deviate from its strategy to achieve a better outcome.
-However, the advanced computational capabilities of LLMs enable agents to process vast amounts of data and communicate in natural language, allowing for the simulation and analysis of complex, multi-round games that are difficult to model in human society.
-Markov games are multi-round, sequential decision-making games in which agents must consider not only immediate rewards but also long-term benefits by anticipating future environmental states and corresponding actions.
-This framework is well-suited to decision-making problems in dynamic environments, where optimal strategies depend on both current conditions and projected future developments.
-For example, Perolat et al.
-[351] examined the allocation of common-pool resources—such as fisheries, pastures, and irrigation systems—within the framework of non-cooperative games, illustrating how individual decision-making in dynamic environments can result in either the "tragedy of the commons" or cooperative outcomes.
-In a more applied context, Markov games have been used to model multi-vehicle interactions in autonomous driving [352].
-These models aim to shift agent (i.e., vehicle) behavior away from short-term reward maximization toward long-term benefit optimization—for instance, by slowing down or yielding at intersections to avoid potential future collisions, thereby improving overall driving strategies.
-Beyond the influence of short- or long-term interest orientations, information symmetry is also a key factor influencing multi-agent interactions and the design of regulatory governance.
-From a game-theoretic perspective, information asymmetry can create strategic advantages, as part of an agent’s payoff may derive from possessing superior information.
-Consequently, decision-making by agents is influenced by the distribution of information.
-Moreover, asymmetry in information can lead to signaling behaviors, where agents convey indirect information through their actions.
-Observing these signals allows other agents to infer hidden information, thereby adapting their own strategies accordingly.
-In games with incomplete or imperfect information, agents may deviate from normative value principles in pursuit of individual gain—for example, by withholding critical information to obtain short-term advantages at the expense of long-term trust, using misleading signals to distort others’ decisions, or exploiting information asymmetries to amplify systemic risk.
-To address these challenges, regulators must design mechanisms based on the principle of incentive compatibility [353].
-When the regulated entities are humans, regulatory bodies often struggle to obtain private information from companies.
-To address this, they design incentive-compatible mechanisms aimed at maximizing social welfare or other objectives.
-However, the situation changes when AI agents become the regulated entities.
-In the context of value alignment in multi-agent systems, the agents’ objective functions may implicitly reflect the designers’ values.
-Ensuring consistency between these embedded values and broader social goals introduces new constraints.
-One key consideration is that private information may take the form of data or model parameters.
-Unlike humans, artificial agents operate based on algorithms, often exhibiting higher rationality and information-processing capabilities.
-They may share information through collaborative techniques such as federated learning without compromising privacy, potentially enabling algorithmic collusion and altering the nature of information asymmetry.
-Additionally, we must account for the differences in behavioral patterns between agents and humans.
-While humans might misrepresent information due to personal gain or emotion, an agent’s decisions are based on predefined algorithms and objective functions, likely leading them to precisely compute optimal strategies.
-Consequently, the design of incentive-compatible mechanisms may require different mathematical modeling approaches, such as incorporating reward mechanisms from reinforcement learning.
-Furthermore, the dynamic and scalable nature of intelligent agent systems is another critical factor.
-Mechanisms might need to be adjusted in real-time to adapt to agents’ rapid learning and evolving environmental conditions.
-Cooperative game.
-Cooperative game theory provides a foundational framework for addressing issues of rights allocation and responsibility attribution in multi-agent system interactions.
-From a rights-based perspective, cooperative games are particularly relevant for the design of rule-based allocation mechanisms within the system.
-In practice, agents—often representing organizations with divergent objectives—may initially engage in non-cooperative interactions, such as the classic Prisoner’s Dilemma, leading to suboptimal or mutually disadvantageous outcomes.
-In response, agents and their affiliated enterprises may choose to form cooperative coalitions to enhance overall utility and achieve more favorable outcomes through coordinated strategies.
-Cooperative games offer an effective approach to addressing efficiency losses that arise from strategic interactions among multiple agents.
-The primary focus lies in maximizing collective utility and ensuring a fair distribution of gains among participants.
-Compared to the often complex task of identifying equilibrium solutions in non-cooperative settings, cooperative games typically reduce the problem to a more tractable optimization framework.
-When there is no loss associated with transfer payments among agents in the system, fairness serves as the primary principle guiding allocation.
-However, in cases where such transfers incur losses, both fairness and overall system efficiency must be jointly considered as key objectives.
-Li et al.
-[354] proposed a nucleolus-based credit assignment method grounded in cooperative game theory for dynamic coalition formation in multi-agent reinforcement learning.
-In this approach, agents are dynamically grouped into smaller coalitions to handle subtasks according to task demands.
-The nucleolus is employed to minimize the maximum dissatisfaction among coalitions, ensuring that each agent receives a fair reward for its contribution.
-This method effectively addresses the fairness and stability challenges of credit assignment, helping to prevent coalition fragmentation.
-From the perspective of responsibility attribution, cooperative game theory provides a valuable framework for analyzing how moral responsibility should be assigned in multi-agent systems.
-Assigning responsibility is a foundational step toward designing autonomous agents that comply with legal and ethical norms.
-Building on Halpern and Kleiman-Weiner’s [355] work on responsibility attribution in single-agent settings, Friedenberg and Halpern [356] extended the framework to multi-agent scenarios by introducing a definition of collective responsibility.
-In their model, a group of agents is treated as a single entity, and the degree of responsibility for an undesirable outcome is determined by comparing the probability of that outcome under different epistemic states, while also considering the cost of coordinated actions.
-They further employed the Shapley value from cooperative game theory to distribute the group’s responsibility among individual agents, based on each agent’s marginal contribution across all possible coalitions.
-Triantafyllou et al.
-[357] observed that some widely used attribution methods, such as those based on the Shapley value, may fail to incentivize agents to adopt better strategies, while others, such as the Banzhaf index, risk over-attributing responsibility.
-To address these limitations, they drew on concepts from cooperative game theory—including cost-sharing, data valuation, and coalition formation—to develop a framework for responsibility attribution in accountable multi-agent sequential decision-making.
-This line of work holds significant implications for enhancing fairness, transparency, and accountability in multi-agent systems.
-5.2 Value Considerations in the Design of Multi-agent Communication Structures Informed by Organizational Model
-5.2.1 Types of Communication Structure
-In a multi-agent system, the communication structure refers to the framework that governs interactions among agents to facilitate coordination and achieve shared or individual objectives.
-It defines the control relationships and communication mechanisms among agents, thereby influencing the system’s efficiency, adaptability, and scalability.
-Common communication structures include centralized, decentralized, and hierarchical models.
-In a centralized structure, all agents are connected to a central agent, which serves as a hub for managing interactions and coordinating collaboration among participants [112, 324].
-The central agent is primarily responsible for overseeing control, coordination, and task allocation.
-Centralized systems are relatively easy to design and implement.
-Through centralized control, they facilitate information aggregation, task scheduling, and the allocation of responsibilities and rights among agents, enabling globally efficient and optimized decision-making [358].
-However, this architecture is vulnerable to single points of failure: if the central node malfunctions, the entire system risks collapse.
-Additionally, as the number of agents increases, the central node may become a performance bottleneck [358].
-Some studies adopt a central agent to coordinate multi-LLM communication, integrating techniques such as pairwise ranking, visual interfaces with interactive features, knowledge extraction, and graph aggregation to enhance system performance in workload distribution, answer integration, coordination strategy design, querying and search-answering processes, and knowledge graph construction [359, 360, 361, 362, 363, 364].
-In decentralized multi-agent systems, control and decision-making are distributed among agents without relying on a central controller.
-Communication is based on a peer-to-peer model, where agents are directly connected and operate autonomously using local information and limited neighborhood communication.
-These systems commonly adopt topologies such as chains, rings, small-world networks, or random graphs [358, 365] .
-The decentralized architecture is highly adaptable, robust, and scalable.
-Its resilience to failure allows the system to function even if individual agents malfunction.
-However, decentralization introduces challenges.
-The lack of a central coordinator increases communication overhead and coordination complexity, as agents must engage in repeated negotiation and feedback cycles to mitigate information silos [324].
-Moreover, since agents primarily rely on local information, system-wide consistency and decision-making efficiency may suffer from information asymmetry and communication delays, making it difficult to achieve global optima [324].
-This architecture has been applied in diverse domains including world scenario simulation, domain-specific collaboration, and creative ideation [243, 366, 367].
-It is also suited for scenarios requiring autonomous agent decision-making like real-time optimization of large-scale traffic networks [368, 369].
-The hierarchical structure organizes agents in a tiered structure, where interactions occur between agents within the same layer or across adjacent layers.
-Higher-level agents coordinate or supervise lower-level ones [358], forming a top-down command and coordination pathway with feedback loops—resembling the relationship between management and execution layers in traditional organizations [366].
-Tasks are distributed across different layers, with each layer fulfilling distinct functions.
-Higher-level agents typically handle global planning, task decomposition, and oversight, while lower-level agents execute specific tasks and relay feedback upward.
-This structure facilitates a clear command chain and well-defined information flow, enabling orderly task allocation and efficient execution in complex systems [112, 370, 371].
-The modular organization and hierarchical control reduce dependence on a central node and support efficient resource allocation, with relatively low sensitivity to bandwidth constraints or coordination latency.
-However, hierarchical systems are complex and prone to latency.
-The reliability of edge devices significantly impacts system stability—failures at the lower tiers may lead to system-wide disruptions or functionality loss [324] .
-Moreover, bottlenecks or errors at higher levels may impair the overall coordination efficiency [370].
-5.2.2 Value Considerations and Future Directions in the Organizational Models
-The interaction mechanisms in multi-agent systems focus on pairwise competition or cooperation between agents, while communication structures adopt a system-wide perspective, emphasizing the complex network formed by all agents.
-As the foundational framework for coordination and information exchange, communication structure plays a critical role in shaping both system-wide values and the alignment of individual agents’ values.
-In exploring value-related issues arising from communication architectures in multi-agent systems, relevant theories from organizational management can offer valuable insights and guidance for structural design and value alignment.
-Organizational structure refers to the framework of an organization, reflecting how resources are allocated within it [372].
-Essentially, it integrates dispersed individuals or departments into a goal-oriented organic whole through the design of functional divisions, authority distribution, and collaboration rules.
-Centralized authority patterns and chains of command are important characteristics of organizational models.
-Organizational structure is a critical vehicle for achieving organizational goals.
-Its design directly influences members’ behavior patterns and value orientations [373].
-The design of communication architectures in multi-agent systems can, to some extent, correspond to characteristics of organizational models.
-This mapping forms a crucial foundation for applying organizational management theories to multi-agent system research.
-Organizational models primarily affect both value orientations and normative principles, as well as the degree of difficulty in achieving value alignment.
-First, the design of multi-agent system organizational structures will orient values in distinct ways.
-It’s crucial to note that the value orientation shaped by organizational structure design differs significantly from the fundamental setting of value principles.
-From a dynamic evolutionary perspective, organizational structure design itself will guide agent behavior to evolve in different directions.
-Unlike explicit, pre-set value rules, the implicit guiding effect produced by organizational structures exhibits path dependency and self-reinforcing characteristics.
-This has already been validated in agent simulation experiments conducted at the Stanford University Organizational Science Lab [374].
-Based on Burns and Stalker’s dichotomy of organizational structures [375], values within multi-agent systems can be categorized into those aligned with mechanistic structures and those with organic structures.
-These two types of structures exhibit distinct value characteristics.
-In a mechanistic structure’s value propagation mechanism, three key aspects stand out:
-  • Emphasis on vertical authority chains reinforces collectivism. This mimics a bureaucratic, centralized structure, where value objectives are broken down via a chain of command.
-  • Standardized processes solidify value judgments. Pre-set operational norms are internalized as default value principles through repetitive execution.
-  • Limited information channels narrow value perception. Information filtering between hierarchical levels can lead to fragmented value judgments by agents at the lower end.
-In the emergence of values within an organic structure, three key characteristics stand out:
-  • Networked collaboration fosters adaptive values. A flatter structure promotes cross-unit interaction, which helps form collaborative values like "shared risk."
-  • Dynamic role allocation shapes a sense of responsibility.
-Temporary empowerment mechanisms, based on task demands, enable agents to develop context-specific value judgment capabilities.
-This might even lead to a stronger sense of moral responsibility than agents in mechanistic structures.
-  • Open information flow expands the value horizon.
-A fully connected communication network encourages the integration of diverse value perceptions, thereby enhancing agents’ ability to resolve complex ethical dilemmas.
-Second, organizational structure and values will engage in an interactive evolutionary process.
-For instance, in the value-anchoring effect of authority and responsibility allocation, contingency theory [376] suggests that decision-making power distribution directly impacts value weighting.
-A centralized structure locks in value priorities through a "means-end" chain, while a decentralized structure forms a dynamic value equilibrium via local interactions.
-Beyond this, the structure of communication networks will also exhibit a value diffusion process.
-Based on social network analysis theory [377], a star network significantly increases the speed of value dissemination from core nodes, but at the cost of reduced diversity.
-In contrast, fully connected networks show a slower rate of value convergence, yet they enhance the ethical soundness of solutions.
-Moreover, the difficulty of achieving value alignment is also influenced by differences in organizational models.
-Specifically, in mechanistic structure, a central agent coordinates and controls the behaviors and goals of other agents, concentrating decision-making authority.
-Non-central agents must strictly obey the central agent’s commands.
-Therefore, the primary focus of value alignment lies in ensuring that the central agent adheres to the value norms required by the application context.
-Misalignment at the center can lead to system-wide behaviors that deviate from human interests.
-In contrast, agents in organic structure communicate peer-to-peer and make decisions autonomously based on their individual goals and local information [378].
-The absence of a central controller complicates value alignment.
-On one hand, decentralization enhances resilience against adversarial attacks or localized value contamination since deviations in a single agent’s values have limited impact on the overall system.
-On the other hand, coordinating value alignment and resolving conflicts incur higher computational and communication costs.
-Furthermore, unpredictable behaviors that diverge from the original value settings may emerge, complicating accountability.
-Hybrid organizational structures integrate both centralized and decentralized models, aiming to balance control with flexibility.
-Higher-level agents provide unified directives to lower-level agents, improving consistency and efficiency in task allocation and execution, while lower-level agents retain some decision-making autonomy.
-For value alignment, a unified set of value norms can be embedded at the top level to guide decision-making throughout the system.
-However, challenges arise in transmitting values accurately across levels—distortions or oversimplifications may occur.
-Additionally, conflicts in values between different layers may require trade-offs and reconciliation.
-5.3 Multi-Level Value Evaluation Framework and Dataset for LLM-Based Agent Systems
-5.3.1 Establishment of a Multi-Level Agent Value Alignment Evaluation System
-In the intelligent era, where AI deeply intervenes in public affairs decision-making, value alignment has transcended a mere technical optimization category, evolving into a core proposition for modernizing social governance systems.
-While large-scale agent clusters enhance social operational efficiency, their value misalignments can trigger systemic governance crises.
-A multi-level agent value alignment evaluation system aims to address the issue of agent "goal conflicts" from a social governance perspective.
-This involves harmonizing values across macroscopic, mesoscopic, and microscopic levels, comprehensively considering the conflicts between "individual optimum" and "collective optimum" under varying resource constraints.
-From a social governance perspective, the construction of a value alignment framework manifests as different levels of governance entities participating in the evolutionary process of value alignment.
-Governments are responsible for benchmarking values, establishing indicator systems for principles such as public safety, social equity, and privacy protection.
-Industry organizations, leveraging their unique characteristics, develop industry-specific value standards to foster the healthy development of agents within their sectors.
-Lastly, enterprises or other entities developing AI agents are tasked with forming goal-oriented values tailored to specific application scenarios, ensuring the achievement of concrete objectives.
-Among these three governance entities, governments and industry organizations primarily engage in external oversight of value alignment in agent systems.
-Conversely, enterprises or other units developing AI agents are responsible for internal value configuration within these systems.
-Future research can focus on building a governance ecosystem that simultaneously satisfies the value alignment requirements of different levels of governing bodies.
-Regarding cross-cultural alignment theory, this involves establishing a value mapping system applicable to diverse cultures, addressing the compatibility challenges of varied cultural algorithms.
-For value evolution prediction algorithms, research can develop models based on big data of societal changes to predict value evolution, enabling proactive alignment of AI agent systems.
-Additionally, it’s worth noting that current research on value alignment for application-oriented LLM agent systems primarily focuses on generalized scenarios.
-As agents become more integrated into enterprise organizations, they’ll encounter more specific and specialized business contexts.
-This will necessitate greater consideration of micro-level value alignment issues.
-5.3.2 Open Sharing of High-Quality Value Alignment Data
-At the critical juncture where AI is becoming deeply integrated into societal operations, the open sharing of high-quality data plays a pivotal role in facilitating value alignment for AI agents from the perspective of social governance.
-With the increasing trend of privatized AI deployment, barriers erected by data privatization have led to the fragmentation of high-quality data and exacerbated the dilemma of “localized value alignment” in Agentic AI Systems.
-Therefore, this study advocates for a forward-looking balance: clearly delineating the boundaries of core proprietary data while promoting the open sharing of high-quality datasets for value alignment.
-This approach is identified as a key pathway for enabling value alignment across macro, meso, and micro levels.
-Constructing high-quality datasets for value alignment is fundamentally about ensuring the credibility and universality of values.
-This requires that data authentically, comprehensively, and deeply reflect the core ethical principles universally recognized at the macro level (e.g., responsibility, fairness and justice, trust), the policy and regulatory frameworks and cultural orientations at the meso level, and the critical normative requirements specific to concrete application scenarios at the micro level.
-The construction of such datasets necessitates the establishment of rigorous provenance and verification mechanisms to ensure the ethical legitimacy of data sources.
-Promoting the open sharing of high-quality value alignment data centers on the continuous transmission of values.
-To this end, a primary objective for future research is to extract the core logic of ethical decision-making from proprietary data and transform it into securely shareable formats.
-In this process, it is essential to precisely preserve key ethical features embedded in the data (e.g., “fairness,” “harmlessness”) to ensure the authenticity and reliability of the shared values.
-Moreover, effective data sharing must be grounded in accurate identification of data attributes and scenario-based authorization management, thereby establishing a trustworthy sharing mechanism.
-Sustaining a long-term ecosystem of open sharing depends on fostering a virtuous cycle of interaction between contributors and beneficiaries.
-The focus of sharing should shift from the competition over “data ownership” to the operationalization of “value-based service capabilities,” ultimately transforming open sharing into a self-sustaining infrastructure that balances commercial viability with ethical objectives.
-6 Conclusion
-This review deeply explores value alignment in LLM-based agent systems, integrating AI development with the demands of social governance.
-It systematically outlines hierarchical value principles from macro to micro levels, surveys diverse application scenarios, and evaluates value alignment datasets and methodologies.
-Looking ahead, we advocate for establishing a multi-level evaluation system and emphasize value coordination among multiple agents, particularly in designing interaction mechanisms and organizational models.
-We also suggest sharing micro-level value propositions from enterprise datasets, contingent on maintaining confidentiality.
-This review aims to provide a thorough understanding of value alignment in LLM agent systems from the perspective of practical deployment and application, offering thought-provoking insights to foster deeper exploration in this field.